### v2

# A Construction of Infinite Families of Self-Orthogonal Quasi-Cyclic Codes Using Constituent Codes

Abstract.
Quasi-cyclic codes have been recently employed in the constructions of quantum error-correcting codes. In this paper, we propose a construction of infinite families of quasi-cyclic codes which are self-orthogonal with respect to the Euclidean and Hermitian inner products. In particular, their dimension and a lower bound for their minimum distance are computed using their constituent codes defined over field extensions of $\mathbb{F}_{q}$. We also show that the lower bound for the minimum distance satisfies the square-root-like lower bound and also show how self-dual quasi-cyclic codes can arise from our construction. Using the CSS construction, we show the existence of quantum error-correcting codes with good parameters.
Key words and phrases:
quasi-cyclic codes, self-dual codes, self-orthogonal codes, CSS construction, quantum error correcting codes
2020 Mathematics Subject Classification:
94B60
1. Introduction
Quantum error correction plays a central role in the development of quantum computing architectures. With the rapid advancement of technologies involving quantum computation and communication, there is a growing demand for effective quantum error-correcting codes (QECCs). These codes are essential for protecting quantum information from noise, which can arise from various types of quantum channels. In the celebrated work of Calderbank, Shor, and Steane [Calderbank-Shor1996, Steane1996], the authors proposed the CSS construction that utilizes the duality of certain classical error-correcting codes to construct quantum error-correcting codes. This method is commonly used to construct QECCs over finite fields [Ketkar-Klappenecker-Kumar-Sarvepalli2006]. QECCs from the CSS construction are most often constructed using self-orthogonal error-correcting codes and its dual, where self-orthogonal codes have applications in lattice theory [W98], row-self-orthogonal matrices [Massey92, Massey98], linear complementary dual codes [Massey98], and many others. Beyond self-orthogonality, the notions of dual-containing and self-dual codes are also significant. Self-dual codes notably exhibit deep connections with group theory [Hoehn00, JLLX13], lattices [SloaneChapter99], and design theory [Sole90], and are recognized for being asymptotically good [MACWILLIAMS1972153]. Because of these applications, it is a worthwhile endeavor to determine the existence and constructions of self-orthogonal, dual-containing, and self-dual classical error-correcting codes.
The focus of this paper is on constructing quasi-cyclic (QC) error-correcting codes that are self-orthogonal and self-dual under the Euclidean and Hermitian inner products. Quasi-cyclic codes are a natural generalization of cyclic codes and are well-studied in classical coding theory literature. Quasi-cyclic codes have a rich algebraic structure [seguin, cem, lingsole] and have strong connections with convolutional codes [SolomonTilborg]. Moreover, there exist QC codes that are asymptotically good [longcodes]. In 2022, the authors in [RB] found an upper bound for the locality of QC locally recoverable codes using generator matrices in systematic form where the first block is the identity matrix and the remaining blocks are comprised of smaller circulant matrices. More recently in  [BastosAlvarezFloresSalerno], a more concise way of finding an upper bound on the locality of these codes was studied, and a construction of optimal locally recoverable QC codes was presented.
Very recently, some constructions of self-dual, dual-containing, and self-orthogonal QC codes have been explored to develop effective quantum error-correcting codes. For example in [Abdukhalikov-Bag-Panario2023, GalindoHernando, benjwal], the authors obtained quantum codes with good parameters by studying the generators of the dual of a QC code to determine the conditions where such codes are dual-containing, self-orthogonal, or self-dual. Consequently, as these properties are intrinsically related to the CSS construction, QECCs with good parameters are obtained.
In this paper, we propose a construction of infinite families of self-orthogonal and self-dual QC codes using their constituent code representation described in [cem]. We then describe how the codes in our construction yield the existence of QECCs with good parameters. The paper is organized as follows: Section 2 provides background results related to QC codes and their constituent codes. This section also includes the foundations of constructing QECCs from classical ones. In Section 3, we present the main results of this paper, namely, the use of constituent codes of QC codes to construct infinite families of QC codes that are self-orthogonal (or self-dual) with respect to the Euclidean inner product. We show that the self-orthogonal QC codes in this construction satisfy the square-root-like lower bound to their minimum distance. We also introduce the notion of Galois-closed codes to obtain QC codes that are self-orthogonal with respect to the Hermitian inner product. Lastly in Section 4, we show how some codes from our construction are used to obtain quantum error-correcting codes with the best known parameters so far. All computations throughout this paper have been done using SageMath or Mathematica.
2. Preliminaries
2.1. Quasi-Cyclic Codes and Their Representations
Let $\mathbb{F}_{q}$ be a finite field with $q$ elements, where $q=p^{t}$, for $p$ a positive prime and $t\in\mathbb{Z}^{+}$. An $n,k,d]_{q}-$linear code $C$ is a $k$-dimensional $\mathbb{F}_{q}$-vector subspace of $\mathbb{F}_{q}^{n}$, where the minimum weight of nonzero codewords is $d$. In this work, only linear codes are presented, so we will refer to them as codes only. The relationship between the parameters $n,k$ and $d$ is the following Singleton bound
(1) $d\leq n-k+1$
and the code is called a maximum distance separable code, or an MDS code, if equality in (1) is achieved. A quintessential example of a maximum distance separable code is the Reed-Solomon code described in [RS1958]. More generally, if we choose nonzero elements $v_{1},...,v_{n}\in\mathbb{F}_{q}$ and distinct elements $\alpha_{1},...,\alpha_{n}\in\mathbb{F}_{q}$, we define the Generalized Reed-Solomon Codes as
(2) $GRS_{n,k}(\mathbf{\alpha},\mathbf{v})=\{(v_{1}f(\alpha_{1}),...,v_{n}f(\alpha_{n}))\mid f(x)\in\mathbb{F}_{q}[x]_{k}\}$
where $\textbf{v}=(v_{1},...,v_{n})$, $\textbf{$\alpha$}=(\alpha_{1},...,\alpha_{n})$, and $\mathbb{F}_{q}[x]_{k}$ is the set of polynomials in $\mathbb{F}_{q}[x]$ of degree less than $k$. This family of codes is another example of an MDS code that will be useful in Example 3.8.
This paper focuses on quasi-cyclic codes, which are generalizations of cyclic codes. More formally, a quasi-cyclic (QC) code of index $\ell$ is a code $\mathcal{C}\subset\mathbb{F}_{q}^{m\ell}$ where for all $\textbf{c}\in\mathcal{C}$, $T^{\ell}(\textbf{c})\in\mathcal{C}$, where $T$ is the standard shift operator on $\mathbb{F}_{q}^{m\ell}$ and $\ell$ is the smallest positive integer with this property. If $\mathcal{C}$ is a QC code in $\mathbb{F}_{q}^{m\ell}$, then any codeword $\textbf{c}\in\mathcal{C}$ can be represented as an $m\times\ell$ array
(3) $\textbf{c}=\begin{pmatrix}c_{00}&c_{01}&...&c_{0,\ell-1}\\ c_{10}&c_{11}&...&c_{1,\ell-1}\\ \vdots&\vdots&\ddots&\vdots\\ c_{m-1,0}&c_{m-1,1}&...&c_{m-1,\ell-1}\end{pmatrix}$
where the action of the operator $T^{\ell}$ is equivalent to a row shift in (3). We can also “flatten the array” by its rows to obtain the vector
(4) $\mathbf{c}=(c_{0,0},c_{0,1},\ldots,c_{0,\ell-1},c_{1,0},c_{1,1},\ldots,c_{1,\ell-1},\ldots,c_{m-1,0},c_{m-1,1},\ldots,c_{m-1,\ell-1})\in\mathbb{F}_{q}^{m\ell}.$
Moreover, as in [lingsole], we consider the ring $R_{q,m}^{\ell}:=\left(\mathbb{F}_{q}[x]/\langle x^{m}-1\rangle\right)^{\ell}$ where $\operatorname{gcd}(m,q)=1$. To any codeword c as in (3), we associate an element of $R_{q,m}^{\ell}$:
(5) $\vec{\textbf{c}}(x):=\left(c_{0}(x),c_{1}(x),...,c_{\ell-1}(x)\right)\in R_{q,m}^{\ell}$
where for each $0\leq j\leq\ell-1$,
(6) $c_{j}(x):=c_{0j}+c_{1j}x+c_{2j}x^{2}+\cdots+c_{m-1,j}x^{m-1}\in R_{q,m}.$
By Lemma 3.1 in [lingsole], the map $\phi:\mathbb{F}_{q}^{m\ell}\to R_{q,m}^{\ell}$ defined by $\phi(\textbf{c})=\vec{\textbf{c}}(x)$ is an $R$-module isomorphism. Hence, any QC code over $\mathbb{F}_{q}$ of length $m\ell$ and index $\ell$ can be viewed as an $R$-submodule of $R_{q,m}^{\ell}$. Throughout this paper, we will move freely between the representations in (4) and (5).
Given a polynomial $f(x)=f_{n}x^{n}+f_{n-1}x^{n-1}+\cdots+f_{0}\in\mathbb{F}_{q}[x]$, its reciprocal polynomial, $f^{*}$, is given by $f^{*}(x)=x^{n}f(x^{-1})$. As seen in [lingsole], we may decompose $x^{m}-1\in\mathbb{F}_{q}[x]$ into irreducible factors
(7) $x^{m}-1=\delta\prod_{i=1}^{t}g_{i}(x)g_{i}^{*}(x)\prod_{i=1}^{s}f_{i}(x),$
where $\delta$ is a non-zero element in $\mathbb{F}_{q}$, $f_{i}(x)$ is a self-reciprocal polynomial for each $1\leq i\leq s$, and $g_{i}^{*}(x)$ is the reciprocal polynomial of $g_{i}(x)$, where $g_{i}(x)\neq g_{i}^{*}(x)$, for each $1\leq i\leq t$. With this decomposition and via the Chinese remainder theorem (CRT), the ring $R_{q,m}$ may be written as
(8) $R_{q,m}=\frac{\mathbb{F}_{q}[x]}{\langle x^{m}-1\rangle}\cong\Bigg(\bigoplus_{j=1}^{t}\bigg(\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}\bigg)\Bigg)\oplus\bigg(\bigoplus_{i=1}^{s}\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\bigg),$
where we have suppressed the explicit dependence on $x$ in the $g_{j}$ and $f_{i}$. Let $\alpha$ be an $m$-th primitive root of the unity in an extension of $\mathbb{F}_{q}$ and let
$F_{i}:=\dfrac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\cong\mathbb{F}_{q}(\alpha^{u_{i}}),\quad G^{\prime}_{j}:=\dfrac{\mathbb{F}_{q}[x]}{\langle g_{i}\rangle}\cong\mathbb{F}_{q}(\alpha^{v_{j}}),\quad\text{and}\quad G^{\prime\prime}_{j}:=\dfrac{\mathbb{F}_{q}[x]}{\langle g^{*}_{j}\rangle}\cong\mathbb{F}_{q}(\alpha^{v^{*}_{j}}).$
for each $1\leq i\leq s$ and $1\leq j\leq t$. Hence,
(9) $R_{q,m}^{\ell}\cong\Bigg(\bigoplus_{j=1}^{t}(G^{\prime}_{j}\oplus G^{\prime\prime}_{j})^{\ell}\Bigg)\oplus\bigg(\bigoplus_{i=1}^{s}F_{i}^{\ell}\bigg).$
Define codes $C_{i}\subset F_{i}^{\ell}$, $C^{\prime}_{j}\subset G_{j}^{{}^{\prime}\ell}$, and $C^{\prime\prime}_{j}\subset G_{j}^{{}^{\prime\prime}\ell}$ for $1\leq i\leq s$ and $1\leq j\leq t$. We call these codes the constituent codes of the $R_{q,m}$-QC code $\mathcal{C}\subset R_{q,m}^{\ell}$ which has the following CRT decomposition:
(10) $C\cong\Bigg(\bigoplus_{j=1}^{t}(C_{j}^{\prime}\oplus C_{j}^{\prime\prime})\Bigg)\oplus\bigg(\bigoplus_{i=1}^{s}C_{i}\bigg).$
The following lemma gives an easy way to compute the dimension of a QC code $\mathcal{C}$ from its constituent codes. We slightly modified the notation of the following result from [BastosAlvarezFloresSalerno] to align with the notation in this paper.
Lemma 2.1.
[BastosAlvarezFloresSalerno, Lemma 2] Let $\mathcal{C}\subset\mathbb{F}_{q}^{m\ell}$ be an $m\ell,k]$ QC code and let $C_{j}\subseteq\mathbb{F}_{q}^{\ell}(\alpha^{u_{j}})\cong\left(\frac{\mathbb{F}_{q}[x]}{\langle b_{j}\rangle}\right)^{\ell}$ be its respective constituent $\ell,k_{j}]$ codes. Suppose that the nontrivial constituent codes of $\mathcal{C}$ are given by $C_{1},\ldots,C_{h}$. Then,
(11) $k=\sum_{j=1}^{h}k_{j}\deg(b_{j}).$
Let $s\in\mathbb{Z}_{m}$ and let $C_{s}=\{s,sq,sq^{2},...,sq^{r-1}\}\subset\mathbb{Z}_{m}$ be the $q$-cyclotomic coset of $s$, where $r$ is the smallest positive integer so that $sq^{r}\equiv s\mod m$. It is important to note that the $q$-cyclotomic cosets split $\mathbb{Z}_{m}.$ This classic concept is seen in constructions of cyclic codes, as well as in the next theorem (formally known as the trace representation [lingsole] of QC codes) where we slightly modified the notation in the statement to better align with the notation in this paper. Moreover, if $\mathbb{F}_{q}$ is a field and $\mathbb{F}_{q^{s}}$ a a field extension of $\mathbb{F}_{q}$, the trace map $\operatorname{Tr}_{\mathbb{F}_{q^{s}}/\mathbb{F}_{q}}:\mathbb{F}_{q^{s}}\to\mathbb{F}_{q}$ is defined as
(12) $\operatorname{Tr}_{\mathbb{F}_{q^{s}}/\mathbb{F}_{q}}(a):=\sum_{i=0}^{s-1}a^{q^{i}}.$
Theorem $2.$
[lingsole, Theorem 5.1] Let $\mathbb{F}_{q}$ be a finite field and $\gcd(m,q)=1$. Then, for any $\ell$, the quasi-cyclic codes over $\mathbb{F}_{q}$ of length $m\ell$ and of index $\ell$ are precisely given by the following construction. Decompose $x^{m}-1$ as per (7) . Write $\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}=F_{i}$, $\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}=G_{j}^{\prime}$, and $\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}=G_{j}^{\prime\prime}$. Let $U_{i}$ (resp. $V_{j}$ and $W_{j}$) denote the cyclotomic coset of $\mathbb{Z}/m\mathbb{Z}$ corresponding to $F_{i}$ (resp. $G_{j}^{\prime}$ and $G_{j}^{\prime\prime}$) and fix $u_{i}\in U_{i}$, $v_{j}\in V_{j}$, and $w_{j}\in W_{j}$. For each $i$, let $C_{i}$ be a code of length $\ell$ over $F_{i}$, and for each $j$, let $C_{j}^{\prime}$ be a code of length $\ell$ over $G_{j}^{\prime}$ and let $C_{j}^{\prime\prime}$ be a code of length $\ell$ over $G_{j}^{\prime\prime}$. For $\mathbf{x_{i}}\in C_{i}$, $\mathbf{y_{j}^{\prime}}\in C_{j}^{\prime}$, and $\mathbf{y_{j}^{\prime\prime}}\in C_{j}^{\prime\prime}$, and for each $0\leq g\leq m-1$, let
(13) $c_{g}((\mathbf{x_{i}}),(\mathbf{y_{j}^{\prime}}),(\mathbf{y_{j}^{\prime\prime}}))=\sum_{i=1}^{s}\operatorname{Tr}_{F_{i}/\mathbb{F}_{q}}(\mathbf{x_{i}}\alpha^{-gu_{i}})+\sum_{j=1}^{t}\left(\operatorname{Tr}_{G_{j}^{\prime}/\mathbb{F}_{q}}(\mathbf{y_{j}^{\prime}}\alpha^{-gv_{j}})+\operatorname{Tr}_{G_{j}^{\prime\prime}/\mathbb{F}_{q}}(\mathbf{y_{j}^{\prime\prime}}\alpha^{-gw_{j}})\right),$
where $\alpha$ is a primitive $m$-th root of unity in some extension of $\mathbb{F}_{q}$. Then the code
(14) $\mathcal{C}=\{(c_{0}((\mathbf{x_{i}}),(\mathbf{y_{j}^{\prime}}),(\mathbf{y_{j}^{\prime\prime}})),...,c_{m-1}((\mathbf{x_{i}}),(\mathbf{y_{j}^{\prime}}),(\mathbf{y_{j}^{\prime\prime}}))):\forall\;\mathbf{x_{i}}\in C_{i},\forall\;\mathbf{y_{j}^{\prime}}\in C_{j}^{\prime},\mbox{ and }\forall\;\mathbf{y_{j}^{\prime\prime}}\in C_{j}^{\prime\prime}\}$
is a quasi-cyclic code over $\mathbb{F}_{q}$ of length $m\ell$ and of index $\ell$. Conversely, every quasi-cyclic code over $\mathbb{F}_{q}$ of length $m\ell$ and of index $\ell$ is obtained through this construction.
$2.$ The Duals of Quasi-Cyclic Codes
Recently, QC codes have been used in constructions of QECCs where orthogonality conditions on QC codes play a central role [Abdukhalikov-Bag-Panario2023, GalindoHernando, benjwal] . To study the duality of codes in $\mathbb{F}_{q}^{m\ell}$, we define the following inner products: Let $\textbf{a},\textbf{b}\in\mathbb{F}_{q}^{m\ell}$ where
a $=(a_{00},a_{01},...,a_{0,\ell-1},a_{10},...,a_{1,\ell-1},...,a_{m-1,0},...,a_{m-1,\ell-1})$
b $=(b_{00},b_{01},...,b_{0,\ell-1},b_{10},...,b_{1,\ell-1},...,b_{m-1,0},...,b_{m-1,\ell-1}).$
The Euclidean inner product on $\mathbb{F}_{q}^{m\ell}$ is
(15) $\mathbf{a}\cdot\mathbf{b}=\langle\textbf{a},\textbf{b}\rangle_{E}=\sum_{i=0}^{m-1}\sum_{j=0}^{\ell-1}a_{ij}b_{ij}.,$
and the Hermitian inner product on $\mathbb{F}_{q^{t}}^{m\ell}$, for even $t$ is
(16) $\langle\textbf{a},\textbf{b}\rangle_{H}=\sum_{i=0}^{m-1}\sum_{j=0}^{\ell-1}a_{ij}b_{ij}^{\sqrt{q^{t}}}.$
In particular, when $\textbf{a},\textbf{b}\in\mathbb{F}_{q^{2}}^{m\ell}$, then $\langle\textbf{a},\textbf{b}\rangle_{H}=\sum_{i=0}^{m-1}\sum_{j=0}^{\ell-1}a_{ij}b_{ij}^{q}$.
On $R^{\ell}$, the Hermitian inner product is as follows: for $\textbf{x}=(x_{0},x_{1},...,x_{\ell-1})$, and $\textbf{y}=(y_{0},y_{1},...,y_{\ell-1})$,
(17) $\langle\textbf{x},\textbf{y}\rangle_{H}=\sum_{j=0}^{\ell-1}x_{j}\overline{y_{j}},$
where $\bar{\;}$ denotes the “conjugation map” on $R$ that sends $x$ to $x^{-1}=x^{m-1}$ and is extended $\mathbb{F}_{q}$-linearly. The following proposition is useful for determining when orthogonality in $R^{\ell}$ implies orthogonality in $\mathbb{F}_{q}^{m\ell}$.
Proposition 2.3.
[lingsole, Proposition 3.2] Let $\mathbf{a},\mathbf{b}\in\mathbb{F}_{q}^{m\ell}$. Then $(T^{k\ell}(\mathbf{a}))\cdot\mathbf{b}=0$ for all $0\leq k\leq m-1$ if and only if $\langle\phi(\mathbf{a}),\phi(\mathbf{b})\rangle_{H}=0$, where $\phi$ is the aforementioned $R$-module isomorphism.
Let $\mathcal{C}$ be a QC code in $\mathbb{F}_{q}^{m\ell}$ and define the Euclidean dual, $\mathcal{C}^{\perp_{E}}$ of the code $\mathcal{C}$ by
(18) $\mathcal{C}^{\perp_{E}}:=\{\textbf{x}\in\mathbb{F}_{q}^{m\ell}\mid\langle\textbf{x},\textbf{c}\rangle_{E}=0,\forall\;\textbf{c}\in\mathcal{C}\}.$
Similarly, the Hermitial dual, $\mathcal{C}^{\perp_{H}}$ of the code $\mathcal{C}$ is given by
(19) $\mathcal{C}^{\perp_{H}}:=\{\textbf{x}\in\mathbb{F}_{q^{t}}^{m\ell}\mid\langle\textbf{x},\textbf{c}\rangle_{H}=0,\forall\;\textbf{c}\in\mathcal{C}\}.$
We say that $\mathcal{C}$ is self-orthogonal (ESO) with respect to the Euclidean inner product if $\mathcal{C}\subseteq\mathcal{C}^{\perp_{E}}$, dual-containing (EDC) with respect to the Euclidean inner product if $\mathcal{C}^{\perp_{E}}\subseteq\mathcal{C}$, and self-dual (ESD) with respect to the Euclidean inner product if $\mathcal{C}^{\perp_{E}}=\mathcal{C}$. Similarly, $\mathcal{C}$ is self-orthogonal (HSO), dual-containing (HDC), and self-dual (HSD) with respect to the Hermitian inner product if $\mathcal{C}\subseteq\mathcal{C}^{\perp_{H}},\mathcal{C}^{\perp_{H}}\subseteq\mathcal{C}$, and $\mathcal{C}=\mathcal{C}^{\perp_{H}}$, respectively.
Corollary 2.4.
[lingsole, Corollary 3.3] Let $\mathcal{C}$ be a quasi-cyclic code over $\mathbb{F}_{q}$ of length $m\ell$ and of index $\ell$ and let $\phi(\mathcal{C})$ be its image in $R_{q,m}^{\ell}$ under $\phi$. Then $\phi(\mathcal{C})^{\perp}=\phi(\mathcal{C}^{\perp})$, where the dual in $\mathbb{F}_{q}^{m\ell}$ is taken with respect to the Euclidean inner product, while the dual in $R_{q,m}^{\ell}$ is taken with respect to the Hermitian inner product. In particular, a quasi-cyclic code $\mathcal{C}$ over $\mathbb{F}_{q}$ is self-dual with respect to the Euclidean inner product if only if $\phi(\mathcal{C})$ is self-dual over $R_{q,m}$ with respect to the Hermitian inner product.
The following result, which we slightly modified to better align with the notation in this paper, allows us to characterize the duality of QC codes using constituent codes and will be used throughout this paper.
Theorem 2.5.
[lingsole, Theorem 4.2] An $\ell$ quasi-cyclic code $\mathcal{C}$ of length $m\ell$ over $\mathbb{F}_{q}$ is self-dual with respect to the Euclidean inner product if and only if
(20) $\mathcal{C}\cong\left(\bigoplus_{j=1}^{t}(C_{j}^{\prime}\oplus(C^{\prime}_{j})^{\perp_{E}})\right)\oplus\left(\bigoplus_{i=1}^{s}C_{i}\right)$
where for $1\leq i\leq s$, $C_{i}$ is a self-dual code over $F_{i}$ of length $\ell$ (with respect to the Hermitian inner product) and, for $1\leq j\leq t$, $C^{\prime}_{j}$ is a code of length $\ell$ over $G^{\prime}_{j}$ and $C_{j}^{{}^{\prime\prime}}=(C^{\prime}_{j})^{\perp_{E}}$ is its dual with respect to the Euclidean inner product.
Remark 1.
  1. (i)
    As observed in [HKS, page 136], the irreducible factors $f_{i}$ which are self-reciprocal imply that the cardinality $q_{i}$ of each field extension $F_{i}$ is an even power of $q$, for all ii. However, there are two exceptions. The first one, for all $m$ and $q$, is the field extension coming from the irreducible factor x1x-1 of $x^{m}-1$. The other exception occurs when $q$ is odd and $m$ is even in which $x+1$ is another self-reciprocal factor of $x^{m}-1$. In both exceptions, $q_{i}=q$ and we equip these fields with the Euclidean inner product. Otherwise, we equip $F_{i}$ with the Hermitian inner product.
  2. (ii)
    When the constituent codes $C_{i}$ over finite extension fields coming from self-reciprocal factors are self-dual, self-orthogonal, or dual-containing (under the Hermitian or Euclidean inner products), for all $1\leq i\leq s$, then the respective $\ell$-QC code $\mathcal{C}$ is self-dual, self-orthogonal, or dual-containing, respectively. Therefore, the duality of $\mathcal{C}$ is managed by the duality of its constituent codes over finite extension fields obtained from self-reciprocal polynomials.
The next result extends [lingsole, Theorem 4.2] to the dual of a QC code.
Proposition 2.6.
[HKS, Proposition 7.3.5] Let $\mathcal{C}$ be a QC code with CRT decomposition as (10). Then its dual code $\mathcal{C}^{\perp}$ (under the Euclidean inner product) is of the form
(21) $\mathcal{C}^{\perp_{E}}\cong\left(\bigoplus_{j=1}^{t}(C_{j}^{\prime}\oplus C_{j}^{{}^{\prime}\perp_{E}})\right)\oplus\left(\bigoplus_{i=1}^{s}C_{i}^{\perp_{H}}\right)$
Since the duality of $\mathcal{C}$ can be determined from its constituent codes, the following results provide insight to extended constructions of self-orthogonal codes.
Lemma 2.7.
[HillNewton, Lemma 2.10] Suppose $C_{1}$ and $C_{2}$ are $n_{1},k,d_{1}]$ and $n_{2},k,d_{2}]$ codes over $\mathbb{F}_{q}$, respectively. If $C_{1}$ and $C_{2}$ have generator matrices $G_{1}$ and $G_{2}$, respectively, then $G_{1}|G_{2}]$ generates an $n_{1}+n_{2},k,d_{1}+d_{2}]$ code over $\mathbb{F}_{q}$ .
In particular, a version of Lemma 2.7 regarding self-orthogonal codes is presented below.
Theorem 2.8.
[Li, Theorem 1]
  1. (i)
    Suppose $C_{1}$ and $C_{2}$ are $n_{1},k,d_{1}]$ and $n_{2},k,d_{2}]$ self-orthogonal codes over $\mathbb{F}_{q}$, respectively. If $C_{1}$ and $C_{2}$ have generator matrices $G_{1}$ and $G_{2}$, respectively, then $G_{1}|G_{2}]$ generates an $n_{1}+n_{2},k,d_{1}+d_{2}]$ self-orthogonal code over $\mathbb{F}_{q}$.
  2. (ii)
    Suppose $C_{1}$ and $C_{2}$ are $n_{1},k,d_{1}]$ and $n_{2},k-1,d_{2}]$ self-orthogonal codes over $\mathbb{F}_{q}$, respectively. If $C_{1}$ contains a codeword of weight at least $d_{1}+d_{2}$, then there exists an $n_{1}+n_{2},k,d_{1}+d_{2}]$ self-orthogonal code over $\mathbb{F}_{q}$.
In the special case when $C_{1}=C_{2}$, the code with generator matrix $G_{1}|G_{1}]$ will be denoted as $2C_{1}$, and such an argument may be naturally extended to $mC_{1}$, for any positive integer $m$. In this work, for any code $C$, we will refer to the code mCmC as $m$-copies of $C$ which will play an important role in the next section.
2.3. A Lower Bound of the Minimum Distance of a QC Code
Let $C_{1}\subset\mathbb{F}_{q}^{\ell}(\alpha^{u_{a_{1}}}),...,C_{h}\subset\mathbb{F}_{q}^{\ell}(\alpha^{u_{a_{h}}})$ be the nonzero constituent codes of a $\ell$-QC code $\mathcal{C}\subset\mathbb{F}_{q}^{m\ell}$. By [cem, Lemma 4.1], the trace-formula described in (13) may be taken from an specific finite extension field $\mathbb{F}_{q^{w}}$, where $\mathbb{F}_{q}(\alpha^{u_{a_{i}}})\subset\mathbb{F}_{q^{w}}$, for all $1\leq i\leq h$, given the existence of elements $w_{i}\in\mathbb{F}_{q}(\alpha^{u_{a_{i}}})$ so that $\operatorname{Tr}_{\mathbb{F}_{q^{w}}/\mathbb{F}_{q}(\alpha^{u_{a_{i}}})}(w_{i})=1$. In this case, looking $\mathcal{C}$ as a collection of arrays, the columns of any codeword $\textbf{c}\in\mathcal{C}$ are codewords of a $m$-length cyclic code $D:=D_{1,2,..,h}$, whose dual’s basic zero set is
$BZ(D^{\perp})=\{\alpha^{-u_{a_{i}}}\mid i=1,2,...,h\}.$
In [BastosAlvarezFloresSalerno], the authors called $D$ the cyclic code associated to $\mathcal{C}$. Let $I=\{i_{1},i_{2},...,i_{t}\}$ be a nonempty subset of $\{1,2,...,h\}$ where
$1\leq i_{1}<i_{2}<\cdots<i_{t}\leq h.$
Let $D_{I}=D_{i_{1},...,i_{t}}\subset D$ be the cyclic subcode of $D$ whose dual’s basic zero set is
(22) $BZ(D_{I}^{\perp})=\{\alpha^{-u_{a_{i}}}\mid i\in I\}.$
Define the following quantity:
(23) $R_{I}(\mathcal{C})$ $=R_{i_{1},...,i_{t}}(\mathcal{C})$
(24) $:=\begin{cases}d(C_{i_{1}})d(D_{i_{1}})&\text{if $t=1$}\\ \\ \left(d(C_{i_{1}})-d(C_{i_{2}})\right)d(D_{i_{1}})+\left(d(C_{i_{2}})-d(C_{i_{3}})\right)d(D_{i_{1},i_{2}})+\\ \hskip 128.0374pt\vdots&\text{if $t\geq 2$}\\ +\left(d(C_{i_{t-1}})-d(C_{i_{t}})\right)d(D_{i_{1},i_{2},...i_{t-1}})+d(C_{i_{t}})d(D_{i_{1},i_{2},...,i_{t}}).\end{cases}$
The next result uses the $R_{I}$ to provide a lower bound to the minimum distance of a $\ell$-QC code $\mathcal{C}$ based on the minimum distances of its constituent codes and the cyclic code (and its cyclic subcodes) associated to $\mathcal{C}$. This lower bound is usually referred as GO-bound where relevant details can be checked in [cem]. First, we assume that the minimum distances of the nonzero constituent codes $C_{1},C_{2},...,C_{h}$ of $\mathcal{C}$ satisfy the following condition
(25) $d(C_{1})\geq d(C_{2})\geq...\geq d(C_{h}).$
Next, we use the slightly modified version of the original theorem which will be useful in determining the minimum distances of self-orthogonal (resp. dual containing or self dual) QC codes obtained from our construction in Section 3.
Theorem 2.9 (Güneri-Özbudak bound).
[cem, Theorem 4.8] Let $\mathcal{C}$ be a QC code as in (10) with the assumption (25). Then we have
(26) $d(\mathcal{C})\geq d_{\text{GO}}(\mathcal{C}):=\min\{R_{h}(\mathcal{C}),R_{h-1,h}(\mathcal{C}),R_{h-2,h-1,h}(\mathcal{C}),...,R_{1,2,...,h}(\mathcal{C})\}.$
Remark 2.
  1. (i)
    If the number of nonzero constituent codes is two, the GO-bound (26) becomes
    $d(\mathcal{C})\geq d_{\text{GO}}(\mathcal{C}):=\min\{d(C_{2})d(D_{2}),\left(d(C_{1})-d(C_{2})\right)d(D_{1})+d(C_{2})d(D)\},$
    where $D=D_{1,2}$.
  2. (i)
    If the number of nonzero constituent codes is three, it becomes
    $d(\mathcal{C})\geq d_{\text{GO}}(\mathcal{C}):=\min\{R_{3},R_{2,3},R_{1,2,3}\},$
    where
    $R_{3}(\mathcal{C})$ $=$ $d(C_{3})d(D_{3})$
    $R_{2,3}(\mathcal{C})$ $=$ $(d(C_{2})-d(C_{3}))d(D_{2})+d(C_{3})d(D_{2,3})$
    $R_{1,2,3}(\mathcal{C})$ $=$ $(d(C_{1})-d(C_{2}))d(D_{1})+(d(C_{2})-d(C_{3}))d(D_{1,3})+d(C_{3})d(D),$
    and $D=D_{1,2,3}$.
2.4. Constructions of Quantum Error-Correcting Codes from Classical Ones
From [Ketkar-Klappenecker-Kumar-Sarvepalli2006], the use of classical codes has became one of the standard methods for constructing quantum error-correcting codes. A quantum error-correcting code (QECC) $Q$ is a $K$-dimensional subspace of $(\mathbb{C}^{q})^{\otimes n}$. If $Q$ has minimum distance $d$, then we say that $Q$ is an $n,K,d]]_{q}$ code. If $K=q^{k}$, we write $n,k,d]]_{q}$. The length $n$, the dimension $K$, and minimum distance $d$ are the parameters of $Q$.
If $Q$ is a quantum code with minimum distance $d$, $Q$ can detect up to d1d-1 quantum errors, and can correct up to $\lfloor\frac{d-1}{1}\rfloor$ quantum errors. The celebrated Singleton bound [LaGuardia2021, Lemma 3.5.1] says that the parameters of an $n,K,d]]_{q}$ quantum code satisfy
(27) $\log_{q}(K)\leq n-2d+2$
or equivalently $k+2d\leq n+2$. If a quantum code $Q$ attains the quantum Singleton bound, $Q$ is called a quantum maximum distance separable (quantum MDS) code.
This paper focuses on the construction and existence of quantum stabilizer codes, where a stabilizer (quantum) code $Q\neq\{\textbf{0}\}$ is the common eigenspace of a commutative subgroup of the error group generated by a nice basis of $(\mathbb{C}^{q})^{\otimes n}$. The code $Q$ is said to be pure to d, or simply pure, if and only if its stabilizer group does not contain non-scalar matrices of weight less than its minimum distance $d$. (See Definitions 3.5.7 and 3.5.8 in [LaGuardia2021].) Moreover, a code $Q$ is said to be impure if and only if there are non-identity stabilizer elements of weight less than the minimum distance. In 1996, the following construction was introduced by Calderbank and Shor [Calderbank-Shor1996], and Steane [Steane1996], and is notably the most direct link between classical and quantum coding theory.
Theorem 2.10 (CSS construction [Ketkar-Klappenecker-Kumar-Sarvepalli2006]).
Let $C_{1}$ and $C_{2}$ denoted two classical codes with parameters $n,k_{1},d_{1}]_{q}$ and $n,k_{2},d_{2}]_{q}$ such that $C_{2}^{\perp}\subseteq C_{1}$. Then there exists $n,k_{1}+k_{2}-n,d]]_{q}$ stabilizer code with minimum distance $d=\min\{\operatorname{wt}(\mathbf{c})\mid\mathbf{c}\in(C_{1}\backslash C_{2}^{\perp})\cup(C_{2}\backslash C_{1}^{\perp})\}$ that is pure to $\min\{d_{1},d_{2}\}$.
In the special case when $C_{2}=C_{1}$, we have that $C_{1}$ is a dual-containing code. This yields
Corollary 2.11.
[Aly-Klappenecker-Sarvepalli2007, Ketkar-Klappenecker-Kumar-Sarvepalli2006] Let $C$ be a classical $[n,k,d]$ code over $\mathbb{F}_{q}$. If $C$ is dual-containing, then there exists a quantum stabilizer code $Q$ with parameters $n,2k-n,\geq d]]_{q}$ that is pure to $d$. Moreover, if the minimum distance of $C^{\perp}$ exceeds $d$, then the quantum code $Q$ has minimum distance $d$.
In recent years, researchers in coding theory have worked diligently to construct good (e.g., quantum MDS) quantum codes from classical Euclidean, Hermitian, or symplectic self-orthogonal, dual-containing, or self-dual codes [Abdukhalikov-Bag-Panario2023, Ball21, Cao24, DinhLeNguyenTansuchat21, FangFu19, GalindoHernando, galindo, GaoSunYanZhao25, JinXing13, LiuLiu21, ShiYueChang18, TianZhu19, WanZhengZhu24]. Moreover, the following results (which will be used in examples in Section 4) present new quantum code constructions utilizing the lengthening, shortening, and dimension reduction of a given stabilizer quantum code.
Lemma 2.12.
[Ketkar-Klappenecker-Kumar-Sarvepalli2006, Lemmas 69, 70, 71, and Corollary 73]
  1. (1)
    If an $n,k,d]]_{q}$ stabilizer code exists for $k>0$, then there exists an impure $n+1,k,d]]_{q}$ stabilizer code.
  2. (2)
    If a pure $n,k,d]]_{q}$ stabilizer code exists with $n\geq 2$ and $d\geq 2$, then there exists a $n-1,k+1,d-1]]_{q}$ pure stabilizer code.
  3. (3)
    If a (pure) $n,k,d]]_{q}$ stabilizer code exists, with $k\geq 2$ $(k\geq 1)$, then there exists an $n,k-1,d^{*}]]_{q}$ stabilizer code (pure to $d$) such that $d^{*}\geq d$.
  4. (4)
    Suppose that an $n,q^{k},d]]_{q}$ and an $n^{\prime},q^{k^{\prime}},d^{\prime}]]$ stabilizer code exist. Then there exists an $n+n^{\prime},q^{k+k^{\prime}},\min\{d,d^{\prime}\}]]_{q}$ stabilizer code.
3. A Construction of ESO/EDC/ESD QC codes from their constituent codes
In this section, we discuss the main results of this paper where we construct infinite families $\{\mathcal{C}_{u}\}_{u\geq 1}$ of QC codes that are self-orthogonal (or self-dual, or dual-containing) with respect to the Euclidean inner product. Using the notion of Galois closed codes, we also propose a construction of QC codes that are self-orthogonal (or self-dual, or dual-containing) with respect to the Hermitian inner product. Such constructions of QC codes will be utilized to show the existence of QECCs with good parameters in Section 4.
3.1. Constructions of Infinite Families of QC Codes
We propose a recursive construction of infinite families $\{\mathcal{C}_{u}\}_{u\geq 1}$ of ESO QC codes using its constituent codes. To provide a careful description of our construction, we first describe a general case of the construction and then specify certain conditions on the constituent codes to explicitly describe the code dimension and lower bound of its minimum distance.
Let $x^{m}-1$ factor into $2t+s$ irreducible factors, namely
(28) $x^{m}-1=\delta g_{1}g_{1}^{*}\cdots g_{t}g_{t}^{*}f_{1}\cdots f_{s-1}f_{s},$
where $\delta\in\mathbb{F}_{q}$ is nonzero, $g_{1},g_{1}^{*},...,g_{t},g_{t}^{*}$ are the factors grouped in pairs with their reciprocal polynomials for each $1\leq j\leq t$, and $f_{1},...,f_{s}$ are self-reciprocal polynomials for $1\leq i\leq s$, where $f_{s}=x-1$. Then,
(29) $R^{\ell}_{q,m}\cong\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{1}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{1}^{*}\rangle}\right)^{\ell}\oplus\cdots\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{t}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{t}^{*}\rangle}\right)^{\ell}\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle f_{1}\rangle}\right)^{\ell}\oplus\cdots\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle f_{s-1}\rangle}\right)^{\ell}\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{\ell}.$
3.1.1. Recursive Construction:
Let $\mathcal{C}_{1}\subseteq\mathbb{F}_{q}^{m\ell}$ be an $\ell$-QC whose first $2t$ constituent codes are direct sums of arbitrary codes and their duals, i.e.,
(30) $\bigoplus_{j=1}^{t}\left(C^{\prime}_{1j}\oplus{C^{\prime}}_{1j}^{\perp_{E}}\right)\subset\bigoplus_{j=1}^{t}\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}\right)^{\ell}$
and whose last $s$ constituent codes are
(31) $\left(\bigoplus_{i=1}^{s-1}C_{1i}\right)\oplus C_{1s}\subset\left(\bigoplus_{i=1}^{s-1}\left(\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\right)^{\ell}\right)\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{\ell}$
where the codes $C_{1i}$ are HSO codes for $1\leq i\leq s-1$ and $C_{1s}$ is ESO (see Remark 1). Thus,
(32) $\mathcal{C}_{1}\cong\bigoplus_{j=1}^{t}\left(C^{\prime}_{1j}\oplus{C^{\prime}}_{1j}^{\perp_{E}}\right)\oplus\left(\bigoplus_{i=1}^{s-1}C_{1i}\right)\oplus C_{1s}\subset R_{q,m}^{\ell}.$
We also assume that the minimum distances of these constituent codes satisfy the inequalities in (25), i.e.,
(33) $d(C_{11}^{\prime})\geq d((C_{11}^{\prime})^{\perp_{E}})\geq...\geq d(C_{1t}^{\prime})\geq d((C_{1t}^{\prime})^{\perp_{E}})\geq d(C_{11})\geq...\geq d(C_{1,s-1})\geq d(C_{1s}).$
As the $C_{1i}$ are HSO for each $1\leq i\leq s-1$ and $C_{1s}$ is ESO, Theorem 2.5 tells us that $\mathcal{C}_{1}$ is an ESO $\ell$-QC code of length $m\ell$. Now using $\mathcal{C}_{1}$, we construct another QC code $\mathcal{C}_{2}$ as follows: Let
(34) $R^{m\ell}_{q,m}\cong\bigoplus_{j=1}^{t}\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}\right)^{m\ell}\oplus\left(\bigoplus_{i=1}^{s-1}\left(\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\right)^{m\ell}\right)\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{m\ell}.$
Let $\mathcal{C}_{2}\subseteq\mathbb{F}_{q}^{m^{2}\ell}$ be the $\ell$-QC code whose first $2t$ constituent codes are the following direct sum of arbitrary codes and their duals,
(35) $\bigoplus_{j=1}^{t}\left(C^{\prime}_{2j}\oplus{C^{\prime}}_{2j}^{\perp_{E}}\right)\subset\bigoplus_{j=1}^{t}\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}\right)^{m\ell},$
and whose last $s$ constituent codes are
(36) $\left(\bigoplus_{i=1}^{s-1}C_{2i}\right)\oplus\mathcal{C}_{1}\subset\left(\bigoplus_{i=1}^{s-1}\left(\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\right)^{m\ell}\right)\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{m\ell}$
where the $C_{2i}$ are HSO codes for $1\leq i\leq s-1$ and $\mathcal{C}_{1}\subset\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{m\ell}\cong\mathbb{F}_{q}^{m\ell}$ is an ESO constituent code of $\mathcal{C}_{2}$. Thus,
(37) $\mathcal{C}_{2}\cong\bigoplus_{j=1}^{t}\left(C^{\prime}_{2j}\oplus{C^{\prime}}_{2j}^{\perp_{E}}\right)\oplus\left(\bigoplus_{i=1}^{s-1}\;C_{2i}\right)\oplus\mathcal{C}_{1}$
Again by Theorem 2.5, $\mathcal{C}_{2}$ is an ESO $\ell$-QC code of length $m^{2}\ell$. To extend this recursive construction, we have that any $u\geq 2$,
(38) $R^{m^{u-1}\ell}_{q,m}\cong\bigoplus_{j=1}^{t}\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}\right)^{m^{u-1}\ell}\oplus\left(\bigoplus_{i=1}^{s-1}\left(\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\right)^{m^{u-1}\ell}\right)\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{m^{u-1}\ell}.$
and define an $\ell$-QC code $\mathcal{C}_{u}$ in $\mathbb{F}_{q}^{m^{u}\ell}$ as
(39) $\mathcal{C}_{u}\cong\left(C^{\prime}_{u1}\oplus{C^{\prime}}_{u1}^{\perp}\right)\oplus\left(C^{\prime}_{u2}\oplus{C^{\prime}}_{u2}^{\perp}\right)\oplus\cdots\oplus\left(C^{\prime}_{u,t}\oplus{C^{\prime}}_{u,t}^{\perp}\right)\oplus C_{u1}\oplus\cdots\oplus C_{u,s-1}\oplus\mathcal{C}_{u-1},$
where for each $u$, the first $2t$ arbitrary constituent codes are
(40) $\bigoplus_{j=1}^{t}\left(C^{\prime}_{uj}\oplus{C^{\prime}}_{uj}^{\perp_{E}}\right)\subset\bigoplus_{j=1}^{t}\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}\right)^{m^{u-1}\ell}$
and its last $s$ constituent codes are
(41) $\left(\bigoplus_{i=1}^{s-1}\;C_{ui}\right)\oplus\mathcal{C}_{u-1}\subset\left(\bigoplus_{i=1}^{s-1}\left(\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\right)^{m^{u-1}\ell}\right)\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{m^{u-1}\ell},$
where the $C_{ui}$ are HSO codes for all $1\leq i\leq s-1$ and $\mathcal{C}_{u-1}\subset\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{m^{u-1}\ell}\cong\mathbb{F}_{q}^{m^{u-1}\ell}$ is an ESO constituent code of $\mathcal{C}_{u}$.
In the result, we determine a formula for the dimension of $\mathcal{C}_{u}$ in our recursive construction. To compute the dimension in Lemma 3.1, we focus only on the special case of this construction where the HSO constituent codes of $\mathcal{C}_{u}$ are $m^{u-1}$-copies of the HSO constituent codes of $\mathcal{C}_{1}$, namely, $C_{ui}=m^{u-1}C_{1i}$, for all $1\leq i\leq s-1$. We invoke this condition to have some control on the dimension of each $\mathcal{C}_{u}$. We note that since each $C_{1i}$ are HSO, each of the $C_{ui}$ are also HSO by a generalization of Part (i) of Theorem 2.8. Moreover, if each $C_{1i}$ has parameters $\ell,k_{1i},d_{1i}]$, then the respective parameters for $C_{ui}$ are $m^{u-1}\ell,k_{1i},m^{u-1}d_{1i}]$ for all $1\leq i\leq s-1$.
Lemma 3.1.
Let
$R_{q,m}^{\ell}\cong\bigoplus_{j=1}^{t}\left(\dfrac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\dfrac{\mathbb{F}_{q}[x]}{\langle g^{*}_{j}\rangle}\right)^{\ell}\oplus\bigg(\bigoplus_{i=1}^{s}\dfrac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\bigg)^{\ell},$
where $(g_{j},g_{j}^{*})$ are pairs of reciprocal irreducible factors, for $1\leq j\leq t$, and $f_{i}$ are self-reciprocal irreducible factors, for $1\leq i\leq s$, with $f_{s}=x-1$. Then, for $u\geq 1$, there exists a $m^{u}\ell$-length QC code $\mathcal{C}_{u}$ with constituent codes (40) and (41), whose its dimension is
$\dim(\mathcal{C}_{u})=\ell\left(\frac{m^{u}-1}{m-1}\right)\left(\sum_{j=1}^{t}\deg(g_{j})\right)+u\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}\right)+k_{1s}.$
where $k_{1i}=\dim(C_{1i})$ for all $1\leq i\leq s$.
Proof.
Let $\mathcal{C}_{1}$ be the QC code with decomposition on constituent codes as
(42) $\mathcal{C}_{1}\cong\left(\bigoplus_{j=1}^{t}\left({C_{1j}^{{}^{\prime}}}\oplus{C_{1j}^{{}^{\prime}}}^{\perp_{E}}\right)\right)\oplus\left(\bigoplus_{i=1}^{s}C_{1i}\right).$
Using Lemma 2.1, the fact that $\deg(g_{j}(x))=\deg(g_{j}^{*}(x))$ and $\dim({C^{\prime}}_{1j}^{\perp_{E}})=\ell-k^{\prime}_{1j}$ for all $1\leq j\leq t$, the dimension of $\mathcal{C}_{1}$ is
(43) $\dim(\mathcal{C}_{1})$ $=$ $\left(\sum_{j=1}^{t}\deg(g_{j})k_{1j}^{\prime}+\deg(g_{j}^{*})(\ell-k_{1j}^{\prime})\right)+\left(\sum_{i=1}^{s}\deg(f_{i})k_{1i}\right)$
$=$ $\ell\sum_{j=1}^{t}\deg(g_{j})+\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}\right)+k_{1s}.$
For any $u\geq 2$, let
(44) $\mathcal{C}_{u}=\bigoplus_{j=1}^{t}\left(C_{uj}^{\prime}\oplus{C^{\prime}}_{uj}^{\perp_{E}}\right)\oplus\left(\bigoplus_{i=1}^{s-1}C_{ui}\right)\oplus\mathcal{C}_{u-1},$
where $C_{ui}$ is $m^{u-1}$-copies of $C_{1i}$, namely, $C_{ui}=m^{u-1}C_{ui}$, for all $1\leq i\leq s$. Given its recursive argument, $\dim(\mathcal{C}_{u})$ is computed via induction. For $u=1$, it is done. Now assume that the argument is valid for $u$, i.e., there exists a QC code $\mathcal{C}_{u}\subset\mathbb{F}_{q}^{m^{u}\ell}$ with the following dimension
(45) $\dim(\mathcal{C}_{u})=\ell\left(\frac{m^{u}-1}{m-1}\right)\left(\sum_{j=1}^{t}\deg(g_{j})\right)+u\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}\right)+k_{1s}$
where $k_{1i}=\dim(C_{1i})=\dim(m^{u-1}C_{1i})=\dim(C_{ui})$ for all $1\leq i\leq s-1$. Now let $\mathcal{C}_{u+1}\subset\mathbb{F}_{q}^{m^{u+1}\ell}$ be the QC code with the following $m^{u}\ell$-length constituent codes
(46) $\mathcal{C}_{u+1}\cong\left(\bigoplus_{j=1}^{t}\left(C_{u+1,j}^{\prime}\oplus C_{u+1,j}^{{}^{\prime}\perp_{E}}\right)\right)\oplus\left(\bigoplus_{i=1}^{s-1}{C_{u+1,i}}\right)\oplus\mathcal{C}_{u}$
where each $C_{u+1,i}=m^{u}C_{1i}$ are also HSO for all $1\leq i\leq s-1$. (See Theorem 2.8.) Using equation  (45), we have
$\dim(\mathcal{C}_{u+1})$ $=\sum_{j=1}^{t}k^{\prime}_{u+1,j}\deg(g_{j})+\sum_{j=1}^{t}\deg(g_{j}^{*}){k^{\prime}}_{u+1,j}^{\perp_{E}}+\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}\right)+\dim(\mathcal{C}_{u})$
$=\sum_{j=1}^{t}\left[\deg(g_{j})k^{\prime}_{u+1,j}+\deg(g_{j})\left(m^{u}\ell-k^{\prime}_{u+1,j}\right)\right]+\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}\right)+\dim(\mathcal{C}_{u})$
$=m^{u}\ell\sum_{j=1}^{t}\deg(g_{j})+\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}\right)+\dim(\mathcal{C}_{u})$
$=m^{u}\ell\sum_{j=1}^{t}\deg(g_{j})+\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}\right)+\ell\left(\frac{m^{u}-1}{m-1}\right)\left(\sum_{j=1}^{t}\deg(g_{j})\right)+u\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}\right)+k_{1s}$
$=\ell\left(m^{u}+\frac{m^{u}-1}{m-1}\right)\left(\sum_{j=1}^{t}\deg(g_{j})\right)+(u+1)\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}\right)+k_{1s}$
$=\ell\left(\frac{m^{u+1}-1}{m-1}\right)\left(\sum_{j=1}^{t}\deg(g_{j})\right)+(u+1)\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}\right)+k_{1s}$
as desired. ∎
In the following lemma, we utilize Theorem 2.9 to find a lower bound for the minimum distance $d(\mathcal{C}_{u})$ for each $u\geq 2$. To obtain our lower bound, we include the condition that for each $u\geq 2$, the last HSO constituent codes $C_{ui}$ in (41) satisfy
(47) $d(C_{ui})\geq md(C_{u-1,i}),$
for all $1\leq i\leq s-1$. Note that repeated application of (47) yields $d(C_{ui})\geq m^{u-1}d(C_{1i})$ for each ii. This condition can be satisfied, for instance, by setting $C_{ui}=mC_{u-1,i}$. However, unlike in Lemma 3.1, taking copies of constituent codes is not required to obtain our result.
Lemma 3.2.
Let
$R_{q,m}^{\ell}\cong\bigoplus_{j=1}^{t}\left(\dfrac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\dfrac{\mathbb{F}_{q}[x]}{\langle g^{*}_{j}\rangle}\right)^{\ell}\oplus\bigg(\bigoplus_{i=1}^{s}\dfrac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\bigg)^{\ell},$
where $(g_{j},g_{j}^{*})$ are pairs of reciprocal irreducible factors, for $1\leq j\leq t$, and $f_{i}$ are self-reciprocal irreducible factors, for $1\leq i\leq s$ with $f_{s}=x-1$. Given the constituent codes provided in (40) and (41), then, for $u\geq 1$, there exists an $m^{u}\ell$-length QC code $\mathcal{C}_{u}$ with $d(\mathcal{C}_{u})\geq m^{u-1}d_{1s}$, where $d_{1s}=d(C_{1s})$.
Proof.
We again use the induction argument along with the associated cyclic codes per (22). We first note that $D_{2t+s}$ is the $m$-length repetition code, so $d(D_{2t+s})=m$. For $u=1$, equation (24) and the fact that all nonzero constituent codes of $\mathcal{C}_{1}$ satisfy (25) yields
$R_{2t+s}(\mathcal{C}_{1})$ $=$ $d(D_{2t+s})d(C_{1s})=md_{1s}\geq d_{1s}$
$R_{2t+s-1,2t+s}(\mathcal{C}_{1})$ $=$ $(d(C_{1,s-1})-d(C_{1s}))d(D_{2t+s-1})+d(C_{1s})d(D_{2t+s-1,2t+s})\geq d(C_{1s})=d_{1s}$
$\vdots$
$R_{1,2,...,2t+s}(\mathcal{C}_{1})$ $=$ $(d(C_{11}^{\prime})-d(C_{11}^{{}^{\prime}\perp_{E}}))d(D_{1})+(d(C_{11}^{{}^{\prime}\perp_{E}})-d(C_{12}^{\prime}))d(D_{1,2})+\ldots+d(C_{1s})d(D_{1,2,...,2t+s})$
$\geq$ $d(C_{1s})=d_{1s}.$
By Theorem 2.9, this shows $d(\mathcal{C}_{1})\geq d_{GO}(\mathcal{C}_{1})=\min\{R_{2t+s}(\mathcal{C}_{1}),R_{2t+s-1,2t+s}(\mathcal{C}_{1}),\ldots,R_{1,2,...,2t+s}(\mathcal{C}_{1})\}\geq d_{1s}$.
Now assume that for $u\in\mathbb{N}$, $d(\mathcal{C}_{u})\geq m^{u-1}d_{1s}$. Let $\mathcal{C}_{u+1}$ be the $m^{u+1}\ell$-length QC code constructed from $\mathcal{C}_{u}$, with constituent-code decomposition (46) and whose last constituent code is $\mathcal{C}_{u}$. Additionally, the minimum distances of the constituent codes of $\mathcal{C}_{u+1}$ over $\dfrac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}$ must satisfy equation (47), namely
$d(C_{u+1,i})\geq md(C_{u,i}),\quad\text{for all}\quad 1\leq i\leq s-1.$
At this point, we do not have information on how $d(\mathcal{C}_{u})$ compares with the minimum distances of the other constituent codes. As the minimum distances in any decomposition of a QC code should satisfy the descending chain in  (25) to compute the GO-bound, we have the following $2t+s$ cases:
[…]
where $F_{i}:=\dfrac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}$, for each $1\leq i\leq s-1$, and $G^{\prime}_{j}:=\dfrac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle},$ $G^{\prime\prime}_{j}:=\dfrac{\mathbb{F}_{q}[x]}{\langle g^{*}_{j}\rangle}$ for each $1\leq j\leq t$. Notice that in each case, the respective QC code $\mathcal{C}_{u+1}$ is the same regardless of the ordering of its constituent codes (see (13)).
For the first 2t+s12t+s-1 cases, we notice that all terms in the $R_{I}(\mathcal{C}_{u+1})$ in (24) are positive, and in particular, the last terms of $R_{2t+s}(\mathcal{C}_{u+1}),$ $R_{2t+s-1,2t+s}(\mathcal{C}_{u+1}),...,R_{1,2,...,2t+s}(\mathcal{C}_{u+1})$ are multiples of $d(C_{u+1,s-1})$. As, $d(C_{u+1,s-1})\geq m^{u}d(C_{1,s-1})\geq m^{u}d(C_{1,s})$. Thus for these cases, $d(\mathcal{C}_{u})\geq d(C_{u+1,s-1})\geq m^{u}d_{1s}$.
For $(2t+s)$-th case, we have
$R_{2t+s}(\mathcal{C}_{u+1})$ $=$ $d(\mathcal{C}_{u})d(D_{2t+s})\geq m^{u-1}d_{1s}m=m^{u}d_{1s},\mbox{ from induction hypothesis,}$
$R_{2t+s-1,2t+s}(\mathcal{C}_{u+1})$ $=$ $(d(C_{u+1,s-1})-d(\mathcal{C}_{u}))d(D_{2t+s-1})+d(\mathcal{C}_{u})d(D_{2t+s-1,2t+s})$
$=$ $d(C_{u+1,s-1})d(D_{2t+s-1})-d(\mathcal{C}_{u})\underbrace{(d(D_{2t+s-1})-d(D_{2t+s-1,2t+s}))}_{>0}$
$\geq$ $d(C_{u+1,s-1})d(D_{2t+s-1})-d(C_{u+1,s-1})(d(D_{2t+s-1})-d(D_{2t+s-1,2t+s}))$
$=$ $d(C_{u+1,s-1})d(D_{2t+s-1,2t+s})$
$\geq$ $m^{u}d_{1s},$
$\vdots$
$R_{1,2,...,2t+s}(\mathcal{C}_{u+1})$ $=$ $(d(C_{u1}^{\prime})-d(C_{u+1,1}^{{}^{\prime}\perp}))d(D_{1})+\ldots+(d(C_{u+1,s-1})-d(\mathcal{C}_{u}))d(D_{1,2,...,2t+s-1})$
$+d(\mathcal{C}_{u})d(D_{1,2,...,2t+s})$
$=$ $\underbrace{(d(C_{u+1,1}^{\prime})-d(C_{u+1,1}^{{}^{\prime}\perp}))d(D_{1})+\ldots}_{\geq 0}+d(C_{u+1,s-1})d(D_{1,2,...,2t+s-1})$
- $d(\mathcal{C}_{u})\underbrace{(d(D_{1,2,...,2t+s-1})-d(D_{1,2,...,2t+s}))}_{>0}$
$\geq$ $d(C_{u+1,s-1})d(D_{1,2,...,2t+s-1})-d(C_{u+1,s-1})(d(D_{1,2,...,2t+s-1})-d(D_{1,2,...,2t+s}))$
$=$ $d(C_{u+1,s-1})d(D_{1,2,...,2t+s})$
$\geq$ $m^{u}d_{1s}$
where $d(D_{1,2,...,2t+s-1})-d(D_{1,2,...,2t+s})>0$ from Equation (4.19) in [cem]. Therefore by Theorem 2.9,
$d(\mathcal{C}_{u+1})\geq d_{GO}(\mathcal{C}_{u+1})=\min\{R_{2t+s}(\mathcal{C}_{u+1}),R_{2t+s-1,2t+s}(\mathcal{C}_{u+1}),\ldots,R_{1,2,...,2t+s}(\mathcal{C}_{u+1})\}\geq m^{u}d_{1s},$
as desired. ∎
The following theorem provides a summary of our construction of a QC code in $\mathbb{F}_{q}^{m^{u}\ell}$ that is self-orthogonal with respect to the Euclidean inner product.
Theorem $3.$
Let
$R_{q,m}^{\ell}\cong\bigoplus_{j=1}^{t}\left(\dfrac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\dfrac{\mathbb{F}_{q}[x]}{\langle g^{*}_{j}\rangle}\right)^{\ell}\oplus\bigg(\bigoplus_{i=1}^{s}\dfrac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\bigg)^{\ell},$
where $(g_{j},g_{j}^{*})$ are pairs of reciprocal irreducible factors, for $1\leq j\leq t$, and $f_{i}$ are self-reciprocal irreducible factors, for $1\leq 1\leq s$, with $f_{s}=x-1$. Let $\mathcal{C}_{u}$ be the $m^{u}\ell$-length QC code whose constituent codes are in (40) and (41), where $C_{ui}=m^{u-1}C_{1i}$, for all $1\leq i\leq s-1$. Then, we have an infinite family $\{\mathcal{C}_{u}\}_{u\geq 1}$ of ESO $q$-ary $\ell$-QC codes with parameters
(48) $\left[m^{u}\ell,\ell\left(\frac{m^{u}-1}{m-1}\right)\left(\sum_{j=1}^{t}\deg(g_{j})\right)+u\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}\right)+k_{1s},\geq m^{u-1}d_{1s}\right].$
Proof.
The parameters and duality of $\mathcal{C}_{u}$ comes from the recursive construction, and Lemmas 3.13.2. ∎
The next result is a particular case of Theorem 3.3, namely, where we remove some restrictions on the constituent codes which allow us also to construct infinite families of not only ESO QC codes, but also ones that are ESD and EDC. From now on, we assume $m$ is not even, when $q$ is odd (See Remark 1). In these special cases, the unique self-reciprocal irreducible factor of $x^{m}-1$ is $x-1\in\mathbb{F}_{q}[x]$. Hence, we no longer need to use copies of the constituent codes of $\mathcal{C}_{u}$ to obtain the constituent codes of $\mathcal{C}_{u+1}$. Consequently, we have no dimensional restriction on these codes, which allows for them to be ESD/EDC.
Corollary 3.4.
Let
$R_{q,m}^{\ell}\cong\left(\bigoplus_{i=1}^{t}\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{i}\rangle}\ \oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{i}^{*}\rangle}\right)^{\ell}\right)\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{\ell},$
where $(g_{i},g_{i}^{*})$ are pairs of distinct reciprocal polynomials, for all $1\leq i\leq t$. Let $\mathcal{C}_{1}\subset\mathbb{F}_{q}^{m\ell}$ be the ESO (resp. EDC or ESD) $\ell$-QC code with constituent codes $C_{1i}\oplus C_{1i}^{\perp}\subset\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{i}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{i}^{*}\rangle}\right)^{\ell}$ and a ESO (resp. EDC or ESD) $\ell,k_{1,t+1},d_{1,t+1}]_{q}$ code $C_{1,t+1}\subset\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{\ell}$, where $d(C_{11})\geq d^{\perp}(C_{11})\geq...\geq d(C_{1t})\geq d^{\perp}(C_{1t})\geq d(C_{1,t+1}).$ Then there exists an infinite family $\{\mathcal{C}_{u}\}_{u\geq 1}$ of ESO (resp. EDC or ESD) $q$-ary QC codes with parameters
(49) $\left[m^{u}\ell,\frac{\ell}{2}\left(m^{u}-1\right)+k_{1,t+1},\geq m^{u-1}d_{1,t+1}\right].$
Proof.
This is a special case of Theorem 3.3 where $s=1$, and the unique self-reciprocal irreducible factor of $x^{m}-1$ is x1x-1. Hence,
(50) $\sum_{j=1}^{t}\deg(g_{j})=\frac{1}{2}\left[(m-\left(\sum_{i=1}^{s-1}\deg(f_{i})\right)-1\right]\overset{s=1}{=}\frac{1}{2}(m-1),$
namely, ${\dim\mathcal{C}_{u}=\ell\left(\frac{m^{u}-1}{m-1}\right)\left(\sum_{j=1}^{t}\deg(g_{j})\right)+k_{1,t+1}=\frac{\ell}{2}\left(m^{u}-1\right)+k_{1,t+1}}$.
The duality of all constituent codes may be taken under the Euclidean inner product (see Remark 1 - item (1)) and it is defined exclusively by the $q$-ary constituent code $C_{1,t+1}$. Indeed, if $C_{1,t+1}$ is ESO (resp. EDC or ESD), then $\mathcal{C}_{1}$ also is. By the recursive construction, for any $u\geq 1$. the duality of $\mathcal{C}_{1}$ defines the duality of $\mathcal{C}_{2}$, and so on. This avoids relying on copies of constituent codes, which would otherwise restrict the construction to only ESO codes because of the dimensional limitations of self-dual and dual-containing codes. ∎
Remark 3.
We note that if we want $x^{m}-1$ to have three distinct irreducible factors in $\mathbb{F}_{q}[x]$, of which x1x-1 is one, we need that $m$ is prime or the square of a prime, in addition to $m$ and $q$ being coprime. Otherwise, $m$ would have at least two distinct prime factors so that in the factorization
(51) $x^{m}-1=\prod_{d\mid m}\Phi_{d}(x),$
where $\Phi_{d}$ is the $d$-th cyclotomic polynomial, there would be at least four cyclotomic polynomial factors (11, $m$, and $m$’s distinct prime factors). This would result in at least four irreducible factors. If $m$ is prime,
(52) $x^{m}-1=\Phi_{1}(x)\Phi_{m}(x)=(x-1)\Phi_{m}(x),$
so we must consider the irreducible factors of $\Phi_{m}$. The polynomial $\Phi_{m}$ has $\frac{\phi(m)}{\operatorname{ord}_{m}(q)}=\frac{m-1}{\operatorname{ord}_{m}(q)}$ irreducible factors. Thus, if $x^{m}-1$ has three irreducible factors, we need $\operatorname{ord}_{m}(q)=\frac{m-1}{2}$, i.e. $q^{(m-1)/2}\equiv 1\ (\operatorname{mod}\ m)$. If $m$ is the square of a prime, say $p^{2}$ where $p$ is prime,
(53) $x^{m}-1=\Phi_{1}(x)\Phi_{p}(x)\Phi_{p^{2}}(x).$
We need that both $\Phi_{p}$ and $\Phi_{p^{2}}$ are both irreducible, i.e., $1=\frac{\phi(p)}{\operatorname{ord}_{p}(q)}=\frac{p-1}{\operatorname{ord}_{p}(q)}$ so that $\operatorname{ord}_{p}(q)=p-1$ and $1=\frac{\phi(p^{2})}{\operatorname{ord}_{p^{2}}(q)}=\frac{p(p-1)}{\operatorname{ord}_{p^{2}}(q)}$ so that ${\operatorname{ord}_{p^{2}}(q)}=p(p-1)$. The following tables provide some examples of $m$ and $q$ pairs where $x^{m}-1\in\mathbb{F}_{q}[x]$ decomposes into one pair of reciprocal polynomials and x1x-1. One caveat is that over $\mathbb{F}_{q^{2}}[x]$, i.e., when $q$ is the square of a prime, $\Phi_{p}$ is irreducible if and only if $p=2$ by Fermat’s little theorem, and so $x^{p^{2}}-1$ does not have three irreducible factors over $\mathbb{F}_{q^{2}}$.
$q$ $1\leq m\leq 100$
2 7, 17, 23, 41, 47, 71, 79, 97
3 11, 23, 37, 47, 59, 71, 83, 97
5 11, 19, 29, 41, 59, 61, 79, 89
7 3, 31, 47, 53, 59, 83
11 7, 53, 79, 83, 97
Table 1. Some values of $m$ and $q$ for which $x^{m}-1\in\mathbb{F}_{q}[x]$ decomposes into three irreducible factors over $\mathbb{F}_{q}$.
$q^{2}$ $1\leq m\leq 100$
4 3,5,7,11,13,19,23,29,37,47,53,59,61,67,71,79,83
9 5,7,11,17,19,23,29,31,43,47,53,59,71,79,83,89
25 3,7,11,17,19,23,37,43,47,53,59,73,79,83,97
49 3,5,11,13,17,23,31,41,47,59,61,67,71,79,83,89,97
121 3,7,13,17,23,29,31,41,47,59,67,71,73,79,83
Table 2. Some values of $m$ and $q$ for which $x^{m}-1\in\mathbb{F}_{q^{2}}[x]$ decomposes into three irreducible factors over $\mathbb{F}_{q^{2}}$.
Example 3.5.
Let
$R_{3,11}^{5}$ $\cong$ $\left(\frac{\mathbb{F}_{3}[x]}{\langle x^{5}+2x^{3}+x^{2}+2x+2\rangle}\right)^{5}\oplus\left(\frac{\mathbb{F}_{3}[x]}{\langle x^{5}+x^{4}+2x^{3}+x^{2}+2\rangle}\right)^{5}\oplus\left(\frac{\mathbb{F}_{3}[x]}{\langle x-1\rangle}\right)^{5}$
$=$ $\left(G_{1}^{\prime}\oplus G_{1}^{\prime\prime}\right)^{5}\oplus\left(\mathbb{F}_{3}\right)^{5},$
where $G_{1}^{\prime}=\mathbb{F}_{3}(\alpha)$. In addition, let $C_{11}^{\prime}=\langle(1,2,1,2,1),(\alpha,\alpha^{2},\alpha^{3},\alpha^{4},\alpha^{5})\rangle$ be the $5,2,4]_{3^{5}}-$code, $C_{11}^{\prime\prime}=C_{11}^{{}^{\prime}\perp_{E}}\subset(G_{1}^{\prime\prime})^{5}$ the $5,3,3]_{3^{5}}-$code, and $C_{11}=\langle(1,1,1,0,0),(1,2,0,1,0)\rangle$ the self-orthogonal $5,2,3]_{3}-$code, which are the constituent codes of the QC code $\mathcal{C}_{1}$. Thus, we may obtain the $m^{u}\ell$-length QC codes
(54) $\mathcal{C}_{u}\cong 11^{u-1}C_{11}^{\prime}\oplus 11^{u-1}C_{11}^{{}^{\prime}\perp}\oplus\mathcal{C}_{u-1},\mbox{ for }u\geq 2.$
According to Corollary 3.4, it is possible to get a infinite family $\{\mathcal{C}_{u}\}_{u\geq 1}$ of ESO QC codes with parameters $\left[5\cdot 11^{u},\frac{5\cdot 11^{u}-1}{2},\geq 3\cdot 11^{u-1}\right]_{3}$.
3.2. Constructing ESO/ESD/EDC QC Codes with Square-Root-Like Lower Bound
Self-dual codes have applications in several mathematical problems such as in the construction of designs, lattices, and quantum error-correcting codes (see Subsection 2.4). Notably, the construction of infinite families of $n_{i},k_{i},d_{i}]_{q}$ self-dual codes is itself an interesting problem. Additionally, code families where $d_{i}\geq c\sqrt{n_{i}}$ for some constant $c$ are said to satisfy the square-root-like lower bound. In the special case when $c=1$, we say that the code family satisfies the square-root lower bound. According to [infiniteselfdualfamilies], there are very few constructions of infinite families of self-dual codes whose minimum distances meet the square-root lower bound, which are depicted in that paper. Notably, the work in  [Calderbank] presents a unique construction of an infinite family of self-dual QC codes.
To the best of our knowledge, the authors of [infiniteselfdualfamilies] in 2025 presented the first infinite families of non-binary ESD (and HSD) cyclic codes satisfying the square-root-like lower bound to their minimum distances. Using results from [lingsole] and developments presented in this work (see Corollary 3.4), we construct several infinite families of ESD QC codes which satisfy the square-root like lower bound. Moreover, we also construct HSD QC codes using results in the next subsection. We first state the following helpful lemma.
Lemma 3.6.
Let $m$ and $u$ be positive integers. If $u\geq 2$, then $\sqrt{m^{u}}\leq m^{u-1}.$
Proof.
$u\geq 2\Rightarrow u+2\leq 2u\Rightarrow m^{u+2}\leq m^{2u}\Rightarrow m^{u}\leq m^{(u-1)^{2}}\Rightarrow\sqrt{m^{u}}\leq m^{u-1}$. ∎
Let $\{\mathcal{C}_{u}\}_{u\geq 1}$ be the infinite family of QC codes proposed in Corollary 3.4. As mentioned in the statement of that corollary, $\mathcal{C}_{u}$ being ESO/EDC/ESD comes from the duality of the last constituent code over $\mathbb{F}_{q}$. Based on this observation, we have the following result.
Theorem 3.7.
Let $R_{q,m}^{\ell}\cong\bigoplus_{i=1}^{t}\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{i}(x)\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{i}^{*}(x)\rangle}\right)^{\ell}\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{\ell}$, where $(g_{i}(x),g_{i}^{*}(x))$ are pairs of distinct reciprocal polynomials, for all $1\leq i\leq t$. Let $\mathcal{C}_{1}\subset\mathbb{F}_{q}^{m\ell}$ be the ESD $\ell$-QC code with constituent codes $C_{1i}\oplus C_{1i}^{\perp}\subset\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{i}(x)\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{i}^{*}(x)\rangle}\right)^{\ell}$ and a ESD code $C_{1,t+1}\subset\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{\ell}$, where
$d(C_{11})\geq d^{\perp}(C_{11})\geq...\geq d(C_{1t})\geq d^{\perp}(C_{1t})\geq d(C_{1,t+1})\geq c\sqrt{\ell}.$
Then there exists an infinite family $\{\mathcal{C}_{u}\}_{u\geq 2}$ of ESD QC codes with parameters
(55) $\left[m^{u}\ell,\ell\left(\frac{m^{u}-1}{m-1}\right)\left(\sum_{j=1}^{t}\deg(g_{j})\right)+k_{1s},\geq m^{u-1}d_{1,t+1}\right],$
where all these codes have square-root-like bound on their minimum distances.
Proof.
According to Corollary 3.4 and the fact that the last constituent code of $\mathcal{C}_{1}$ is self-dual, it follows that all $\mathcal{C}_{u}$ is self-dual as well, for $u\geq 1$. Additionally, from Lemma 3.6, for each $m^{u}\ell$-length QC code $\mathcal{C}_{u}$, with $u\geq 2$, we have
(56) $d(\mathcal{C}_{u})\geq m^{u-1}d_{1,t+1}\geq\sqrt{m^{u}}d_{1,t+1}\geq\sqrt{m^{u}}c\sqrt{\ell}=c\sqrt{m^{u}\ell},$
namely, the minimum distances in $\{\mathcal{C}_{u}\}_{u\geq 2}$ satisfy the square-root-like bound, and the result follows. ∎
Remark 4.
Any $c\leq m^{u/2-1}$ will work for square-root-like bound. If we would like a uniform $c$, then $c<\sqrt{m}$ works for all $u\geq 1$.
Example 3.8.
Let
$R_{5,11}^{6}$ $\cong$ $\left(\frac{\mathbb{F}_{5}[x]}{\langle x^{5}+2x^{4}+4x^{3}+x^{2}+x+4\rangle}\right)^{6}\oplus\left(\frac{\mathbb{F}_{5}[x]}{\langle x^{5}+4x^{4}+4x^{3}+x^{2}+3x+4\rangle}\right)^{6}\oplus\left(\frac{\mathbb{F}_{5}[x]}{\langle x-1\rangle}\right)^{6}$
$=$ $\left(G_{1}^{\prime}\oplus G_{1}^{\prime\prime}\right)^{6}\oplus\left(\mathbb{F}_{5}\right)^{6},$
where $G_{1}^{\prime}=\mathbb{F}_{5}(\alpha)$. Let $C_{11}^{\prime}$ be a $6,3,4]_{5^{5}}-$GRS code, and once $C_{11}^{{}^{\prime\prime}}=C_{11}^{{}^{\prime}\perp_{E}}$, then $C_{11}^{\prime\prime}$ is also a $6,3,4]_{5^{5}}-$GRS code. In addition, let $C_{11}=\langle(1,0,0,2,2,4),(0,1,0,2,4,2),(0,0,1,4,2,2)\rangle$ be a self-dual $6,3,4]_{5}$ code. Notice all constituent codes are MDS. Therefore, according to Theorem 3.7, it is possible to get a infinite family $\{\mathcal{C}_{u}\}_{u\geq 2}$ of ESD QC codes with parameters $\left[4\cdot 11^{u},2\cdot 11^{u},\geq 4\cdot 11^{u-1}\right]$ for $u\geq 2$, with square-root-like bound for their minimum distances.
We conclude this subsection by emphasizing that, alongside the seminal constructions of infinite families of self-dual cyclic codes satisfying the square-root-like bound in [infiniteselfdualfamilies], our work introduces a novel construction of infinite families of QC codes that also attain this square-root-like lower bound on minimum distance.
$3.$ Galois Closed QC Codes and Their Duals
Having QC codes whose Hermitian and Euclidean duals are equal is a valuable property as it allows us to move freely between the two inner products. To achieve this property, we introduce the notion of the Galois closure of a code $C$: By  [Jin2010], given a positive integer $m\geq 1$ and $c=(c_{1},c_{2},...,c_{n})\in C\subset\mathbb{F}_{q}^{n}$ (in this case, $C$ is just a subset), we define $c^{m}=(c_{1}^{m},c_{2}^{m},...,c_{n}^{m})$ and $C^{m}=\{c^{m}:c\in C\}\subset\mathbb{F}_{q}^{n}.$ If $C$ is a subspace of $\mathbb{F}_{q}^{n}$ and $m$ is a power of the characteristic of $\mathbb{F}_{q}$, i.e., $m=q^{t}$ for some $t\in\mathbb{Z}^{+}$, then $C^{m}$ is also a subspace of $\mathbb{F}_{q}^{n}$.
Let $\mathbb{F}_{q}\subset\mathbb{F}_{q^{r}}$ and $C$ be an $\mathbb{F}_{q^{r}}$-code. We call $C$ Galois closed if $C=C^{q}$. See [Bierbrauer] for further information. The Frobenius map $C\mapsto C^{q}$ respects the quasi-cyclic property of $C$ so that $C^{q}$ is also quasi-cyclic whenever $C$ is. Furthermore, if $\phi(C),\phi(C^{q})\subset\left(\frac{\mathbb{F}_{q^{2}}[x]}{\langle x^{m}-1\rangle}\right)^{\ell}$, then we can look at the constituent codes of $C$ and $C^{q}$ and assume that the irreducible factors of $x^{m}-1$ are aligned for both decompositions.
Remark 5.
For any code $C$ over $\mathbb{F}_{q^{2}}$, we observe that the Hermitian dual $C^{\perp_{H}}$ is equal to the Euclidean dual $\left(C^{q}\right){{}^{\perp_{E}}}$ of $C^{q}$. Hence, $C$ is HSO if and only if $C\subset(C^{q})^{\perp_{E}}$, i.e., $C^{q}\subset C^{\perp_{E}}$. Moreover, if $C$ is Galois closed, then $C^{\perp_{H}}=C^{\perp_{E}}$.
The following two results give a classification and construction for the Galois closed codes in $\mathbb{F}_{q^{2}}^{\ell}$, namely that they are direct sums of one-dimensional subspaces spanned by vectors with entries 0 or $v_{i}$ satisfying $v_{i}^{q-1}=\beta$ for some fixed $\beta\in\mathbb{F}_{q^{2}}$. This allows for easy construction of Galois closed codes and demonstrates that they are not rare but rather plentiful, making them indeed useful and desirable.
Proposition 3.9.
If $C\subset\mathbb{F}_{q^{2}}^{\ell}$ is Galois closed and $C=\langle\textbf{v}\rangle$ where $\textbf{v}=(v_{1},\ldots,v_{\ell})$, then $v_{i}=0$ or $v_{i}^{q-1}=\beta$ for some fixed $\beta\in\mathbb{F}_{q^{2}}$ for all $1\leq i\leq s$.
Proposition 3.10.
Let $C\subseteq\mathbb{F}_{q^{2}}^{\ell}$ be a $k$-dimensional code. Then $C$ is Galois closed if and only if $C=\bigoplus_{i}^{k}C_{i}$ where $C_{i}=C_{i}^{q}$ and $\dim(C_{i})=1$, i.e., $C$ is Galois closed if any only if can be decomposed as Galois closed codes of dimension 11.
Proposition 3.11.
Let $C\subset\mathbb{F}_{q^{2}}$ be a Galois closed code. $C$ is HSO code if and only if $C$ is ESO.
Proof.
$C$ is HSO if and only if $C^{q}\subset C^{\perp_{E}}$. Since $C^{q}=C$, the result follows. ∎
In general, $C\neq C^{q}$, but they have the same parameters. According to Remark 5, if $C$ is Galois closed, then its Hermitian and Euclidean duals are exactly the same, i.e. $C^{\perp_{H}}=C^{\perp_{E}}$. Furthermore, the following theorems shows that if $C$ is Galois closed, then its Euclidean dual is also Galois closed.
Proposition 3.12.
Let $S=\left\{(c_{1},...,c_{n}),\left(c_{1}^{q},...,c_{n}^{q}\right),...,\left(c_{1}^{q^{r-1}},...,c_{n}^{q^{r-1}}\right)\right\}\subset\mathbb{F}_{q^{r}}^{n}$, and $C=\langle S\rangle$ be the respective $n,r\geq k]-$code. Then $C$ and its Euclidean dual are Galois closed.
Proof.
Without loss of generality, let $B=\{(c_{1},...,c_{n}),(c_{1}^{q},...,c_{n}^{q}),...,(c_{1}^{q^{k-1}},...,c_{n}^{q^{k-1}})\}=\{\mathbf{c_{0}},\mathbf{c_{1}},...,\mathbf{c_{k-1}}\}$ be a basis of $C$. Given $\mathbf{c_{\gamma}}=\sum_{i=0}^{k-1}\gamma_{i}\mathbf{c_{i}}$, where $\gamma_{i}\in\mathbb{F}_{q^{r}}$ for $0\leq i\leq k-1$, we have
(57) $\mathbf{c_{\gamma}}^{q}$ $=$ $(\sum_{i=0}^{k-1}\gamma_{i}\mathbf{c_{i}})^{q}=\sum_{i=0}^{k-1}\gamma_{i}^{q}\mathbf{c_{i+1}}=\sum_{i=0}^{k-2}\gamma_{i}^{q}\mathbf{c_{i+1}}+\gamma_{k-1}^{q}\sum_{i=0}^{k-1}\beta_{i}\mathbf{c_{i}},\mbox{ where }\beta_{i}\in\mathbb{F}_{q^{r}},i=0,...,k-1;$
$=$ $\gamma_{k-1}^{q}\beta_{0}\mathbf{c_{0}}+\sum_{i=1}^{k-1}(\gamma_{i-1}^{q}+\beta_{i})\mathbf{c_{i}}\in C,$
namely, $C^{q}\subset C$. Since $C$ and $C^{q}$ have the same dimension; therefore, $C$ is Galois closed.
Finally, it is enough to prove that $C^{\perp_{E}}\subset(C^{q})^{\perp_{E}}$, since $(C^{q})^{\perp_{E}}\mbox{ and }C^{\perp_{E}}$ have the same dimension. Let $\mathbf{x}=(x_{1},...,x_{n})\in C^{\perp_{E}}$. Therefore,
(58) $\langle\mathbf{x}^{q},\mathbf{c}_{\gamma}\rangle_{E}=\sum_{i=0}^{k-1}\langle\mathbf{x}^{q},\gamma_{i}\mathbf{c_{i}}\rangle_{E}=\left(\sum_{i=0}^{k-1}\langle\mathbf{x},\gamma_{i}^{q^{r-1}}\mathbf{c_{i-1}}\rangle_{E}\right)^{q}=0$
where $\mathbf{c_{-1}}$ is written as linear combination of $\mathbf{c_{0}},...,\mathbf{c_{k-1}}$. ∎
The Trace Representation (Theorem 2.2) of $\ell$-QC codes depends on the constituent codes and $q$-cyclotomic cosets. It turns out that the the constituent codes of a QC code being Galois closed directly determines whether $\mathcal{C}\subset\mathbb{F}_{q}^{m\ell}$ is Galois closed.
Theorem 3.13.
An $\ell$-QC code $\mathcal{C}\subset\mathbb{F}_{q^{2}}^{m\ell}$ is Galois closed if and only if its respective constituent codes are Galois closed.
Proof.
Let $\mathcal{C}$ be an $\ell$-QC code with decomposition in (10). Then
(59) $\mathcal{C}^{q}\cong\Bigg(\bigoplus_{j=1}^{t}({(C^{q})}{{}^{\prime}}_{j}\oplus{(C^{q})}{{}^{\prime\prime}}_{j})\Bigg)\oplus\bigg(\bigoplus_{i=1}^{s}(C^{q})_{i}\bigg).$
As the corresponding constituent codes of $\mathcal{C}^{q}$ and $\mathcal{C}$ are aligned, we have that $(C^{q})^{\prime}_{j}=C^{\prime q}_{j}$ and $(C^{q})^{\prime\prime}_{j}=C^{\prime\prime q}_{j}$ for all $1\leq j\leq t$, and $(C^{q})_{i}=C_{i}^{q}$ for each $1\leq i\leq s$. Thus $\mathcal{C}=\mathcal{C}^{q}$ as codes if and only if $(C^{q})_{i}=C_{i}^{q}=C_{i}$ for each $1\leq i\leq s$, and $(C^{q})^{\prime}_{j}=C^{\prime q}_{j}=C^{\prime}_{j}$ and $(C^{q})^{\prime\prime}_{j}=C^{\prime\prime q}_{j}=C^{\prime\prime}_{j}$ for all $1\leq j\leq t$. This can also be seen via the Trace Representation (14) in conjunction with properties of the trace map.
Theorem 3.14.
Let
$R_{q^{2},m}^{\ell}\cong\bigoplus_{j=1}^{t}\left(\dfrac{\mathbb{F}_{q^{2}}[x]}{\langle g_{j}\rangle}\oplus\dfrac{\mathbb{F}_{q^{2}}[x]}{\langle g^{*}_{j}\rangle}\right)^{\ell}\oplus\bigg(\bigoplus_{i=1}^{s}\dfrac{\mathbb{F}_{q^{2}}[x]}{\langle f_{i}\rangle}\bigg)^{\ell},$
where $(g_{j},g_{j}^{*})$ are pairs of reciprocal irreducible factors, for $1\leq j\leq t$, and $f_{i}$ are self-reciprocal irreducible factors, for $1\leq i\leq s$, with $f_{s}=x-1$. Let $\mathcal{C}_{u}$ be the $m^{u}\ell$-length QC code whose constituent codes (40) and (41) are all Galois closed. Additionally, assume that the constituent codes (41) are HSO. Then, we have an infinite family $\{\mathcal{C}_{u}\}_{u\geq 2}$ of HSO $q^{2}$-ary QC codes satisfying the square-root-like lower bound for their minimum distances.
4. Constructions of QECCs from QC Codes
Given the construction of $\ell$-QC codes and results seen in [cem, lingsole], along with the collection of quantum constructions in Subsection 2.4, we present in this section a construction of quantum codes with good parameters.
Example 4.1.
Let $q=4$, $\ell=3$, $m=7$, and $\alpha$ be a primitive 77-th root of unity. Let
$\left(\frac{\mathbb{F}_{4}[x]}{\langle x^{7}-1\rangle}\right)^{3}$ $\cong$ $\left(\frac{\mathbb{F}_{4}[x]}{\langle x^{3}+x+1\rangle}\right)^{3}\oplus\left(\frac{\mathbb{F}_{4}[x]}{\langle x^{3}+x^{2}+1\rangle}\right)^{3}\oplus\left(\frac{\mathbb{F}_{4}[x]}{\langle x-1\rangle}\right)^{3}$
$\cong$ $\mathbb{F}_{4}^{3}(\alpha)\oplus\mathbb{F}_{4}^{3}(\alpha^{3})\oplus\mathbb{F}_{4}^{3}.$
For any element $\gamma\in\mathbb{F}_{4}(\alpha)\setminus\mathbb{F}_{4}$, let $C_{11}^{\prime}=\left\langle(\gamma,\gamma,\gamma)\right\rangle\subset\mathbb{F}_{4}^{3}(\alpha)$ , $C_{11}^{\prime\prime}=(C_{11}^{\prime})^{\perp}\subset\mathbb{F}_{4}^{3}(\alpha^{3})$, and $C_{11}=\mathbb{F}_{4}^{3}$ be the constituent codes, whose parameters are $3,1,3]_{64}$, $3,2,2]_{64}$, and $3,3,1]_{4}$, respectively. Notice that $C_{11}^{\prime},C_{11}^{\prime\prime}$, and $C_{11}$ are Galois closed by Proposition 3.12, and $C_{11}$ is trivially HDC. Let $\mathcal{C}\subset\mathbb{F}_{4}^{m\ell}$ be the respective 33-QC-code which by Theorem 3.13 is also Galois-closed. The minimum distance of the cyclic code associated to $\mathcal{C}$ and all its associated cyclic subcodes are described below
$\begin{array}[]{lclclclclcl}BZ(D_{1}^{\perp})&=&\{\alpha^{-1}=\alpha^{6}\}&\Rightarrow&D_{1}&=&\left\langle(x^{3}+x^{2}+1)(x+1)\right\rangle&\Rightarrow&d(D_{1})&=&4\\ BZ(D_{2}^{\perp})&=&\{\alpha^{-3}=\alpha^{4}\}&\Rightarrow&D_{2}&=&\left\langle(x^{3}+x+1)(x+1)\right\rangle&\Rightarrow&d(D_{2})&=&4\\ BZ(D_{3}^{\perp})&=&\{1\}&\Rightarrow&D_{3}&=&\left\langle(x^{3}+x+1)(x^{3}+x^{2}+1)\right\rangle&\Rightarrow&d(D_{3})&=&7\\ BZ(D_{1,2}^{\perp})&=&\{\alpha^{4},\alpha^{6}\}&\Rightarrow&D_{1,2}&=&\left\langle x+1\right\rangle&\Rightarrow&d(D_{1,2})&=&2\\ BZ(D_{1,3}^{\perp})&=&\{1,\alpha^{6}\}&\Rightarrow&D_{1,3}&=&\left\langle x^{3}+x^{2}+1\right\rangle&\Rightarrow&d(D_{1,3})&=&3\\ BZ(D_{2,3}^{\perp})&=&\{1,\alpha^{4}\}&\Rightarrow&D_{2,3}&=&\left\langle x^{3}+x+1\right\rangle&\Rightarrow&d(D_{2,3})&=&3\\ BZ(D_{1,2,3}^{\perp})&=&\{1,\alpha^{4},\alpha^{6}\}&\Rightarrow&D_{1,2,3}&=&\left\langle 1\right\rangle&\Rightarrow&d(D_{1,2,3})&=&1\end{array}$
By Theorem 2.9 (See Remark 2), the minimum distance of the $\ell$-QC code $\mathcal{C}$ is lower-bounded by
(60) $d(\mathcal{C})\geq\min\{7,7,8\}=7,$
and $\mathcal{C}$ is a $21,12,\geq 7]_{4}-$EDC $\ell$-QC code by [lingsole, Theorem 4.2]. As seen in Corollary 2.11, from $\mathcal{C}$ it is possible to obtain a $21,3,d^{\perp}\geq 7]]_{4}-$quantum code $Q$. According to [table], the best known quaternary QECC for $n=21$ and $k=3$ has minimum distance $7.$
By the Lemma 2.12, it is still possible to construct more good QECCs from the explicit code given above which are listed below:
$16,8]]_{4} [[$$17,7]]_{4} [[$$18,6]]_{4} [[$$19,5]]_{4} [[$$20,4]]_{4}$
Ex. 4.1 and Lemma 2.12 - item (2) $\geq 2$ $\geq 3$ $\geq 4$ $\geq 5$ $\geq 6$
Code Table [table] 4-5 5-6 5-7 6-8 7-8
$22,3]]_{4} [[$$23,3]]_{4} [[$$24,3]]_{4} [[$$25,3]]_{4} [[$$26,3]]_{4}$
Ex. 4.1 and Lemma 2.12 - item (1) $\geq 7$ $\geq 7$ $\geq 7$ $\geq 7$ $\geq 7$
Code Table [table] 7-10 8-11 8-11 9-11 9-12
Although there is no significant improvement in the minimum distances for any of the above parameters, most of our codes achieve values that are within one or two of the best-known lower bounds. It is also worth noting that our example exhibits lower bounds of the minimum distance, so it may be possible to obtain codes with larger minimum distances.
Example 4.2.
Still working with the decomposition of $R_{4,7}^{3}$ provided in Example 4.1, assume the same constituent codes $C_{11}^{\prime}$ and $C_{11}^{{}^{\prime\prime}}$, and let $C_{11}=\langle(0,1,\beta),(1,0,\beta^{2})\rangle$ be the last constituent code, where $\beta\in\mathbb{F}_{4}$ so that $\beta^{2}=\beta+1$, whose parameters are $3,2,2]_{4}$, respectively. Observe that $C_{11}$ is EDC.
Let $\mathcal{C}\subset\mathbb{F}_{4}^{m\ell}$ be the respective 33-QC-code. By Theorem 3.7, we have an infinity family $\{\mathcal{C}_{u}\}_{u\geq 1}$ of EDC QC codes whose their parameters are
(61) $\left[3\cdot 7^{u},3\left(\frac{7^{u}+1}{2}\right),\geq 2\cdot 7^{u-1}\right]_{4}.$
From Corollary 2.11, it is possible to get the infinity family of quantum codes $\{Q_{u}\}_{u\geq 1}$ with parameters $\left[\left[3\cdot 7^{u},3,\geq 2\cdot 7^{u-1}\right]\right]_{4}$.
Example 4.3.
Let $q=2$, $m=7$, $\ell=8$, and $\alpha$ be a primitive 77-th root of unity. Assume the following decomposition of $R_{2,7}^{8}=\left(\frac{\mathbb{F}_{2}[x]}{\langle x^{7}-1\rangle}\right)^{8}$. Thus
$\left(\frac{\mathbb{F}_{2}[x]}{\langle x^{7}-1\rangle}\right)^{8}$ $\cong$ $\left(\frac{\mathbb{F}_{2}[x]}{\langle x^{3}+x+1\rangle}\right)^{8}\oplus\left(\frac{\mathbb{F}_{2}[x]}{\langle x^{3}+x^{2}+1\rangle}\right)^{8}\oplus\left(\frac{\mathbb{F}_{2}[x]}{\langle x-1\rangle}\right)^{8}$
$\cong$ $\mathbb{F}_{2}^{8}(\alpha)\oplus\mathbb{F}_{2}^{8}(\alpha^{3})\oplus\mathbb{F}_{2}^{8}.$
Let $C_{11}^{\prime}$ be a $8,3,6]_{8}$ MDS code (its construction can be checked in [table]), $C_{11}^{{}^{\prime\prime}}=C_{11}^{{}^{\prime}\perp}$ a $8,5,4]_{8}$ (also MDS) code, and $C_{11}$ be a EDC $8,6,2]_{2}-$code with basis
(62) $B$ $=$ $\left\{1,0,0,1,0,0,0,0),(0,1,0,1,0,0,0,0),(0,0,1,1,0,0,0,0),(0,0,0,0,1,0,0,1),(0,0,0,0,0,1,0,1),\right$.
$\left.(0,0,0,0,0,0,1,1)\right\}$
From the minimum distances of the associated cyclic subcodes provided in the Example 4.1, together with the minimum distances of the constituent codes described above, the minimum distance of $\mathcal{C}$ in $\mathbb{F}_{2}^{56}$ is $d\geq d_{GO}=14.$ Moreover, by Lemma 2.1, the dimension of $\mathcal{C}$ is 30, namely, $\mathcal{C}$ is a 88-QC EDC $56,30,\geq 14]_{2}-$code and, based on Corollary 2.11, it is possible to construct a QECC with parameters $56,4,\geq 14]]_{2}$. Notice that in [table], the best $56,4]]_{2}-$QECC has minimum distance 13.
By the Lemma 2.12, we list more good quantum codes from this code. We have an improvement on some lower bounds from this explicit example: the best minimum distance for $55,5]]_{2}$ can be improved to 13–18 and $57,4]]_{2}$ can be improved to 14–19.
$51,9]]_{2} [[$$52,8]]_{2} [[$$53,7]]_{2} [[$$4,6]]_{2} [[55$$5,5]]_{2}$
Ex. 4.3 and Lemma 2.12 - item (2) $\geq 9$ $\geq 10$ $\geq 11$ $\geq 12$ $\geq 13$
Code Table [table] 11–15 11–16 11–17 12–17 12–18
$57,4]]_{2} [[$$58,4]]_{2} [[$$59,4]]_{2} [[$$60,4]]_{2} [[$$61,4]]_{2}$
Ex. 4.3 and Lemma 2.12 - item (1) $\geq 14$ $\geq 14$ $\geq 14$ $\geq 14$ $\geq 14$
Code Table [table] 13–19 14–19 14–19 14–20 14–20
5. Conclusion
In general, describing the parameters of infinity families of codes is not an easy task. In this paper, we have presented families of QC codes over $\mathbb{F}_{q}$, making explicit their dimension and a lower bound to their minimum distances. In addition, based on their constituent code decomposition, we also explored their dualities according to Euclidean and Hermitian inner products. We have also addressed when this lower bound to minimum distance is greater than the square-root-like bound, which there are only a few constructions in the literature achieving this goal. Finally, some quantum codes have been reported, which are comparable with those ones depicted on the Code Table [table].
For future work, we aim to gain a deeper understanding of how using MDS codes as constituent codes affects the minimum distance of the respective QC code. It is also important to explore whether the GO-bound can be improved, since doing so would lead to better lower-bound estimates for the minimum distances of both our QC codes and their corresponding quantum codes.
6. acknowledgements
The authors would like to thank the Institute for Computational and Experimental Research in Mathematics (ICERM) for providing funding and space for us to work in a mathematically rich environment during our week-long visit in Summer 2025. The first two authors would also like to thank Simons Laufer Mathematical Sciences Institute (SLMath) for providing funding for post-programmatic travel in 2024 following our Summer Research in Mathematics visit in 2023. The second author was partially supported by the AMS-Simons Research Enhancement Grants for PUI Faculty while working on this project. The three authors also thank Dr. Lara Ismert for her conversations that inspired us to use Galois closed codes to move freely between the Hermitian and Euclidean inner products.