### v4

Abstract
Quasi-cyclic codes have been recently employed in the constructions of quantum error-correcting codes. In this paper, we propose a construction of infinite families of quasi-cyclic codes over $\mathbb{F}_{q}$ which are self-orthogonal with respect to the Euclidean and Hermitian inner products. In particular, their dimension and a lower bound for their minimum distance are computed using their constituent codes defined over field extensions of $\mathbb{F}_{q}$. We also show that the lower bound for the minimum distance satisfies the square-root-like lower bound and also show how dual-containing and self-dual quasi-cyclic codes can arise from our construction. Using the CSS construction, we show the existence of quantum error-correcting codes with good parameters.
keywords:
[…]
1 Introduction
In the study of infinite families of codes, the growth rate of the minimum distance with respect to the code length is an important measure of performance. The square-root bound is a useful parameter for measuring how good a classical error-correcting code is; namely, given $n$ and $d$ the code length and the minimum distance of a code $C$, we say $C$ is a good code if $d\geq\sqrt{n}.$ This definition was extended to the square-root-like bound when there is a positive constant $c$ such that $d\geq c\sqrt{n}$. Codes satisfying square-root-like bounds are desirable as they exhibit nontrivial distance growth while maintaining strong algebraic structure. Such bounds guarantee that the error-correcting capability of the codes increases with the code length. Moreover, the square-root bound frequently arises as a natural benchmark for structured algebraic code families, where achieving linear distance growth is often difficult. The construction of infinite families of codes satisfying such bounds has attracted significant attention, and for instance, the various codes constructed in Chen2023FourIF; squarerootbound; 28fcd52cbe024304880942df5f3617ce; WLZZ25 satisfy the square-root-like bound.
Beyond minimum distance considerations, structural properties of codes such as self-orthogonality, dual-containment, and self-duality play important roles in coding theory. Self-orthogonal codes arise in several areas, including lattice theory W98, row-self-orthogonal matrices Massey92; Massey98, and linear complementary dual codes Massey98. Self-dual codes also exhibit deep connections with group theory Hoehn00; JLLX13, lattices SloaneChapter99, and design theory Sole90, and are known to be asymptotically good MACWILLIAMS1972153. The construction of codes that simultaneously satisfy the square-root-like bound together with certain duality conditions are also of interest. For instance, the authors in infiniteselfdualfamilies provide infinite families of (Euclidean and Hermitian) self-dual cyclic codes satisfying the square-root bound, and in goodsocycliccodes, the authors explore constructions of self-orthogonal cyclic codes whose minimum distances are larger than the square-root bound. Moreover, classical codes with self-orthogonal or dual-containing properties are useful in the construction of quantum error-correcting codes through celebrated work in Calderbank-Shor1996; Steane1996, known as the CSS construction. Because of these connections and applications, determining the existence and constructions of self-orthogonal, dual-containing, and self-dual classical error-correcting codes is an important problem in coding theory.
According to infiniteselfdualfamilies, there are very few constructions of infinite families of self-dual codes whose minimum distances meet the square-root lower bound, which are depicted in that paper. To the best of our knowledge, the authors of infiniteselfdualfamilies in 2025 were the first to present infinite families of non-binary self-dual cyclic codes achieving a square-root-type lower bound on their minimum distances. Additionally, Calderbank provides a distinct construction of an infinite family of self-dual quasi-cyclic codes.
Quasi-cyclic (QC) are have been suitable for constructing structured families of codes with desirable distance and duality properties. QC codes are a natural generalization of cyclic codes and are well-studied in the classical coding theory literature. They possess a rich algebraic structure seguin; cem; lingsole and have strong connections with convolutional codes SolomonTilborg. Moreover, there exist QC codes that are asymptotically good longcodes. Very recently, some constructions of self-dual, dual-containing, and self-orthogonal QC codes have been explored to develop effective quantum error-correcting codes. For example in Abdukhalikov-Bag-Panario2023; GalindoHernando; benjwal, the authors obtained quantum codes with good parameters by studying the generators of the dual of a QC code to determine the conditions where such codes are dual-containing, self-orthogonal, or self-dual. As these properties are intrinsically related to the CSS construction, QECCs with good parameters are obtained.
In this paper, we propose a construction of infinite families of self-orthogonal, dual-containing, and self-dual QC codes using their constituent code representation described in lingsole, and whose lower bound for their minimum distances are determined using their concatenated description described in jensen; OO24. We then show that the codes in our construction yield minimum distances that satisfy a square-root-like lower bound. We also explore the existence of QECCs with good parameters arising from these constructions. The paper is organized as follows: Section 2 provides background results related to QC codes and their constituent codes, and also includes the foundations of constructing QECCs from classical codes. In Section 3, we present the main results of this paper, namely the use of constituent codes of QC codes to construct infinite families of QC codes that are self-orthogonal (dual-containing, or self-dual) with respect to the Euclidean inner product. In Section 4, we show that minimum distances of such codes satisfy the square-root-like lower bound. We then introduce the notion of Galois closed codes in Section 5 to obtain QC codes that are self-orthogonal (dual-containing, or self-dual) with respect to the Hermitian inner product. Lastly, in Section 6, we show how some codes from our construction yield quantum error-correcting codes with good parameters using the CSS construction. All computations throughout this paper have been done using Magma, SageMath, and Mathematica.
uantum error correction plays a central role in the development of quantum computing architectures. With the rapid advancement of technologies involving quantum computation and communication, there is a growing demand for effective quantum error-correcting codes (QECCs). These codes are essential for protecting quantum information from noise, which can arise from various types of quantum channels. In the celebrated work of Calderbank, Shor, and Steane Calderbank-Shor1996; Steane1996, the authors proposed the CSS construction that utilizes the duality of certain classical error-correcting codes to construct quantum error-correcting codes. This method is commonly used to construct QECCs over finite fields Ketkar-Klappenecker-Kumar-Sarvepalli2006. QECCs from the CSS construction are most often constructed using self-orthogonal error-correcting codes and its dual, where self-orthogonal codes have applications in lattice theory W98, row-self-orthogonal matrices Massey92; Massey98, linear complementary dual codes Massey98, and many others. Beyond self-orthogonality, the notions of dual-containing and self-dual codes are also significant. Self-dual codes notably exhibit deep connections with group theory Hoehn00; JLLX13, lattices SloaneChapter99, and design theory Sole90, and are recognized for being asymptotically good MACWILLIAMS1972153. Because of these applications, it is a worthwhile endeavor to determine the existence and constructions of self-orthogonal, dual-containing, and self-dual classical error-correcting codes.
The square-root bound is a useful parameter for measuring how good a classical error-correcting code is. Given $n$ and $d$ the code length and the minimum distance of a code $C$, we say $C$ is a good code if $d\geq\sqrt{n}.$ This definition was extended for square-root-like bound, when there is a positive constant $c$ so that $d\geq c\sqrt{n}$. Currently, the academic community have been focused on construct infinite families of codes satisfying the square-root bound, and their dualities. As exponents in this research area, we quote infiniteselfdualfamilies , where the authors provide infinity families of (Euclidean and Hermitian) self-dual cyclic codes satisfying the square-root. In goodsocycliccodes, the authors explore constructions of self-orthogonal cyclic codes whose their minimum distances are larger than the square-root bound.
The focus of this paper is on constructing quasi-cyclic (QC) error-correcting codes that are self-orthogonal, dual-containing and self-dual under the Euclidean and Hermitian inner products. Quasi-cyclic codes are a natural generalization of cyclic codes and are well-studied in classical coding theory literature. Quasi-cyclic codes have a rich algebraic structure seguin; cem; lingsole and have strong connections with convolutional codes SolomonTilborg. Moreover, there exist QC codes that are asymptotically good longcodes.
Very recently, some constructions of self-dual, dual-containing, and self-orthogonal QC codes have been explored to develop effective quantum error-correcting codes. For example in Abdukhalikov-Bag-Panario2023; GalindoHernando; benjwal, the authors obtained quantum codes with good parameters by studying the generators of the dual of a QC code to determine the conditions where such codes are dual-containing, self-orthogonal, or self-dual. Consequently, as these properties are intrinsically related to the CSS construction, QECCs with good parameters are obtained.
In this paper, we propose a construction of infinite families of self-orthogonal and self-dual QC codes using their constituent code representation described in lingsole and their concatenated description described in jensen; OO24. We then describe how the codes in our construction yield codes whose minimum distances are also better than the square-root bound. We also explore the existence of QECCs with good parameters. The paper is organized as follows: Section 2 provides background results related to QC codes and their constituent codes. This section also includes the foundations of constructing QECCs from classical ones. In Section 3, we present the main results of this paper, namely, the use of constituent codes of QC codes to construct infinite families of QC codes that are self-orthogonal (or self-dual) with respect to the Euclidean inner product. We show that the self-orthogonal QC codes in this construction satisfy the square-root-like lower bound to their minimum distance. We also introduce the notion of Galois-closed codes to obtain QC codes that are self-orthogonal with respect to the Hermitian inner product. Lastly in Section 6, we show how some codes from our construction are used to obtain quantum error-correcting codes with the best known parameters so far. All computations throughout this paper have been done using SageMath or Mathematica.
uantum error correction plays a central role in the development of quantum computing architectures. With the rapid advancement of technologies involving quantum computation and communication, there is a growing demand for effective quantum error-correcting codes (QECCs). These codes are essential for protecting quantum information from noise, which can arise from various types of quantum channels. In the celebrated work of Calderbank, Shor, and Steane Calderbank-Shor1996; Steane1996, the authors proposed the CSS construction that utilizes the duality of certain classical error-correcting codes to construct quantum error-correcting codes. This method is commonly used to construct QECCs over finite fields Ketkar-Klappenecker-Kumar-Sarvepalli2006. QECCs from the CSS construction are most often constructed using self-orthogonal error-correcting codes and its dual, where self-orthogonal codes have applications in lattice theory W98, row-self-orthogonal matrices Massey92; Massey98, linear complementary dual codes Massey98, and many others. Beyond self-orthogonality, the notions of dual-containing and self-dual codes are also significant. Self-dual codes notably exhibit deep connections with group theory Hoehn00; JLLX13, lattices SloaneChapter99, and design theory Sole90, and are recognized for being asymptotically good MACWILLIAMS1972153. Because of these applications, it is a worthwhile endeavor to determine the existence and constructions of self-orthogonal, dual-containing, and self-dual classical error-correcting codes.
The focus of this paper is on constructing quasi-cyclic (QC) error-correcting codes that are self-orthogonal and self-dual under the Euclidean and Hermitian inner products. Quasi-cyclic codes are a natural generalization of cyclic codes and are well-studied in classical coding theory literature. Quasi-cyclic codes have a rich algebraic structure seguin; cem; lingsole and have strong connections with convolutional codes SolomonTilborg. Moreover, there exist QC codes that are asymptotically good longcodes. In 2022, the authors in RB found an upper bound for the locality of QC locally recoverable codes using generator matrices in systematic form where the first block is the identity matrix and the remaining blocks are comprised of smaller circulant matrices. More recently in  BastosAlvarezFloresSalerno, a more concise way of finding an upper bound on the locality of these codes was studied, and a construction of optimal locally recoverable QC codes was presented.
Very recently, some constructions of self-dual, dual-containing, and self-orthogonal QC codes have been explored to develop effective quantum error-correcting codes. For example in Abdukhalikov-Bag-Panario2023; GalindoHernando; benjwal, the authors obtained quantum codes with good parameters by studying the generators of the dual of a QC code to determine the conditions where such codes are dual-containing, self-orthogonal, or self-dual. Consequently, as these properties are intrinsically related to the CSS construction, QECCs with good parameters are obtained.
We then describe how the codes in our construction yield the existence of QECCs with good parameters. The paper is organized as follows: Section 2 provides background results related to QC codes, their constituent codes, and their different representations that will be used throughout this paper. This section also includes the foundations of constructing QECCs from classical ones. In Section 3, we present the main results of this paper, namely, the use of constituent codes of QC codes to construct infinite families of QC codes that are self-orthogonal (or self-dual) with respect to the Euclidean inner product. We show that the self-orthogonal QC codes in this construction satisfy the square-root-like lower bound to their minimum distance. We also introduce the notion of Galois-closed codes to obtain QC codes that are self-orthogonal with respect to the Hermitian inner product. Lastly in Section 6, we show how some codes from our construction are used to obtain quantum error-correcting codes. All computations throughout this paper have been done using SageMath or Mathematica.
2 Preliminaries
Let $\mathbb{F}_{q}$ be a finite field with $q$ elements, where $q=p^{t}$, for $p$ a positive prime and $t\in\mathbb{Z}^{+}$. An $n,k,d]_{q}-$linear code $C$ is a $k$-dimensional $\mathbb{F}_{q}$-vector subspace of $\mathbb{F}_{q}^{n}$, where the minimum weight of nonzero codewords is $d$. In this work, only linear codes are presented, so we will refer to them as codes only. The relationship between the parameters $n,k$ and $d$ is the following Singleton bound
$d\leq n-k+1$ (1)
and the code is called a maximum distance separable code, or an MDS code, if equality in (1) is achieved. A quintessential example of a maximum distance separable code is the Reed-Solomon code described in RS1958. More generally, if we choose nonzero elements $v_{1},...,v_{n}\in\mathbb{F}_{q}$ and distinct elements $\alpha_{1},...,\alpha_{n}\in\mathbb{F}_{q}$, we define the Generalized Reed-Solomon (GRS) Codes as
$GRS_{n,k}(\mathbf{\alpha},\mathbf{v})=\{(v_{1}f(\alpha_{1}),...,v_{n}f(\alpha_{n}))\mid f(x)\in\mathbb{F}_{q}[x]_{k}\}$ (2)
where $\textbf{v}=(v_{1},...,v_{n})$, $\textbf{$\alpha$}=(\alpha_{1},...,\alpha_{n})$, and $\mathbb{F}_{q}[x]_{k}$ is the set of polynomials in $\mathbb{F}_{q}[x]$ of degree less than $k$. This family of codes is another example of an MDS code that will be useful in Example 4.37.
This paper focuses on quasi-cyclic codes, which are generalizations of cyclic codes. More formally, a quasi-cyclic (QC) code of index $\ell$ is a code $\mathcal{C}\subset\mathbb{F}_{q}^{m\ell}$ where for all $\textbf{c}\in\mathcal{C}$, $T^{\ell}(\textbf{c})\in\mathcal{C}$, where $T$ is the standard shift operator on $\mathbb{F}_{q}^{m\ell}$ and $\ell$ is the smallest positive integer with this property. If $\mathcal{C}$ is a QC code in $\mathbb{F}_{q}^{m\ell}$, then any codeword $\textbf{c}\in\mathcal{C}$ can be represented as an $m\times\ell$ array
$\textbf{c}=\begin{pmatrix}c_{00}&c_{01}&...&c_{0,\ell-1}\\ c_{10}&c_{11}&...&c_{1,\ell-1}\\ \vdots&\vdots&\ddots&\vdots\\ c_{m-1,0}&c_{m-1,1}&...&c_{m-1,\ell-1}\end{pmatrix}$ (3)
where the action of the operator $T^{\ell}$ is equivalent to a row shift in (3).
To describe the constituent code representation of a QC code in lingsole, we consider the ring $R_{q,m}^{\ell}:=\left(\mathbb{F}_{q}[x]/\langle x^{m}-1\rangle\right)^{\ell}$, where $\operatorname{gcd}(m,q)=1$. To any codeword c as in (3), we associate the following element of $R_{q,m}^{\ell}$:
$\vec{\textbf{c}}(x):=\left(c_{0}(x),c_{1}(x),...,c_{\ell-1}(x)\right)\in R_{q,m}^{\ell}$ (4)
where for each $0\leq j\leq\ell-1$,
$c_{j}(x):=c_{0j}+c_{1j}x+c_{2j}x^{2}+\cdots+c_{m-1,j}x^{m-1}\in R_{q,m}.$
By Lemma 3.1 in lingsole, the map $\phi:\mathbb{F}_{q}^{m\ell}\to R_{q,m}^{\ell}$ defined by $\phi(\textbf{c})=\vec{\textbf{c}}(x)$ is an $R_{q,m}$-module isomorphism. Hence, any QC code over $\mathbb{F}_{q}$ of length $m\ell$ and index $\ell$ can be viewed as an $R_{q,m}$-submodule of $R_{q,m}^{\ell}$.
grayThroughout this paper, we will move freely between the representations in (LABEL:flattened_codeword) and (4).
Lemma 1
(lingsole, Lemma 3.1) The map induces a one-to-one correspondence between quasi-cyclic codes over $\mathbb{F}_{q}$ of index $\ell$ and length $\ell m$ and linear codes over $R$ of length $\ell$. AA: The statement of this lemma is redundant from the paragraph directly above. Remove this lemma?G.: Yes. In fact we cite this lemma in the paragraph above
Additionally, given a polynomial $h(x)=h_{n}x^{n}+h_{n-1}x^{n-1}+\cdots+h_{0}\in\mathbb{F}_{q}[x]$, its reciprocal polynomial $h^{*}(x)$ is given by $h^{*}(x)=x^{n}h(x^{-1})$. As seen in lingsole, we may decompose $x^{m}-1\in\mathbb{F}_{q}[x]$ into irreducible factors
$x^{m}-1=\delta\prod_{j=1}^{t}g_{j}(x)g_{j}^{*}(x)\prod_{i=1}^{s}f_{i}(x),$ (5)
where $\delta$ is a non-zero element in $\mathbb{F}_{q}$, $f_{i}(x)$ is a self-reciprocal polynomial for each $1\leq i\leq s$, and $g_{i}^{*}(x)$ is the reciprocal polynomial of $g_{i}(x)$, where $g_{i}(x)\neq g_{i}^{*}(x)$, for each $1\leq j\leq t$. With this decomposition and via the Chinese remainder theorem (CRT), the ring $R_{q,m}$ may be written as
$R_{q,m}=\frac{\mathbb{F}_{q}[x]}{\langle x^{m}-1\rangle}\cong\Bigg(\bigoplus_{j=1}^{t}\bigg(\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}\bigg)\Bigg)\oplus\bigg(\bigoplus_{i=1}^{s}\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\bigg),$ (6)
where we have suppressed the explicit dependence on $x$ in the $g_{j}$ and $f_{i}$ to simplify notation. Let $\alpha$ be an $m$-th primitive root of the unity in an extension of $\mathbb{F}_{q}$ and let
$F_{i}:=\dfrac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\cong\mathbb{F}_{q}(\alpha^{u_{i}}),\quad G^{\prime}_{j}:=\dfrac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\cong\mathbb{F}_{q}(\alpha^{v_{j}}),\quad\text{and}\quad G^{\prime\prime}_{j}:=\dfrac{\mathbb{F}_{q}[x]}{\langle g^{*}_{j}\rangle}\cong\mathbb{F}_{q}(\alpha^{v^{*}_{j}}).$
for each $1\leq i\leq s$ and $1\leq j\leq t$. Hence,
$R_{q,m}^{\ell}$ $\cong\Bigg(\bigoplus_{j=1}^{t}\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}\right)^{\ell}\Bigg)\oplus\left(\bigoplus_{i=1}^{s}\left(\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\right)^{\ell}\right)$ (7)
$\cong\Bigg(\bigoplus_{j=1}^{t}(G^{\prime}_{j}\oplus G^{\prime\prime}_{j})^{\ell}\Bigg)\oplus\bigg(\bigoplus_{i=1}^{s}F_{i}^{\ell}\bigg).$ (8)
Define codes $C_{i}\subseteq F_{i}^{\ell}$, $C^{\prime}_{j}\subseteq G_{j}^{{}^{\prime}\ell}$, and $C^{\prime\prime}_{j}\subseteq G_{j}^{{}^{\prime\prime}\ell}$ for $1\leq i\leq s$ and $1\leq j\leq t$. We call these codes the constituent codes of the QC code $\mathcal{C}$ which has the following CRT decomposition:
$\mathcal{C}\cong\Bigg(\bigoplus_{j=1}^{t}(C_{j}^{\prime}\oplus C_{j}^{\prime\prime})\Bigg)\oplus\bigg(\bigoplus_{i=1}^{s}C_{i}\bigg).$ (9)
The following lemma gives an easy way to compute the dimension of a QC code $\mathcal{C}$ from its constituent codes. We slightly modified the notation of the following result from BastosAlvarezFloresSalerno to align with the notation in this paper.
Lemma 2
(BastosAlvarezFloresSalerno, Lemma 2) Let $\mathcal{C}\subset\mathbb{F}_{q}^{m\ell}$ be an $m\ell,k]$ QC code and let $C_{\mu}\subseteq\left(\frac{\mathbb{F}_{q}[x]}{\langle b_{j}\rangle}\right)^{\ell}$ be its respective constituent $\ell,k_{\mu}]$ codes. If the nontrivial constituent codes of $\mathcal{C}$ are given by $C_{1},\ldots,C_{h}$, then $k=\sum_{\mu=1}^{h}k_{\mu}\deg(b_{\mu}).$
Moreover, if the nonzero constituent codes of $\mathcal{C}$ are denoted by $C_{1},C_{2},...,C_{h}$, we assume from now on that their minimum distances satisfy the following descending-chain condition
$d(C_{1})\geq d(C_{2})\geq\cdots\geq d(C_{h}),$ (10)
which will be utilized when computing a lower bound for the minimum distance of $\mathcal{C}$.
gray
2.1 Jensen’s Concatenated Description for QC Codes
et $r\in\{1,$$\ell\}$ and let $\mathcal{C}\subseteq R_{q,m}^{\ell}$ be a QC code as an $R$-module whose generators are given as
$\left\{\left(a_{1,0}(x),...,a_{1,\ell-1}(x)\right),...,\left(a_{r,0}(x),...,a_{r,\ell-1}(x)\right)\right\}\subset R_{q,m}^{\ell}.$
Then for each i=1,2,,s,i=1,2,...,s, we have G.: $\xi$ or $\alpha$ ? Good catch! Thanks! I was just copying from Jensen’s paper and did not notice the inconsistent notation
$C_{i}=\mathrm{Span}_{F_{i}}\{\left(a_{b,0}(\alpha^{u_{i}}),...,a_{b,\ell-1}(\alpha^{u_{i}})\right)\mid 1\leq b\leq r\}$
where $F_{i}:=\dfrac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\cong\mathbb{F}_{q}(\alpha^{u_{i}})$.
uppose that the nontrivial constituent codes of a QC code $\mathcal{C}$ are given by $C_{1},\ldots,C_{h}$. Let $\theta_{i}$ be the generating primitive idempotent for the minimal cyclic code $\langle\theta_{i}\rangle$ in consideration, where details on $\theta_{i}$ can be found in (OO24, Section 2). We then have the following isomorphism given by the following maps
$\phi:\langle\theta_{i}\rangle\to\mathbb{E}_{i},\;a(x)\mapsto a(\xi^{u_{i}})\;\;\;\text{and}\;\;\psi:\mathbb{E}_{i}\to\langle\theta_{o}\rangle,\delta\mapsto\sum_{k=0}^{m-1}a_{k}x^{k}$
where $a_{k}=\frac{1}{m}\mathrm{Tr}_{\mathbb{E}/\mathbb{F}_{q}}\left(\delta\xi^{-ku_{i}}\right)$. If $C_{i}$ is a constituent code of $\mathcal{C}$ for $1\leq i\leq h$, then we denote its concatenation with $\langle\theta_{i}\rangle$ by $\langle\theta_{i}\rangle\square C_{i}$. Jensen jensen provided the following description of QC codes via concatenation, and we will adopt the formulation presented in OO24.
Theorem 3
(OO24, Theorem 2.1)
  • (i)
    Let $\mathcal{C}$ be a QC code of length $m\ell$ and index $\ell$ over $\mathbb{F}_{q}$. Then for some subset $\mathcal{I}$ of $\{1,...,s\}$, there exist linear codes $C_{i}$ of length $\ell$ over $F_{i}$, which can be explicitly described, such that
    $\mathcal{C}\cong\bigoplus_{i\in\mathcal{I}}\langle\theta_{i}\rangle\square C_{i}.$
  • (ii)
    Conversely, let $C_{i}$ be a linear code in $F_{i}^{\ell}$, for each $i\in\mathcal{I}\subseteq\{1,...,s\}$. Then,
    $\mathcal{C}\cong\bigoplus_{i\in\mathcal{I}}\langle\theta_{i}\rangle\square C_{i}$
    is a $q$-ary QC code of length $m\ell$ and index $\ell$.
As noted in (GO13, Theorem 4.1), the constituent codes in (9) are equal to the outer codes (i.e., the $C_{i}$’s) in Theorem 3.
2.2 The Cyclic Code Associated to a QC Code and a Lower Bound for its Minimum Distance
Let $s\in\mathbb{Z}_{m}$ and let $C_{s}=\{s,sq,sq^{2},...,sq^{r-1}\}\subset\mathbb{Z}_{m}$ be the $q$-cyclotomic coset of $s$, where $r$ is the smallest positive integer so that $sq^{r}\equiv s\mod m$. It is important to note that the $q$-cyclotomic cosets split $\mathbb{Z}_{m}.$ This classic concept is seen in constructions of cyclic codes, as well as in the trace representation of QC codes in  lingsole.
f $\mathbb{F}_{q}$ is a field and $\mathbb{F}_{q^{s}}$ a field extension of $\mathbb{F}_{q}$, the trace map $\operatorname{Tr}_{\mathbb{F}_{q^{s}}/\mathbb{F}_{q}}:\mathbb{F}_{q^{s}}\to\mathbb{F}_{q}$ is defined as
$\operatorname{Tr}_{\mathbb{F}_{q^{s}}/\mathbb{F}_{q}}(a):=\sum_{i=0}^{s-1}a^{q^{i}},$
and we can extend the trace map componentwise to $\mathbb{F}_{q^{s}}^{\ell}$ by defining the $\mathbb{F}_{q}$-linear map $\operatorname{Tr}_{\mathbb{F}_{q^{s}}/\mathbb{F}_{q}}:\mathbb{F}_{q^{s}}^{\ell}\rightarrow\mathbb{F}_{q}^{\ell}$ as
$\operatorname{Tr}_{\mathbb{F}_{q^{s}}/\mathbb{F}_{q}}(a_{1},\dots,a_{\ell})=\big(\operatorname{Tr}_{\mathbb{F}_{q^{s}}/\mathbb{F}_{q}}(a_{1}),\dots,\operatorname{Tr}_{\mathbb{F}_{q^{s}}/\mathbb{F}_{q}}(a_{\ell})\big).$
theorem(lingsole, Theorem 5.1) Let $\mathbb{F}_{q}$ be a finite field and $\gcd(m,q)=1$. Then, for any $\ell$, the quasi-cyclic codes over $\mathbb{F}_{q}$ of length $m\ell$ and of index $\ell$ are precisely given by the following construction. Decompose $x^{m}-1$ as per (5) . Write $\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}=F_{i}$, $\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}=G_{j}^{\prime}$, and $\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}=G_{j}^{\prime\prime}$. Let $U_{i}$ (resp. $V_{j}$ and $W_{j}$) denote the cyclotomic coset of $\mathbb{Z}/m\mathbb{Z}$ corresponding to $F_{i}$ (resp. $G_{j}^{\prime}$ and $G_{j}^{\prime\prime}$) and fix $u_{i}\in U_{i}$, $v_{j}\in V_{j}$, and $w_{j}\in W_{j}$. For each $i$, let $C_{i}$ be a code of length $\ell$ over $F_{i}$, and for each $j$, let $C_{j}^{\prime}$ be a code of length $\ell$ over $G_{j}^{\prime}$ and let $C_{j}^{\prime\prime}$ be a code of length $\ell$ over $G_{j}^{\prime\prime}$. For $\mathbf{x_{i}}\in C_{i}$, $\mathbf{y_{j}^{\prime}}\in C_{j}^{\prime}$, and $\mathbf{y_{j}^{\prime\prime}}\in C_{j}^{\prime\prime}$, and for each $0\leq g\leq m-1$, let
$c_{g}((\mathbf{x_{i}}),(\mathbf{y_{j}^{\prime}}),(\mathbf{y_{j}^{\prime\prime}}))=\sum_{i=1}^{s}\operatorname{Tr}_{F_{i}/\mathbb{F}_{q}}(\mathbf{x_{i}}\alpha^{-gu_{i}})+\sum_{j=1}^{t}\left(\operatorname{Tr}_{G_{j}^{\prime}/\mathbb{F}_{q}}(\mathbf{y_{j}^{\prime}}\alpha^{-gv_{j}})+\operatorname{Tr}_{G_{j}^{\prime\prime}/\mathbb{F}_{q}}(\mathbf{y_{j}^{\prime\prime}}\alpha^{-gw_{j}})\right),$ (11)
where $\alpha$ is a primitive $m$-th root of unity in some extension of $\mathbb{F}_{q}$. Then the code
$\mathcal{C}=\{(c_{0}((\mathbf{x_{i}}),(\mathbf{y_{j}^{\prime}}),(\mathbf{y_{j}^{\prime\prime}})),...,c_{m-1}((\mathbf{x_{i}}),(\mathbf{y_{j}^{\prime}}),(\mathbf{y_{j}^{\prime\prime}}))):\forall\;\mathbf{x_{i}}\in C_{i},\forall\;\mathbf{y_{j}^{\prime}}\in C_{j}^{\prime},\mbox{ and }\forall\;\mathbf{y_{j}^{\prime\prime}}\in C_{j}^{\prime\prime}\}$ (12)
is a quasi-cyclic code over $\mathbb{F}_{q}$ of length $m\ell$ and of index $\ell$. Conversely, every quasi-cyclic code over $\mathbb{F}_{q}$ of length $m\ell$ and of index $\ell$ is obtained through this construction. Let $C_{1}\subset\mathbb{F}_{q}^{\ell}(\alpha^{u_{a_{1}}}),...,C_{h}\subset\mathbb{F}_{q}^{\ell}(\alpha^{u_{a_{h}}})$ be the nonzero constituent codes of an $\ell$-QC code $\mathcal{C}\subset\mathbb{F}_{q}^{m\ell}$. By (cem, Lemma 4.1), the trace-formula described in (lingsole, Theorem 5.1) may be taken from an specific finite extension field $\mathbb{F}_{q^{w}}$, where $\mathbb{F}_{q}(\alpha^{u_{a_{i}}})\subset\mathbb{F}_{q^{w}}$, for all $1\leq i\leq h$, given the existence of elements $w_{i}\in\mathbb{F}_{q}(\alpha^{u_{a_{i}}})$ so that $\operatorname{Tr}_{\mathbb{F}_{q^{w}}/\mathbb{F}_{q}(\alpha^{u_{a_{i}}})}(w_{i})=1$. In this case, when taking $\mathcal{C}$ as a collection of arrays, the columns of any codeword $\textbf{c}\in\mathcal{C}$ are codewords of an $m$-length cyclic code $D:=D_{1,2,..,h}$, whose dual’s basic zero set is
$BZ(D^{\perp})=\{\alpha^{-u_{a_{i}}}\mid i=1,2,...,h\}.$
Recall that if $D$ is a cyclic code of length $m$ over $\mathbb{F}_{q}$ and $\alpha$ is a primitive $m$-th root of unity, the basic zero set of $D^{\perp}$ consists of those powers $\alpha^{-u}$ at which every codeword polynomial of $D^{\perp}$ vanishes. Equivalently, it is the set of roots of the generator polynomial of $D^{\perp}$, described via representatives of their $q$-cyclotomic cosets modulo $m$. Just as in BastosAlvarezFloresSalerno, we call $D$ the cyclic code associated to $\mathcal{C}$. Let $I=\{i_{1},i_{2},...,i_{t}\}$ be a nonempty subset of $\{1,2,...,h\}$ where
$1\leq i_{1}<i_{2}<\cdots<i_{t}\leq h.$
Define $D_{I}:=D_{i_{1},...,i_{t}}\subset D$ to be the cyclic subcode of $D$ whose dual’s basic zero set is
$BZ(D_{I}^{\perp})=\{\alpha^{-u_{a_{i}}}\mid i\in I\}$
so that $D_{I}=D_{i_{1}}\oplus D_{i_{2}}\oplus\cdots\oplus D_{i_{t}}$ and $d(D_{I})\geq d(D)$ for each $I\subset\{1,2,...,h\}$.
gray[AA: I moved this gray text up to the section 2.3] Let $C_{1}\subset\mathbb{F}_{q}^{\ell}(\alpha^{u_{a_{1}}}),...,C_{h}\subset\mathbb{F}_{q}^{\ell}(\alpha^{u_{a_{h}}})$ be the nonzero constituent codes of an $\ell$-QC code $\mathcal{C}\subset\mathbb{F}_{q}^{m\ell}$. By (cem, Lemma 4.1), the trace-formula described in (11) may be taken from an specific finite extension field $\mathbb{F}_{q^{w}}$, where $\mathbb{F}_{q}(\alpha^{u_{a_{i}}})\subset\mathbb{F}_{q^{w}}$, for all $1\leq i\leq h$, given the existence of elements $w_{i}\in\mathbb{F}_{q}(\alpha^{u_{a_{i}}})$ so that $\operatorname{Tr}_{\mathbb{F}_{q^{w}}/\mathbb{F}_{q}(\alpha^{u_{a_{i}}})}(w_{i})=1$. In this case, looking $\mathcal{C}$ as a collection of arrays, the columns of any codeword $\textbf{c}\in\mathcal{C}$ are codewords of an $m$-length cyclic code $D:=D_{1,2,..,h}$, whose dual’s basic zero set is
$BZ(D^{\perp})=\{\alpha^{-u_{a_{i}}}\mid i=1,2,...,h\}.$
[As per reviewer, include an explanation on the “basic zero set"] Recall that if $D$ is a cyclic code of length $m$ over $\mathbb{F}_{q}$ and $\alpha$ is a primitive $m$-th root of unity, the basic zero set of $D^{\perp}$ consists of those powers $\alpha^{-u}$ at which every codeword polynomial of $D^{\perp}$ vanishes. Equivalently, it is the set of roots of the generator polynomial of $D^{\perp}$, described via representatives of their $q$-cyclotomic cosets modulo $m$. Just as in  BastosAlvarezFloresSalerno, the we call $D$ the cyclic code associated to $\mathcal{C}$. In this section, we set $D_{i}=\langle\theta_{i}\rangle$, where $BZ(D_{i}^{\perp})=\{\alpha^{-u_{i}}\}$ for each $1\leq i\leq h$. Note that $D=D_{1}\oplus D_{2},\oplus\cdots D_{h}$, and that each $D_{i}$ is a subcode of $D$. Hence, $d(D_{i})\geq d(D)$ for each $1\leq i\leq h$.
Let $I=\{i_{1},i_{2},...,i_{t}\}$ be a nonempty subset of $\{1,2,...,h\}$ where
$1\leq i_{1}<i_{2}<\cdots<i_{t}\leq h.$
Define $D_{I}:=D_{i_{1},...,i_{t}}\subset D$ to be the cyclic subcode of $D$ whose dual’s basic zero set is
$BZ(D_{I}^{\perp})=\{\alpha^{-u_{a_{i}}}\mid i\in I\}$ (13)
such that $D_{I}=D_{i_{1}}\oplus D_{i_{2}}\oplus\cdots\oplus D_{i_{t}}$.
efine the following quantity:
$R_{I}(\mathcal{C})$ $=R_{i_{1},...,i_{t}}(\mathcal{C})$ (14)
$:=\begin{cases}d(C_{i_{1}})d(D_{i_{1}})&\text{if $t=1$}\\ \\ \left(d(C_{i_{1}})-d(C_{i_{2}})\right)d(D_{i_{1}})+\left(d(C_{i_{2}})-d(C_{i_{3}})\right)d(D_{i_{1},i_{2}})+\\ \hskip 128.0374pt\vdots&\text{if $t\geq 2$}\\ +\left(d(C_{i_{t-1}})-d(C_{i_{t}})\right)d(D_{i_{1},i_{2},...i_{t-1}})+d(C_{i_{t}})d(D_{i_{1},i_{2},...,i_{t}}).\end{cases}$ (15)
The following result uses Jensen’s concatenated description of QC codes (OO24, Theorem 2.1), and their associated cyclic codes and subcodes to establish a lower bound on the minimum distance of a QC code. This lower bound will be referred to Jensen’s bound where relevant details can be be found in jensen; OO24.
Theorem 4
(jensen, Theorem 4) Given a QC code $\mathcal{C}$ of length $m\ell$ and index $\ell$ over $\mathbb{F}_{q}$ with the concatenated structure ${\mathcal{C}\cong\bigoplus_{i=1}^{h}D_{i}\square C_{i}},$ and $d(C_{1})\geq d(C_{2})\geq...\geq d(C_{h})$, we have
$d(\mathcal{C})\geq d_{\text{J}}(\mathcal{C}):=\min\{N_{h},N_{h-1,h},\ldots,N_{1,2,\ldots,h}\},$ (16)
where
$N_{h}(\mathcal{C})$ $=d(C_{h})d(D_{h}),$
$N_{h-1,h}(\mathcal{C})$ $=d(C_{h-1})d(D_{h-1,h}),$
$\vdots$
$N_{1,2,\ldots,h}(\mathcal{C})$ $=d(C_{1})d(D_{1,2,\ldots,h}).$
Using the CRT decomposition (9) and Jensen’s concatenated description, we know that a QC code $\mathcal{C}$ can be decomposed as the following direct sum of concatenations
$\mathcal{C}\cong\left(\bigoplus_{j=1}^{t}\left(D_{j}^{\prime}\square C_{j}^{\prime}\oplus D_{j}^{\prime\prime}\square C_{j}^{\prime\prime}\right)\right)\oplus\left(\bigoplus_{i=1}^{s}D_{i}\square C_{i}\right).$ (17)
We will apply Theorem 4 to the QC code representation in (17) to determine a lower bound for the minimum distance for the codes in our QC family in Section 3.
theorem[Güneri-Özbudak bound](cem, Theorem 4.8) Let $\mathcal{C}$ be a QC code as in (9) with the assumption (10). Then we have
$d(\mathcal{C})\geq d_{\text{GO}}(\mathcal{C}):=\min\{R_{h}(\mathcal{C}),R_{h-1,h}(\mathcal{C}),R_{h-2,h-1,h}(\mathcal{C}),...,R_{1,2,...,h}(\mathcal{C})\}.$ (18)
lightgray
Remark 2.5.
  1. (i)
    If the number of nonzero constituent codes is two, the Jensen bound (4) becomes
    $d(\mathcal{C})\geq d_{\text{J}}(\mathcal{C}):=\min\{d(C_{2})d(D_{2}),d(C_{1})d(D)\},$
  2. (ii)
    If the number of nonzero constituent codes is three, the Jensen bound (4) becomes
    $d(\mathcal{C})\geq d_{\text{J}}(\mathcal{C}):=\min\{N_{3},N_{2,3},N_{1,2,3}\},$
    where $N_{3}=d(C_{3})d(D_{3}),N_{2,3}=d(C_{2})d(D_{2,3})$ and $N_{1,2,3}=d(C_{1})d(D)$.
2.3 The Duals of Quasi-Cyclic Codes
Recently, QC codes have been used in constructions of QECCs where orthogonality conditions on QC codes play a central role Abdukhalikov-Bag-Panario2023; GalindoHernando; benjwal . To study the duality of codes in $\mathbb{F}_{q}^{m\ell}$, we define the following inner products: Let $\textbf{a},\textbf{b}\in\mathbb{F}_{q}^{m\ell}$ where $\textbf{a}=(a_{ij})$ and $\textbf{b}=(b_{ij})$ where $0\leq i\leq m-1$ and $0\leq j\leq\ell-1$. align* a =(a_00, a_01,…,a_0,ℓ-1, a_10,…, a_1, ℓ-1,…,a_m-1,0,…,a_m-1, ℓ-1)
b =(b_00, b_01,…,b_0,ℓ-1, b_10,…, b_1, ℓ-1,…,b_m-1,0,…,b_m-1, ℓ-1). The Euclidean inner product on $\mathbb{F}_{q}^{m\ell}$ is
$\mathbf{a}\cdot\mathbf{b}=\langle\textbf{a},\textbf{b}\rangle_{E}=\sum_{i=0}^{m-1}\sum_{j=0}^{\ell-1}a_{ij}b_{ij},$
and the Hermitian inner product on $\mathbb{F}_{q^{t}}^{m\ell}$, for even $t$ is
$\langle\textbf{a},\textbf{b}\rangle_{H}=\sum_{i=0}^{m-1}\sum_{j=0}^{\ell-1}a_{ij}b_{ij}^{\sqrt{q^{t}}}.$
In particular, when $\textbf{a},\textbf{b}\in\mathbb{F}_{q^{2}}^{m\ell}$, then $\langle\textbf{a},\textbf{b}\rangle_{H}=\sum_{i=0}^{m-1}\sum_{j=0}^{\ell-1}a_{ij}b_{ij}^{q}$. On $R_{q,m}^{\ell}$, the Hermitian inner product is as follows: for $\textbf{x}=(x_{0},x_{1},...,x_{\ell-1})$, and $\textbf{y}=(y_{0},y_{1},...,y_{\ell-1})$,
$\langle\textbf{x},\textbf{y}\rangle_{H}=\sum_{j=0}^{\ell-1}x_{j}\overline{y_{j}},$
where $\bar{\;}$ denotes the “conjugation map" on $R_{q,m}$ that sends $x$ to $x^{-1}=x^{m-1}$ and is extended $\mathbb{F}_{q}$-linearly. The following proposition is useful for determining when orthogonality in $R_{q,m}^{\ell}$ implies orthogonality in $\mathbb{F}_{q}^{m\ell}$.
Proposition 2.6.
(lingsole, Proposition 3.2) Let $\mathbf{a},\mathbf{b}\in\mathbb{F}_{q}^{m\ell}$. Then $(T^{k\ell}(\mathbf{a}))\cdot\mathbf{b}=0$ for all $0\leq k\leq m-1$ if and only if $\langle\phi(\mathbf{a}),\phi(\mathbf{b})\rangle_{H}=0$, where $\phi$ is the aforementioned $R_{q,m}$-module isomorphism.
Let $\mathcal{C}$ be a QC code in $\mathbb{F}_{q}^{m\ell}$ and define its Euclidean dual $\mathcal{C}^{\perp_{E}}$ by
$\mathcal{C}^{\perp_{E}}:=\{\textbf{x}\in\mathbb{F}_{q}^{m\ell}\mid\langle\textbf{x},\textbf{c}\rangle_{E}=0,\forall\;\textbf{c}\in\mathcal{C}\}.$
Similarly, the Hermitial dual, $\mathcal{C}^{\perp_{H}}$ of $\mathcal{C}$ is given by
$\mathcal{C}^{\perp_{H}}:=\{\textbf{x}\in\mathbb{F}_{q^{t}}^{m\ell}\mid\langle\textbf{x},\textbf{c}\rangle_{H}=0,\forall\;\textbf{c}\in\mathcal{C}\}.$
We say that $\mathcal{C}$ is self-orthogonal with respect to the Euclidean inner product (or ESO) if $\mathcal{C}\subseteq\mathcal{C}^{\perp_{E}}$, dual-containing with respect to the Euclidean inner product (or EDC) if $\mathcal{C}^{\perp_{E}}\subseteq\mathcal{C}$, and self-dual with respect to the Euclidean inner product (or ESD) if $\mathcal{C}^{\perp_{E}}=\mathcal{C}$. Analogously, $\mathcal{C}$ is self-orthogonal (HSO), dual-containing (HDC), and self-dual (HSD) with respect to the Hermitian inner product if $\mathcal{C}\subseteq\mathcal{C}^{\perp_{H}},\mathcal{C}^{\perp_{H}}\subseteq\mathcal{C}$, and $\mathcal{C}=\mathcal{C}^{\perp_{H}}$, respectively.
lightgray
Corollary 2.7.
(lingsole, Corollary 3.3) Let $\mathcal{C}$ be a quasi-cyclic code over $\mathbb{F}_{q}$ of length $m\ell$ and of index $\ell$ and let $\phi(\mathcal{C})$ be its image in $R_{q,m}^{\ell}$ under $\phi$. Then $\phi(\mathcal{C})^{\perp}=\phi(\mathcal{C}^{\perp})$, where the dual in $\mathbb{F}_{q}^{m\ell}$ is taken with respect to the Euclidean inner product, while the dual in $R_{q,m}^{\ell}$ is taken with respect to the Hermitian inner product. In particular, a quasi-cyclic code $\mathcal{C}$ over $\mathbb{F}_{q}$ is self-dual with respect to the Euclidean inner product if only if $\phi(\mathcal{C})$ is self-dual over $R_{q,m}$ with respect to the Hermitian inner product.
The following result, which we slightly modified to better align with the notation in this paper, allows us to characterize the duality of QC codes using constituent codes and will be used throughout this paper to determine the duality of the QC codes in our code construction in Section 3.
Theorem 2.8.
(lingsole, Theorem 4.2) An $\ell$-quasi-cyclic code $\mathcal{C}$ of length $m\ell$ over $\mathbb{F}_{q}$ is self-dual with respect to the Euclidean inner product if and only if
$\mathcal{C}\cong\left(\bigoplus_{j=1}^{t}(C_{j}^{\prime}\oplus(C^{\prime}_{j})^{\perp_{E}})\right)\oplus\left(\bigoplus_{i=1}^{s}C_{i}\right)$ (19)
where for $1\leq i\leq s$, $C_{i}$ is a self-dual code over $F_{i}$ of length $\ell$ (with respect to the Hermitian inner product) and, for $1\leq j\leq t$, $C^{\prime}_{j}$ is a code of length $\ell$ over $G^{\prime}_{j}$ and $C_{j}^{{}^{\prime\prime}}=(C^{\prime}_{j})^{\perp_{E}}$ is its dual with respect to the Euclidean inner product over $G_{j}^{{}^{\prime\prime}}$.
Remark 2.9.
  1. (i)
    As observed in (HKS, page 136), the irreducible factors $f_{i}$ which are self-reciprocal imply that the cardinality $q_{i}$ of each field extension $F_{i}$ is an even power of $q$, for all $i$. However, there are two exceptions: The first one, for all $m$ and $q$, is the field extension coming from the irreducible factor x1x-1 of $x^{m}-1$. The other exception occurs when $q$ is odd and $m$ is even in which $x+1$ is another self-reciprocal factor of $x^{m}-1$. In both exceptions, $q_{i}=q$, so we can equip these fields with the Euclidean inner product. Otherwise, we equip $F_{i}$ with the Hermitian inner product.
  2. (ii)
    When the constituent codes $C_{i}$ over finite extension fields coming from self-reciprocal factors are self-dual, self-orthogonal, or dual-containing (under the Hermitian or Euclidean inner products), for all $1\leq i\leq s$, then the respective $\ell$-QC code $\mathcal{C}$ is self-dual, self-orthogonal, or dual-containing, respectively. Therefore, the duality of $\mathcal{C}$ is managed by the duality of its constituent codes over finite extension fields obtained from self-reciprocal polynomials.
lightgray Thus, when we apply Theorem 2.8 to Jensen’s concatenated description in (17), we obtain
$\mathcal{C}\cong\left(\bigoplus_{j=1}^{t}\left((D_{j}^{\prime}\square C_{j}^{\prime})\oplus(D_{j}^{\prime}\square C_{j}^{\prime})^{\perp_{E}}\right)\right)\oplus\left(\bigoplus_{i=1}^{s}D_{i}\square C_{i}\right).$ (20)
Such representation of a QC code $\mathcal{C}$ will be vital when computing a lower bound for the minimum distance of $\mathcal{C}$ in Section 3. The next result extends (lingsole, Theorem 4.2) to the dual of a QC code.
Proposition 2.10.
(HKS, Proposition 7.3.5) Let $\mathcal{C}$ be a QC code with CRT decomposition as (9). Then its Euclidean dual $\mathcal{C}^{\perp_{E}}$ is of the form
$\mathcal{C}^{\perp_{E}}\cong\left(\bigoplus_{j=1}^{t}(C_{j}^{\prime}\oplus C_{j}^{{}^{\prime}\perp_{E}})\right)\oplus\left(\bigoplus_{i=1}^{s}C_{i}^{\perp_{H}}\right).$
Since the duality of $\mathcal{C}$ can be determined from its constituent codes, we will use the following results in Li for constructing self-orthogonal codes.
lightgray Thus, when we apply Proposition 2.10 to Jensen’s concatenated description in (17), we obtain
$\mathcal{C}^{\perp_{E}}\cong\left(\bigoplus_{j=1}^{t}\left(D_{j}^{\prime}\square C_{j}^{\prime}\oplus D_{j}^{{}^{\prime}\perp_{E}}\square C_{j}^{{}^{\prime}\perp_{E}}\right)\right)\oplus\left(\bigoplus_{i=1}^{s}D_{i}\square C_{i}^{\perp_{H}}\right).$ (21)
Such representation of a QC code $\mathcal{C}$ is vital when dealing with the duality of quasi-cyclic codes $\mathcal{C}$.
Lemma 2.11.
(HillNewton, Lemma 2.10) Suppose $C_{1}$ and $C_{2}$ are $n_{1},k,d_{1}]$ and $n_{2},k,d_{2}]$ codes over $\mathbb{F}_{q}$, respectively. If $C_{1}$ and $C_{2}$ have generator matrices $G_{1}$ and $G_{2}$, respectively, then $G_{1}|G_{2}]$ generates an $n_{1}+n_{2},k,d_{1}+d_{2}]$ code over $\mathbb{F}_{q}$.
Theorem 2.12.
(Li, Theorem 1)
  1. (i)
    Suppose $C_{1}$ and $C_{2}$ are $n_{1},k,d_{1}]$ and $n_{2},k,d_{2}]$ self-orthogonal codes over $\mathbb{F}_{q}$, respectively. If $C_{1}$ and $C_{2}$ have generator matrices $G_{1}$ and $G_{2}$, respectively, then $G_{1}|G_{2}]$ generates an $n_{1}+n_{2},k,d_{1}+d_{2}]$ self-orthogonal code over $\mathbb{F}_{q}$.
  2. (ii)
    Suppose $C_{1}$ and $C_{2}$ are $n_{1},k,d_{1}]$ and $n_{2},k-1,d_{2}]$ self-orthogonal codes over $\mathbb{F}_{q}$, respectively. If $C_{1}$ contains a codeword of weight at least $d_{1}+d_{2}$, then there exists an $n_{1}+n_{2},k,d_{1}+d_{2}]$ self-orthogonal code over $\mathbb{F}_{q}$.
In the special case when $C_{1}=C_{2}$, the code with generator matrix $G_{1}|G_{1}]$ will be denoted as $2C_{1}$, and such an argument may be naturally extended to $mC_{1}$, for any positive integer $m$. In this work, for any code $C$, we will refer to the code $mC$ as $m$-copies of $C$ which will play an important role in the next section when determining the dimension of the QC codes in our infinite family.
2.4 Constructions of Quantum Error-Correcting Codes from Classical Ones
From Ketkar-Klappenecker-Kumar-Sarvepalli2006, the use of classical codes has became one of the standard methods for constructing quantum error-correcting codes. A quantum error-correcting code (QECC) $Q$ is a $K$-dimensional subspace of $(\mathbb{C}^{q})^{\otimes n}$. If $Q$ has minimum distance $d$, then we say that $Q$ is an $n,K,d]]_{q}$ code. If $K=q^{k}$, we write $n,k,d]]_{q}$. The length $n$, the dimension $K$, and minimum distance $d$ are the parameters of $Q$.
f $Q$ is a quantum code with minimum distance $d$, $Q$ can detect up to d1d-1 quantum errors, and can correct up to $\lfloor\frac{d-1}{2}\rfloor$ quantum errors. The celebrated Singleton bound (LaGuardia2021, Lemma 3.5.1) says that the parameters of an $n,K,d]]_{q}$ quantum code satisfy
$\log_{q}(K)\leq n-2d+2$
or equivalently $k+2d\leq n+2$. If a quantum code $Q$ attains the quantum Singleton bound, $Q$ is called a quantum maximum distance separable (quantum MDS) code.
This paper focuses on the existence of quantum stabilizer codes, where a stabilizer (quantum) code $Q\neq\{\textbf{0}\}$ is the common eigenspace of a commutative subgroup of the error group generated by a nice basis of $(\mathbb{C}^{q})^{\otimes n}$. The code $Q$ is said to be pure to d, or simply pure, if and only if its stabilizer group does not contain non-scalar matrices of weight less than its minimum distance $d$. (See Definitions 3.5.7 and 3.5.8 in LaGuardia2021.) Moreover, a code $Q$ is said to be impure if and only if there are non-identity stabilizer elements of weight less than the minimum distance. n other words,
[…]
Definition 2.13.
The code $Q$ is said to be pure to l if and only if its stabilizer group does not contain non-scalar matrices of weight less than $l$; $Q$ is pure if and only if its pure to its minimum distance. G.: I did not understand this definition :/
Lemma 2.14.
(Ketkar-Klappenecker-Kumar-Sarvepalli2006, Corollary 80) The parameters of an $n,K,d]]_{q}$ stabilizer code with $K>1$ satisfy the inequality
$K\leq q^{n-2d+2}.$ (23)
In 1996, the following construction was introduced by Calderbank and Shor Calderbank-Shor1996, and Steane Steane1996, and is notably the most direct link between classical and quantum coding theory.
Theorem 2.15 (CSS construction Ketkar-Klappenecker-Kumar-Sarvepalli2006).
Let $C_{1}$ and $C_{2}$ denoted two classical codes with parameters $n,k_{1},d_{1}]_{q}$ and $n,k_{2},d_{2}]_{q}$ such that $C_{2}^{\perp}\subseteq C_{1}$. Then there exists $n,k_{1}+k_{2}-n,d]]_{q}$ stabilizer code with minimum distance $d=\min\{\operatorname{wt}(\mathbf{c})\mid\mathbf{c}\in(C_{1}\backslash C_{2}^{\perp})\cup(C_{2}\backslash C_{1}^{\perp})\}$ that is pure to $\min\{d_{1},d_{2}\}$.
theorem[CSS construction] Let $C_{1}$ and $C_{2}$ be classic linear codes over $\mathbb{F}_{q}$. If $C_{2}\subset C_{1}\subset\mathbb{F}_{q}^{n}$, there exisst a stabilizer quantum code with parameters
$n,\dim(C_{1})-\dim(C_{2}),\min\{d(C_{1}\backslash C_{2}),d(C_{2}^{\perp}\backslash C_{1}^{\perp})\}]]_{q}$ (24)
where $C_{i}^{\perp}$ is the Euclidean dual of $C_{i}$, for $i=1,2$. In the special case when $C_{2}=C_{1}$, we have that $C_{1}$ is a dual-containing code. This yields the following result.
Corollary 2.16.
Aly-Klappenecker-Sarvepalli2007; Ketkar-Klappenecker-Kumar-Sarvepalli2006 Let $C$ be a classical $[n,k,d]$ code over $\mathbb{F}_{q}$. If $C$ is dual-containing, then there exists a quantum stabilizer code $Q$ with parameters $n,2k-n,\geq d]]_{q}$ that is pure to $d$. Moreover, if the minimum distance of $C^{\perp}$ exceeds $d$, then the quantum code $Q$ has minimum distance $d$.
In recent years, researchers in coding theory have worked diligently to construct good (e.g., quantum MDS) quantum codes from classical Euclidean, Hermitian, or symplectic self-orthogonal, dual-containing, or self-dual codes Abdukhalikov-Bag-Panario2023; Ball21; Cao24; DinhLeNguyenTansuchat21; FangFu19; GalindoHernando; galindo; GaoSunYanZhao25; JinXing13; LiuLiu21; ShiYueChang18; TianZhu19; WanZhengZhu24. Moreover, the following results (which will be used in the examples in Section 6) present new quantum code constructions utilizing the lengthening, shortening, and dimension reduction of a given stabilizer quantum code.
Lemma 2.17.
(Ketkar-Klappenecker-Kumar-Sarvepalli2006, Lemmas 69, 70, 71, and Corollary 73)
  1. 1.
    If an $n,k,d]]_{q}$ stabilizer code exists for $k>0$, then there exists an impure $n+1,k,d]]_{q}$ stabilizer code.
  2. 2.
    If a pure $n,k,d]]_{q}$ stabilizer code exists with $n\geq 2$ and $d\geq 2$, then there exists a $n-1,k+1,d-1]]_{q}$ pure stabilizer code.
  3. 3.
    If a (pure) $n,k,d]]_{q}$ stabilizer code exists, with $k\geq 2$ $(k\geq 1)$, then there exists an $n,k-1,d^{*}]]_{q}$ stabilizer code (pure to $d$) such that $d^{*}\geq d$.
  4. 4.
    Suppose that an $n,q^{k},d]]_{q}$ and an $n^{\prime},q^{k^{\prime}},d^{\prime}]]_{q}$ stabilizer code exist. Then there exists an $n+n^{\prime},q^{k+k^{\prime}},\min\{d,d^{\prime}\}]]_{q}$ stabilizer code.
3 A Construction of ESO/EDC/ESD QC Codes From Their Constituent Codes
In this section, we discuss the main results of this paper where we propose a recursive construction of infinite families $\{\mathcal{C}_{u}\}_{u\geq 1}$ of good QC codes that are self-orthogonal (or dual-containing or self-dual) with respect to the Euclidean inner product. Such constructions of QC codes will be utilized to show the existence of QECCs in Section 6.
3.1 Recursive Construction of Infinite Families of QC Codes
To provide a careful description of our construction, we first describe a general case of the construction and then specify certain conditions on the constituent codes to explicitly describe the code dimension and lower bound for its minimum distance. Let $x^{m}-1$ factor into $2t+s$ irreducible factors as in (7), where $f_{s}=x-1$. equation* x^m-1 = δg_1 g_1^* ⋯g_t g_t^* f_1⋯f_s-1 f_s, where $\delta\in\mathbb{F}_{q}$ is nonzero, $g_{1},g_{1}^{*},...,g_{t},g_{t}^{*}$ are the factors grouped in pairs with their reciprocal polynomials for each $1\leq j\leq t$, and $f_{1},...,f_{s}$ are self-reciprocal polynomials for $1\leq i\leq s$, where $f_{s}=x-1$. Then,
$R^{\ell}_{q,m}\cong\bigoplus_{j=1}^{t}\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}\right)^{\ell}\oplus\left(\bigoplus_{i=1}^{s-1}\left(\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\right)^{\ell}\right)\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{\ell}.$ (25)
equation* R^ℓ_q,m ≅(Fq[x]⟨g1Fq[x]⟨g1* )^ℓ⊕⋯⊕( Fq[x]⟨gtFq[x]⟨gt* )^ℓ⊕(Fq[x]⟨f1 )^ℓ⊕⋯⊕(Fq[x]⟨fs-1 )^ℓ⊕(Fq[x]⟨x-1 ⟩)^ℓ. Let $\mathcal{C}_{1}\subseteq\mathbb{F}_{q}^{m\ell}$ be an $\ell$-QC code whose first $2t$ constituent codes are direct sums of arbitrary codes and their duals, i.e.,
$\bigoplus_{j=1}^{t}\left(C^{\prime}_{1,j}\oplus{C^{\prime}}_{1,j}^{\perp_{E}}\right)\subset\bigoplus_{j=1}^{t}\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}\right)^{\ell}$
and whose last $s$ constituent codes are
$\left(\bigoplus_{i=1}^{s-1}C_{1,i}\right)\oplus C_{1,s}\subset\left(\bigoplus_{i=1}^{s-1}\left(\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\right)^{\ell}\right)\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{\ell}$
where the codes $C_{1,i}$ are HSO codes for $1\leq i\leq s-1$ and $C_{1,s}$ is ESO (see Remark 2.9). Thus,
$\mathcal{C}_{1}\cong\Bigg(\bigoplus_{j=1}^{t}\left(C^{\prime}_{1,j}\oplus{C^{\prime}}_{1,j}^{\perp_{E}}\right)\Bigg)\oplus\left(\bigoplus_{i=1}^{s-1}C_{1,i}\right)\oplus C_{1,s}.$ (26)
As $C_{1,s}\subset\mathbb{F}_{q}^{\ell}$, we can assume without much loss of generality that the it is the constituent code of $\mathcal{C}_{1}$ with the smallest minimum distance. Moreover, as the $C_{1,i}$ are HSO for each $1\leq i\leq s-1$ and $C_{1,s}$ is ESO, Theorem 2.8 tells us that $\mathcal{C}_{1}$ is an ESO $\ell$-QC code of length $m\ell$. Now using $\mathcal{C}_{1}$, we construct another QC code $\mathcal{C}_{2}$ as follows: Let
$R^{m\ell}_{q,m}\cong\bigoplus_{j=1}^{t}\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}\right)^{m\ell}\oplus\left(\bigoplus_{i=1}^{s-1}\left(\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\right)^{m\ell}\right)\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{m\ell}.$
Let $\mathcal{C}_{2}\subseteq\mathbb{F}_{q}^{m^{2}\ell}$ be the $m\ell$-QC code whose first $2t$ constituent codes are the following direct sums of arbitrary codes and their duals,
$\bigoplus_{j=1}^{t}\left(C^{\prime}_{2,j}\oplus{C^{\prime}}_{2,j}^{\perp_{E}}\right)\subset\bigoplus_{j=1}^{t}\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}\right)^{m\ell},$
and whose last $s$ constituent codes are
$\left(\bigoplus_{i=1}^{s-1}C_{2,i}\right)\oplus\mathcal{C}_{1}\subset\left(\bigoplus_{i=1}^{s-1}\left(\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\right)^{m\ell}\right)\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{m\ell}$
where the $C_{2,i}$ are HSO codes for $1\leq i\leq s-1$ and $\mathcal{C}_{1}\subset\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{m\ell}\cong\mathbb{F}_{q}^{m\ell}$ is an ESO constituent code of $\mathcal{C}_{2}$. Thus,
$\mathcal{C}_{2}\cong\Bigg(\bigoplus_{j=1}^{t}\left(C^{\prime}_{2,j}\oplus{C^{\prime}}_{2,j}^{\perp_{E}}\right)\Bigg)\oplus\left(\bigoplus_{i=1}^{s-1}\;C_{2,i}\right)\oplus\mathcal{C}_{1}$
Again by Theorem 2.8, $\mathcal{C}_{2}$ is an ESO $m\ell$-QC code of length $m^{2}\ell$. To extend this recursive construction, we have that any $u\geq 2$,
$R^{m^{u-1}\ell}_{q,m}\cong\bigoplus_{j=1}^{t}\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}\right)^{m^{u-1}\ell}\oplus\left(\bigoplus_{i=1}^{s-1}\left(\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\right)^{m^{u-1}\ell}\right)\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{m^{u-1}\ell}.$
and define an $m^{u-1}\ell$-QC code $\mathcal{C}_{u}$ in $\mathbb{F}_{q}^{m^{u}\ell}$ equation C_u ≅(C’_u1⊕C’_u1^⟂) ⊕(C’_u2⊕C’_u2^⟂) ⊕⋯⊕(C’_u,t⊕C’_u,t^⟂) ⊕C_u1 ⊕⋯⊕C_u,s-1 ⊕C_u-1, whose first $2t$ arbitrary constituent codes are
$\bigoplus_{j=1}^{t}\left(C^{\prime}_{u,j}\oplus{C^{\prime}}_{u,j}^{\perp_{E}}\right)\subset\bigoplus_{j=1}^{t}\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}\right)^{m^{u-1}\ell}$ (27)
and whose last $s$ constituent codes are
$\left(\bigoplus_{i=1}^{s-1}\;C_{u,i}\right)\oplus\mathcal{C}_{u-1}\subset\left(\bigoplus_{i=1}^{s-1}\left(\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\right)^{m^{u-1}\ell}\right)\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{m^{u-1}\ell},$ (28)
where the $C_{u,i}$ are HSO codes for all $1\leq i\leq s-1$ and $\mathcal{C}_{u-1}\subset\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{m^{u-1}\ell}\cong\mathbb{F}_{q}^{m^{u-1}\ell}$ is an ESO constituent code of $\mathcal{C}_{u}$. So by Theorem 2.8,
$\mathcal{C}_{u}\cong\Bigg(\bigoplus_{j=1}^{t}\left(C_{u,j}^{\prime}\oplus{C^{\prime}}_{u,j}^{\perp_{E}}\right)\Bigg)\oplus\left(\bigoplus_{i=1}^{s-1}C_{u,i}\right)\oplus\mathcal{C}_{u-1},$ (29)
is an ESO $m^{u-1}\ell$-QC code of length $m^{u}\ell$ for all $u\geq 2$.
Remark 3.18.
Since the dual of an ESO code is EDC, our codes $\mathcal{C}_{u}^{\perp}$ will be an EDC $m^{u-1}\ell$-QC code of length $m^{u}\ell$ for all $u\geq 2$, in which case Proposition 2.10 is utilized.
3.2 The Dimension of the Family of Codes $\{\mathcal{C}_{u}\}_{u\geq 1}$
In the next lemma, we determine a formula for the dimension of each $\mathcal{C}_{u}$ in our recursive construction which depend exclusively on the dimensions of constituent codes of $\mathcal{C}_{1}$. To compute the dimension, we restrict to the special case in which the HSO constituent codes of $\mathcal{C}_{u}$ are obtained by taking $m^{u-1}$-fold copies of the HSO constituent codes of $\mathcal{C}_{1}$; that is, $C_{u,i}=m^{u-1}C_{1,i}$ for all $1\leq i\leq s-1$. This choice allows us to systematically control how the parameters of $\mathcal{C}_{u}$ evolve with $u$, since taking copies preserves the dimension while scaling the length and minimum distance in a predictable way; namely, since each $C_{1,i}$ is HSO, it follows from a generalization of Part (i) of Theorem 2.12 that each $C_{u,i}$ is also HSO. Moreover, if $C_{1,i}$ has parameters $\ell,k_{1,i},d_{1,i}]$, then $C_{u,i}$ has parameters $m^{u-1}\ell,\,k_{1,i},\,m^{u-1}d_{1,i}]$ for all $1\leq i\leq s-1$. While this assumption restricts us to a special case, it provides a tractable setting for analysis, and extensions to more general constructions will be explored in future work.
Lemma 3.19.
Let $R_{q,m}^{\ell}$ be as in (25). Then for $u\geq 1$, there exists an $m^{u}\ell$-length QC code $\mathcal{C}_{u}$ with constituent codes (27) and (28), whose dimension is
$\dim(\mathcal{C}_{u})=\ell\left(\frac{m^{u}-1}{m-1}\right)\left(\sum_{j=1}^{t}\deg(g_{j})\right)+u\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1,i}\right)+k_{1,s}.$
where $k_{1,i}=\dim(C_{1,i})$, for all $1\leq i\leq s$.
Proof 3.20.
Let $\mathcal{C}_{1}$ be the QC code with decomposition on constituent codes as
$\mathcal{C}_{1}\cong\left(\bigoplus_{j=1}^{t}\left({C_{1,j}^{{}^{\prime}}}\oplus{C_{1,j}^{{}^{\prime}}}^{\perp_{E}}\right)\right)\oplus\left(\bigoplus_{i=1}^{s}C_{1,i}\right).$
Using Lemma 2 and the facts that $\deg(g_{j})=\deg(g_{j}^{*})$ and $\dim({C^{\prime}}_{1,j}^{\perp_{E}})=\ell-k^{\prime}_{1,j}$ for all $1\leq j\leq t$, we have that
$\dim(\mathcal{C}_{1})$ $=$ $\left(\sum_{j=1}^{t}\deg(g_{j})k_{1,j}^{\prime}+\deg(g_{j}^{*})(\ell-k_{1,j}^{\prime})\right)+\left(\sum_{i=1}^{s}\deg(f_{i})k_{1,i}\right)$
$=$ $\ell\sum_{j=1}^{t}\deg(g_{j})+\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1,i}\right)+k_{1,s}.$
For any $u\geq 2$, let $\mathcal{C}_{u}$ be the QC code defined in (29) where the $C_{u,i}$ are $m^{u-1}$-copies of $C_{1,i}$, namely, $C_{u,i}=m^{u-1}C_{u,i}$, for all $1\leq i\leq s$. Given its recursive argument, $\dim(\mathcal{C}_{u})$ is computed via induction. The case when $u=1$ is done. Now assume that the argument is valid for any integer $u$, i.e., there exists a QC code $\mathcal{C}_{u}\subset\mathbb{F}_{q}^{m^{u}\ell}$ with the following dimension
$\dim(\mathcal{C}_{u})=\ell\left(\frac{m^{u}-1}{m-1}\right)\left(\sum_{j=1}^{t}\deg(g_{j})\right)+u\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1,i}\right)+k_{1,s}$ (30)
where $k_{1,i}=\dim(C_{1,i})=\dim(m^{u-1}C_{1,i})=\dim(C_{u,i})$ for all $1\leq i\leq s-1$. Now let $\mathcal{C}_{u+1}\subset\mathbb{F}_{q}^{m^{u+1}\ell}$ be the QC code with the following $m^{u}\ell$-length constituent codes
$\mathcal{C}_{u+1}\cong\left(\bigoplus_{j=1}^{t}\left(C_{u+1,j}^{\prime}\oplus C_{u+1,j}^{{}^{\prime}\perp_{E}}\right)\right)\oplus\left(\bigoplus_{i=1}^{s-1}{C_{u+1,i}}\right)\oplus\mathcal{C}_{u}$
where each $C_{u+1,i}=m^{u}C_{1,i}$ are also HSO for all $1\leq i\leq s-1$ (see Theorem 2.12). Using (30), we have hus,
$dim_{q}\mathcal{C}_{u+1}$ $=$ $\sum_{j=1}^{t}\deg(g_{j})k_{uj}^{\prime}+(m^{u}\ell-\deg(g_{j}^{*})k_{uj}^{\prime})+\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}+dim_{q}\mathcal{C}_{u}$ (31)
$=$ $\sum_{j=1}^{t}\deg(g_{j})k_{uj}^{\prime}+(m^{u}\ell-\deg(g_{j})k_{uj}^{\prime})+\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}+\left(\frac{m^{u}-1}{m-1}\right)t\ell+k_{1s}+u\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}$
$=$ $m^{u}\ell t+\left(\frac{m^{u}-1}{m-1}\right)t\ell+(u+1)\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}+k_{1s}$
$=$ $\left(\frac{m^{u+1}-1}{m-1}\right)t\ell+(u+1)\sum_{i=1}^{s-1}\deg(f_{i})k_{1i}+k_{1s}.$
$\dim(\mathcal{C}_{u+1})$ $=\sum_{j=1}^{t}k^{\prime}_{u+1,j}\deg(g_{j})+\sum_{j=1}^{t}\deg(g_{j}^{*}){k^{\prime}}_{u+1,j}^{\perp_{E}}+\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1,i}\right)+\dim(\mathcal{C}_{u})$
$=\sum_{j=1}^{t}\left[\deg(g_{j})k^{\prime}_{u+1,j}+\deg(g_{j})\left(m^{u}\ell-k^{\prime}_{u+1,j}\right)\right]+\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1,i}\right)+\dim(\mathcal{C}_{u})$
$=m^{u}\ell\sum_{j=1}^{t}\deg(g_{j})+\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1,i}\right)+\ell\left(\frac{m^{u}-1}{m-1}\right)\left(\sum_{j=1}^{t}\deg(g_{j})\right)+u\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1,i}\right)+k_{1,s}$
$=\ell\left(\frac{m^{u+1}-1}{m-1}\right)\left(\sum_{j=1}^{t}\deg(g_{j})\right)+(u+1)\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1,i}\right)+k_{1,s}$
as desired. $\qed$
3.3 A Lower Bound for the Minimum Distance of the Family of Codes $\{\mathcal{C}_{u}\}_{u\geq 1}$
In this section, we will use Jensen’s bound to find a lower bound for the minimum distance of the QC code $\mathcal{C}_{u}\subset\mathbb{F}_{q}^{m^{u}\ell}$ for each $u\geq 1$. o do so, we will use Jensen’s concatenated description for each $u\geq 2$,
$\mathcal{C}_{u}\cong\left(\bigoplus_{j=1}^{t}\left(D_{uj}^{\prime}\square C_{uj}^{\prime}\oplus D_{uj}^{{}^{\prime}\perp_{E}}\square C_{uj}^{{}^{\prime}\perp_{E}}\right)\right)\oplus\left(\bigoplus_{i=1}^{s-1}D_{ui}\square C_{ui}\right)\oplus D_{u-1}\square\mathcal{C}_{u-1}.$ (32)
Let $\mathcal{C}_{u}$ be a ESO $m^{u-1}\ell$-QC code in $\mathbb{F}_{q}^{m^{u}\ell}$
$\mathcal{C}_{u}\cong\left(\bigoplus_{j=1}^{t}\left(D_{uj}^{\prime}\square C_{uj}^{\prime}\oplus D_{uj}^{{}^{\prime}\perp_{E}}\square C_{uj}^{{}^{\prime}\perp_{E}}\right)\right)\oplus\left(\bigoplus_{i=1}^{s-1}D_{ui}\square C_{ui}\right)\oplus D_{u-1}\square\mathcal{C}_{u-1}.$ (33)
where $C_{ui}$ is HSO for each $1\leq i\leq s-1$.
where each $D^{\prime}_{uj},D^{{}^{\prime}\perp_{E}}_{uj}$, and $D_{ui}$ are the cyclic subcodes of $D_{u}$ associated to $C^{\prime}_{uj},C^{{}^{\prime}\perp_{E}}_{uj}$, and $C_{ui}$ respectively, for $1\leq j\leq t$ for $1\leq i\leq s-1$, and $D_{u-1}$ is the cyclic code associated to $\mathcal{C}_{u-1}$. We will then utilize Jensen’s Bound (Theorem 4) to find a lower bound for the minimum distance $d(\mathcal{C}_{u})$ for each $u\geq 2$. In order to use Theorem 4 in conjunction with our recursive construction of $\mathcal{C}_{u}$ outlined in (27) and (28), the constituent codes of each $\mathcal{C}_{u}$ must satisfy the descending chain condition in (10). We first decompose $\mathcal{C}_{u}$ for $u\geq 2$ as
$\mathcal{C}_{u}\cong\bigoplus_{\lambda=1}^{2t+s}C_{u,\lambda}$
where the $C_{u,\lambda}$ are the nontrivial constituent codes of $\mathcal{C}_{u}$ whose minimum distances satisfy (10), noting that $\mathcal{C}_{u-1}=C_{u,\lambda}$ for some $\lambda$. In general, we do not have enough information to conclude how $d(\mathcal{C}_{u-1})$ compares with the minimum distances of the other constituent codes for $\mathcal{C}_{u}$ and do not know where it appears in the descending chain. Thus, we require the additional condition that the constituent code with the smallest minimum distance in $\mathcal{C}_{u}$ grows by a factor of $m$ from the smallest constituent code in $\mathcal{C}_{u-1}$, i.e.,
$d(\mathcal{C}_{u,2t+s})\geq md(\mathcal{C}_{u-1,2t+s}).$ (34)
This is not a major assumption as $\mathcal{C}_{u}\subset\mathbb{F}_{q}^{m^{u}\ell}$ and $\mathcal{C}_{u-1}\subset\mathbb{F}_{q}^{m^{u-1}\ell}$, so its minimum distance can be a factor of $m$ larger than $\mathcal{C}_{u-1}$. Repeated application of (34) yields
$d(\mathcal{C}_{u,2t+s})\geq m^{u-1}d_{1,s}$ (35)
where $d_{1,s}=d(C_{1,s})$ in (26). We emphasize that having such a condition will be vital in proving a lower bound for the minimum distances of $\mathcal{C}_{u}$ for each $u$ in the following lemma. This condition can be satisfied, for instance, by setting $C_{u,\lambda}=mC_{u-1,\lambda}$ for each $u\geq 2$ and $1\leq\lambda\leq 2t+s$. However, unlike in Lemma 3.19, taking $m$ copies of constituent codes is not required to obtain our result.
Lemma 3.21.
Let $R_{q,m}^{\ell}$ be as in (25). Then for $u\geq 1$, there exists an $m^{u}\ell$-length QC code $\mathcal{C}_{u}$ with constituent codes (27) and (28), with minimum distance $d(\mathcal{C}_{u})\geq m^{u-1}d_{1,s}$, where $d_{1,s}=d(C_{1,s})$.
Proof 3.22.
The case for $\mathcal{C}_{1}$ is trivial as we are assuming that $C_{1,s}$ has the smallest minimum distance in the representation of $\mathcal{C}_{1}$ in (26). For each $u\geq 2$, we will assume that the constituent codes in the decomposition of $\mathcal{C}_{u}$ are ordered according to the descending chain condition (10). The proof will require that we traverse our table horizontally by way of the descending chain condition to the last constituent then up the table vertically to use the inequality (35). For each $u\geq 2$, using the descending chain condition (10) and inequality (35), we have the following:
$N_{2t+s}(\mathcal{C}_{u})$ $=d(C_{u,2t+s})d(D_{2t+s})\geq md(C_{u-1,2t+s})\geq\ldots\geq m^{u-1}d(C_{1,s})=m^{u-1}d_{1,s},$
$N_{2t+s-1,2t+s}(\mathcal{C}_{u})$ $=d(C_{u,2t+s-1})d(D_{2t+s-1,2t+s})\geq d(C_{u,2t+s})\geq m^{u-1}d(C_{1,s})=m^{u-1}d_{1,s},$
$\vdots$
$N_{1,2,...,2t+s}(\mathcal{C}_{u})$ $=d(C_{u,1})d(D_{1,2,...,2t+s})\geq d(C_{u,2t+s})\geq m^{u-1}d(C_{1,s})=m^{u-1}d_{1,s}.$
Thus, $d(\mathcal{C}_{u})\geq d_{\mathrm{J}}(\mathcal{C}_{u})=\min\{N_{2t+s}(\mathcal{C}_{u}),N_{2t+s-1,2t+s}(\mathcal{C}_{u}),\ldots,N_{1,2,\ldots,2t+s}(\mathcal{C}_{u})\}\geq m^{u-1}d_{1,s}$ as desired. $\qed$
Example 3.23.
For sake of clarity, consider the following table of the decompositions of $\mathcal{C}_{u}$ for $2\leq u\leq 4$, where the decomposition is found in (29) and the constituent codes satisfy (34).
$u$ Decomposition
22 $\mathcal{C}_{2}\cong C_{2,1}\oplus C_{2,2}\oplus\cdots\oplus C_{2,2t+s}$
33 $\mathcal{C}_{3}\cong C_{3,1}\oplus C_{3,2}\oplus\cdots\oplus C_{3,2t+s}$
44 $\mathcal{C}_{4}\cong C_{4,1}\oplus C_{4,2}\oplus\cdots\oplus C_{4,2t+s}$
Table 1: Decompositions of $C_{2},\mathcal{C}_{3},\mathcal{C}_{4}$ to obtain lower bound for $d(\mathcal{C}_{4})$
Traversing horizontally through the row, we have that for $I=\{r,r+1,\ldots,2t+s\}\subseteq\{1,2,\ldots,2t+s\}$, $N_{I}(\mathcal{C}_{4})\geq d(C_{4,r})\geq d(C_{4,2t+s})$ by the descending chain condition, and then by equation  (34), traversing vertically up the table along the last constituent codes yields $N_{I}(\mathcal{C}_{4})\geq m^{3}d(C_{1,s})=m^{3}d_{1,s}$ for all $I$. Hence, $d(\mathcal{C}_{4})\geq m^{3}d_{1,s}.$
To obtain our lower bound, we include the condition that for each $u\geq 2$, and d(C’_uj^⟂_E) ≥md(C’_u-1,j^⟂_E), the last HSO constituent codes $C_{ui}$ in (28) satisfy
$d(C_{ui})\geq md(C_{u-1,i}),$ (36)
for all $1\leq i\leq s-1$. Note that repeated application of (34) yields $d(C_{ui})\geq m^{u-1}d(C_{1i})$ for each $i$.
proof We again use the induction argument along with the associated cyclic codes per (2.2). We first note that $D_{2t+s}$ is the $m$-length repetition code, so $d(D_{2t+s})=m$. For $u=1$, equation (16) and the fact that all nonzero constituent codes of $\mathcal{C}_{1}$ satisfy (10) yields
$N_{2t+s}(\mathcal{C}_{1})$ $=$ $d(C_{1s})d(D_{2t+s})=md_{1s}\geq d_{1s},$
$N_{2t+s-1,2t+s}(\mathcal{C}_{1})$ $=$ $d(C_{1,s-1})d(D_{2t+s-1,2t+s})\geq d(C_{1,s-1})\geq d_{1s},$
$\vdots$
$N_{1,2,...,2t+s}(\mathcal{C}_{1})$ $=$ $d(C^{\prime}_{11})d(D_{1,2,...,2t+s})\geq d(C_{1s})=d_{1s}.$
By Theorem 4, this shows $d(\mathcal{C}_{1})\geq\min\{N_{2t+s}(\mathcal{C}_{1}),N_{2t+s-1,2t+s}(\mathcal{C}_{1}),\ldots,N_{1,2,...,2t+s}(\mathcal{C}_{1})\}=d_{1s}$.
Now assume that for $u\in\mathbb{N}$, $d(\mathcal{C}_{u})\geq m^{u-1}d_{1s}$. Let $\mathcal{C}_{u+1}$ be the $m^{u+1}\ell$-length QC code constructed from $\mathcal{C}_{u}$, with constituent-code decomposition (3.20) and whose last constituent code is $\mathcal{C}_{u}$. At this point, we do not have information on how $d(\mathcal{C}_{u})$ compares with the minimum distances of the other constituent codes. As the minimum distance in any decomposition of a QC code should satisfy the descending chain in (10) to compute the Jensen bound, we consider the following two cases:
  • Case 1: $d(\mathcal{C}_{u})$ is not the smallest minimum distance.
    In this case, we notice that all terms
    $N_{2t+s}(\mathcal{C}_{u+1}),N_{2t+s-1,2t+s}(\mathcal{C}_{u+1}),...,N_{1,2,...,2t+s}(\mathcal{C}_{u+1})$
    are greater than or equal $d(C_{u+1,s-1})$ by definition of $N_{I}$, for $I\subseteq\{1,2,\ldots,2t+s\}$, in Theorem 16 and by the descending chain (10) for the constituent codes of $\mathcal{C}_{u+1}$. By (35), $d(C_{u+1,s-1})\geq m^{u}d(C_{1,s-1})$, which yields $m^{u}d(C_{1,s-1})\geq m^{u}d(C_{1,s})$ by the descending chain condition. Hence, $d(\mathcal{C}_{u+1})\geq d(C_{u+1,s-1})\geq m^{u}d_{1s}$.
  • Case 2: $d(\mathcal{C}_{u})$ is the smallest minimum distance
    For this case, we have
    $N_{2t+s}(\mathcal{C}_{u+1})$ $=$ $d(\mathcal{C}_{u})d(D_{2t+s})\geq m^{u-1}d_{1s}m=m^{u}d_{1s},\mbox{ from induction hypothesis,}$
    $N_{2t+s-1,2t+s}(\mathcal{C}_{u+1})$ $=$ $d(C_{u+1,s-1})d(D_{2t+s-1,2t+s})\geq m^{u}d(\mathcal{C}_{1,s-1})\geq m^{u}d_{1,s},$
    $\vdots$
    $N_{1,2,...,2t+s}(\mathcal{C}_{u+1})$ $=$ $d(C_{u+1,1}^{\prime})d(D_{1,2,...,2t+s})\geq d(C_{u+1,s-1})\geq m^{u}d(C_{1,s-1})\geq m^{u}d_{1s}.$
    Therefore by Theorem 4,
    $d(\mathcal{C}_{u+1})\geq\min\{N_{2t+s}(\mathcal{C}_{u+1}),N_{2t+s-1,2t+s}(\mathcal{C}_{u+1}),\ldots,N_{1,2,...,2t+s}(\mathcal{C}_{u+1})\}\geq m^{u}d_{1s},$
Thus for each $u\geq 1$, $d(\mathcal{C}_{u})\geq m^{u-1}d_{1s}$, as desired. dditionally, the minimum distances of the constituent codes of $\mathcal{C}_{u+1}$ over $\dfrac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}$ must satisfy equation (34), namely
$d(C_{u+1,i})\geq md(C_{u,i}),\quad\text{for all}\quad 1\leq i\leq s-1.$
At this point, we do not have information on how $d(\mathcal{C}_{u})$ compares with the minimum distances of the other constituent codes.
Suggestion: Using the Jensen Bound, the only time we really have to be careful is when $\mathcal{C}_{u}$ is the last constituent code. Otherwise, the definition of the $N_{I}$ and the descending chain of minimum distances causes the lower bound of $d_{J}(\mathcal{C}_{u})$ to be straightforwardly true. Hence, should we just consider two major case (Case 1: $\mathcal{C}_{u}$ is not the last constituent code, and Case 2: $\mathcal{C}_{u}$ is the last. Separate into these cases means we do not need to explciitly write out all $2t+s$ cases like below. Thoughts? G.: It makes sense!
Table showing the cases for where $C_{u}$ lives in $C_{u+1}$ according to the minimum distance descending chain condition:
$\mathcal{C}_{1}$
$\mathcal{C}_{2}$
$\mathcal{C}_{3}$
$\vdots$
$\mathcal{C}_{u+1}$
$\qed$
s the minimum distances in any decomposition of a QC code should satisfy the descending chain in  (10) to compute the GO-bound, we have the following $2t+s$ cases:
[…]
where $F_{i}:=\dfrac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}$, for each $1\leq i\leq s-1$, and $G^{\prime}_{j}:=\dfrac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle},$ $G^{\prime\prime}_{j}:=\dfrac{\mathbb{F}_{q}[x]}{\langle g^{*}_{j}\rangle}$ for each $1\leq j\leq t$. Notice that in each case, the respective QC code $\mathcal{C}_{u+1}$ is the same regardless of the ordering of its constituent codes (see (11)).
For the first 2t+s12t+s-1 cases, we notice that all terms $N_{2t+s}(\mathcal{C}_{u+1}),$ $N_{2t+s-1,2t+s}(\mathcal{C}_{u+1}),...,N_{1,2,...,2t+s}(\mathcal{C}_{u+1})$ are greater than or equal $d(C_{u+1,s-1})$ by definition of $N_{I}$ for $I\subseteq\{1,2,\ldots,2t+s\}$ in Theorem 16 and by the descending chain of minimum distances for the constituents of $\mathcal{C}_{u+1}$. As $d(C_{u+1,s-1})\geq m^{u}d(C_{1,s-1})\geq m^{u}d(C_{1,s})$, then $d(\mathcal{C}_{u})\geq d(C_{u+1,s-1})\geq m^{u}d_{1s}$.
For $(2t+s)$-th case, we have
$N_{2t+s}(\mathcal{C}_{u+1})$ $=$ $d(\mathcal{C}_{u})d(D_{2t+s})\geq m^{u-1}d_{1s}m=m^{u}d_{1s},\mbox{ from induction hypothesis,}$
$N_{2t+s-1,2t+s}(\mathcal{C}_{u+1})$ $=$ $d(C_{u+1,s-1})d(D_{2t+s-1,2t+s})\geq m^{u}d(\mathcal{C}_{1,s-1})\geq m^{u}d_{1,s},$
$\vdots$
$N_{1,2,...,2t+s}(\mathcal{C}_{u+1})$ $=$ $d(C_{u+1,1}^{\prime})d(D_{1,2,...,2t+s})\geq d(C_{u+1,s-1})\geq m^{u}d(C_{1,s-1})\geq m^{u}d(C_{1,s}).$
Therefore by Theorem 4,
$d(\mathcal{C}_{u+1})\geq\min\{N_{2t+s}(\mathcal{C}_{u+1}),N_{2t+s-1,2t+s}(\mathcal{C}_{u+1}),\ldots,N_{1,2,...,2t+s}(\mathcal{C}_{u+1})\}\geq m^{u}d_{1s},$
as desired.
The following theorem provides a summary of our construction of a family of ESO QC codes in $\mathbb{F}_{q}^{m^{u}\ell}$.
Theorem 3.24.
Let $R_{q,m}^{\ell}$ be as in (25). Let $\mathcal{C}_{u}$ be the $m^{u}\ell$-length QC code whose constituent codes are in (27) and (28), where $C_{u,i}=m^{u-1}C_{1,i}$, for all $1\leq i\leq s-1$, and that their minimum distances satisfy (10). Then $\{\mathcal{C}_{u}\}_{u\geq 1}$ is an infinite family of ESO $q$-ary QC codes with parameters
$\left[m^{u}\ell,\ell\left(\frac{m^{u}-1}{m-1}\right)\left(\sum_{j=1}^{t}\deg(g_{j})\right)+u\left(\sum_{i=1}^{s-1}\deg(f_{i})k_{1,i}\right)+k_{,1s},\geq m^{u-1}d_{1,s}\right]_{q}.$ (37)
Proof 3.25.
The parameters and duality of $\mathcal{C}_{u}$ comes from the recursive construction, and Lemmas 3.19 and 3.21. $\qed$
The next result is a particular case of Theorem 3.24 which allow us also to construct infinite families of not only ESO (or EDC) QC codes, but also QC codes that are ESD. From now on, we assume $m$ is odd when $q$ is odd (see Remark 2.9) so that the unique self-reciprocal irreducible factor of $x^{m}-1$ is $x-1\in\mathbb{F}_{q}[x]$. Hence in these special cases, we no longer need to use copies of the constituent codes of $\mathcal{C}_{u}$ to obtain the constituent codes of $\mathcal{C}_{u+1}$. Consequently, as we no longer have the dimensional restriction of $\dim(C_{u,i})=\dim(C_{1,i})=k_{1,i}$ for each $1\leq i\leq s-1$, we may allow for the $\mathcal{C}_{u}$ to ESD. Note that $\ell$ must be even in order for $\mathcal{C}_{u}$ to be ESD.
Corollary 3.26.
Let $R_{q,m}^{\ell}$ be as in (25) where the only self-reciprocal factor of $x^{m}-1$ is x1x-1. Let $\mathcal{C}_{1}\subset\mathbb{F}_{q}^{m\ell}$ be the ESO (resp. EDC or ESD) $\ell$-QC code whose constituent codes are $C_{1,j}\oplus C_{1,j}^{\perp_{E}}\subset\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}\right)^{\ell}$ for $1\leq j\leq t$, and an ESO (resp. EDC or ESD) code $C_{1,t+1}\subset\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{\ell}$ with parameters $\ell,k_{1,t+1},d_{1,t+1}]_{q}$. Suppose that $d(C_{1,1})\geq d(C_{1,1}^{\perp_{E}})\geq\cdots\geq d(C_{1,t})\geq d(C_{1,t}^{\perp_{E}})\geq d(C_{1,t+1}).$ Then there exists an infinite family $\{\mathcal{C}_{u}\}_{u\geq 1}$ of ESO (resp. EDC or ESD) $q$-ary QC codes with parameters
$\left[m^{u}\ell,\ell\left(\frac{m^{u}-1}{m-1}\right)\left(\sum_{j=1}^{t}\deg(g_{j})\right)+k_{1,t+1},\geq m^{u-1}d_{1,t+1}\right]_{q}.$ (38)
Proof 3.27.
This is a special case of Theorem 3.24 where $s=1$, and the unique self-reciprocal irreducible factor of $x^{m}-1$ is x1x-1. Hence,
$\sum_{j=1}^{t}\deg(g_{j})=\frac{1}{2}\left[(m-\left(\sum_{i=1}^{s-1}\deg(f_{i})\right)-1\right]=\frac{1}{2}(m-1),$ (39)
namely, ${\dim\mathcal{C}_{u}=\ell\left(\frac{m^{u}-1}{m-1}\right)\left(\sum_{j=1}^{t}\deg(g_{j})\right)+k_{1,t+1}=\frac{\ell}{2}\left(m^{u}-1\right)+k_{1,t+1}}$.
The duality of all constituent codes may be taken under the Euclidean inner product (see Remark 2.9 - item (i)) and it is defined exclusively by the $q$-ary constituent code $C_{1,t+1}$. Indeed, if $C_{1,t+1}$ is ESO (resp. EDC or ESD), then $\mathcal{C}_{1}$ also is. By the recursive construction, for any $u\geq 1$, the duality of $\mathcal{C}_{1}$ defines the duality of $\mathcal{C}_{2}$, and so on. This avoids relying on copies of constituent codes, which would otherwise restrict the construction to only ESO (or EDC) codes because of the dimensional limitations of self-dual and dual-containing codes. $\qed$
We note that if we want $x^{m}-1$ to have three distinct irreducible factors in $\mathbb{F}_{q}[x]$, of which x1x-1 is one, we need that $m$ is prime or the square of a prime. Otherwise, $m$ would have at least two distinct prime factors so that in the factorization
$x^{m}-1=\prod_{d\mid m}\Phi_{d}(x),$ (40)
where $\Phi_{d}(x)$ denotes the $d$-th cyclotomic polynomial, there would be at least four cyclotomic polynomial factors ($1$, $m$, and the distinct prime factors of $m$). This would result in at least four irreducible factors.
If $m$ is prime,
$x^{m}-1=\Phi_{1}(x)\Phi_{m}(x)=(x-1)\Phi_{m}(x),$ (41)
so we must consider the irreducible factors of $\Phi_{m}$. The polynomial $\Phi_{m}$ has $\frac{\phi(m)}{\operatorname{ord}_{m}(q)}=\frac{m-1}{\operatorname{ord}_{m}(q)}$ irreducible factors. Thus, if $x^{m}-1$ has three irreducible factors, we need $\operatorname{ord}_{m}(q)=\frac{m-1}{2}$, i.e. $q^{(m-1)/2}\equiv 1\ (\operatorname{mod}\ m)$.
If $m$ is the square of a prime, say $p^{2}$ where $p$ is prime,
$x^{m}-1=\Phi_{1}(x)\Phi_{p}(x)\Phi_{p^{2}}(x).$ (42)
We need that both $\Phi_{p}$ and $\Phi_{p^{2}}$ are both irreducible, i.e., $1=\frac{\phi(p)}{\operatorname{ord}_{p}(q)}=\frac{p-1}{\operatorname{ord}_{p}(q)}$ so that $\operatorname{ord}_{p}(q)=p-1$ and $1=\frac{\phi(p^{2})}{\operatorname{ord}_{p^{2}}(q)}=\frac{p(p-1)}{\operatorname{ord}_{p^{2}}(q)}$ so that ${\operatorname{ord}_{p^{2}}(q)}=p(p-1)$. Tables 2 and 3 provide some examples of $m$ and $q$ pairs where $x^{m}-1\in\mathbb{F}_{q}[x]$ decomposes into one pair of reciprocal polynomials and x1x-1. One caveat is that over $\mathbb{F}_{q^{2}}[x]$, i.e., when $q$ is the square of a prime, $\Phi_{p}$ is irreducible if and only if $p=2$ by Fermat’s little theorem, and so $x^{p^{2}}-1$ does not have three irreducible factors over $\mathbb{F}_{q^{2}}$.
$q$ $1\leq m\leq 100$
2 7, 17, 23, 41, 47, 71, 79, 97
3 11, 23, 37, 47, 59, 71, 83, 97
5 11, 19, 29, 41, 59, 61, 79, 89
7 3, 31, 47, 53, 59, 83
11 7, 53, 79, 83, 97
Table 2: Some values of $m$ and $q$ for which $x^{m}-1\in\mathbb{F}_{q}[x]$ decomposes into three irreducible factors over $\mathbb{F}_{q}$.
$q^{2}$ $1\leq m\leq 100$
4 3,5,7,11,13,19,23,29,37,47,53,59,61,67,71,79,83
9 5,7,11,17,19,23,29,31,43,47,53,59,71,79,83,89
25 3,7,11,17,19,23,37,43,47,53,59,73,79,83,97
49 3,5,11,13,17,23,31,41,47,59,61,67,71,79,83,89,97
121 3,7,13,17,23,29,31,41,47,59,67,71,73,79,83
Table 3: Some values of $m$ and $q^{2}$ for which $x^{m}-1\in\mathbb{F}_{q^{2}}[x]$ decomposes into three irreducible factors over $\mathbb{F}_{q^{2}}$.
Example 3.28.
Let
$R_{3,11}^{5}$ $\cong$ $\left(\frac{\mathbb{F}_{3}[x]}{\langle x^{5}+2x^{3}+x^{2}+2x+2\rangle}\right)^{5}\oplus\left(\frac{\mathbb{F}_{3}[x]}{\langle x^{5}+x^{4}+2x^{3}+x^{2}+2\rangle}\right)^{5}\oplus\left(\frac{\mathbb{F}_{3}[x]}{\langle x-1\rangle}\right)^{5}$
$=$ $\left(G_{1}^{\prime}\oplus G_{1}^{\prime\prime}\right)^{5}\oplus\left(\mathbb{F}_{3}\right)^{5},$
where $G_{1}^{\prime}=\mathbb{F}_{3}(\alpha)$. In addition, let $C_{1,1}^{\prime}=\langle(1,2,1,2,1),(\alpha,\alpha^{2},\alpha^{3},\alpha^{4},\alpha^{5})\rangle$ be the $5,2,4]_{3^{5}}-$code, $C_{1,1}^{\prime\prime}=C_{1,1}^{{}^{\prime}\perp_{E}}\subset(G_{1}^{\prime\prime})^{5}$ be the $5,3,3]_{3^{5}}-$code, and $C_{11}=\langle(1,1,1,0,0),(1,2,0,1,0)\rangle$ be the self-orthogonal $5,2,3]_{3}-$code. Let these codes be are the constituent codes of the QC code $\mathcal{C}_{1}$. For any $u\geq 2$, let the first two constituent codes of $\mathcal{C}_{u}$ be multiples of $\mathcal{C}_{1,1}^{\prime}$ and $C_{1,1}^{{}^{\prime}\perp_{E}}$ to obtain the $m^{u}\ell$-length QC codes
$\mathcal{C}_{u}\cong 11^{u-1}C_{1,1}^{\prime}\oplus 11^{u-1}C_{1,1}^{{}^{\prime}\perp_{E}}\oplus\mathcal{C}_{u-1}.$ (43)
According to Corollary 3.26, the $\{\mathcal{C}_{u}\}_{u\geq 1}$ is an infinite family of ESO QC codes with parameters
$\left[5\cdot 11^{u},\frac{5\cdot 11^{u}-1}{2},\geq 3\cdot 11^{u-1}\right]_{3}.$
4 Satisfying the Square-Root-Like Lower Bound
Using results from lingsole and developments presented in this paper, we show how the construction of codes in Section 3 satisfy the square-root-like lower bound and then present explicit examples. We first state the following helpful lemma.
Lemma 4.29.
If $m\geq 1$ and $u\geq 2$ are positive integers, then $\sqrt{m^{u}}\leq m^{u-1}.$
Proof 4.30.
If $u\geq 2$ , then $u\leq 2u-2$. For $m\geq 1$, this gives $m^{u}\leq m^{2u-2}$ and so $\sqrt{m^{u}}\leq m^{u-1}$. $\qed$
Theorem 4.31.
Let $\{\mathcal{C}_{u}\}_{u\geq 1}$ be the infinite family of ESO $q$-ary QC codes proposed in Theorem 3.24. Then
  • (i)
    If $d(C_{1,s})\geq c\sqrt{\ell}$ for some constant $c$, then all codes in $\{\mathcal{C}_{u}\}_{u\geq 2}$ satisfy the square-root-like bound for their minimum distances.
  • (ii)
    Otherwise, there exist a positive integer $L>1$ so that the all codes in the subfamily $\{\mathcal{C}_{u}\}_{u\geq L}$ of $\{\mathcal{C}_{u}\}_{u\geq 1}$ satisfies the square-root-like bound for their minimum distances.
Proof 4.32.
  • (i)
    By Lemma 4.29, for each $m^{u}\ell$-length QC code $\mathcal{C}_{u}$ with $u\geq 2$, we have
    $d(\mathcal{C}_{u})\geq m^{u-1}d_{1,s}\geq\sqrt{m^{u}}d_{1,s}\geq\sqrt{m^{u}}c\sqrt{\ell}=c\sqrt{m^{u}\ell},$
    namely, the minimum distances in $\{\mathcal{C}_{u}\}_{u\geq 2}$ satisfy the square-root-like bound and the result follows.
  • (ii)
    Notice that
    $m^{u-1}d_{1s}\geq c\sqrt{m^{u}\ell}\iff m^{2u-2}d_{1s}^{2}\geq c^{2}m^{u}\ell\iff m^{u-2}\geq\frac{c^{2}\ell}{d_{1s}^{2}}.$
    Hence, $u-2\geq\log_{m}\frac{c^{2}\ell}{d_{1s}^{2}}$, and when $u\geq\left\lceil\log_{m}\frac{c^{2}\ell}{d_{1s}^{2}}\right\rceil+2=:L$, the subfamily $\{\mathcal{C}_{u}\}_{u\geq L}$ of $\{\mathcal{C}_{u}\}_{u\geq 1}$ satisfies the square-root-like bound. $\qed$
In the special case where $R_{q,m}^{\ell}$ in (25) has only x1x-1 as the self-reciprocal factor of $x^{m}-1$, we obtain EDC and ESD codes that satisfying the square-root-like bound.
et $\{\mathcal{C}_{u}\}_{u\geq 1}$ be the infinite family of QC codes proposed in Corollary 3.26. As mentioned in the statement of that corollary, $\mathcal{C}_{u}$ being ESO, EDC, or ESD comes from the duality of the last constituent code over $\mathbb{F}_{q}$. Based on this observation, we have the following theorem.
Corollary 4.33.
Let $\{\mathcal{C}_{u}\}_{u\geq 1}$ be the infinite family of ESO/EDC/ESD $q$-ary QC codes proposed in Corollary 3.26. Then
  • (i)
    If $d(C_{1,t+1})\geq c\sqrt{\ell}$ for some constant $c$, then all codes in $\{\mathcal{C}_{u}\}_{u\geq 2}$ satisfy the square-root-like bound for their minimum distances.
  • (ii)
    Otherwise, there exist a positive integer $L>1$ so that the all codes in the subfamily $\{\mathcal{C}_{u}\}_{u\geq L}$ of $\{\mathcal{C}_{u}\}_{u\geq 1}$ satisfies the square-root-like bound for their minimum distances.
lightgray
Proof 4.34.
  • (i)
    According to Corollary 3.26 and the fact that the last constituent code of $\mathcal{C}_{1}$ is ESO (resp. EDC or ESD), it follows that all $\mathcal{C}_{u}$ are ESO (resp. EDC or ESD) as well, for $u\geq 2$. Additionally, from Lemma 4.29, for each $m^{u}\ell$-length QC code $\mathcal{C}_{u}$, with $u\geq 2$, we have
    $d(\mathcal{C}_{u})\geq m^{u-1}d_{1,t+1}\geq\sqrt{m^{u}}d_{1,t+1}\geq\sqrt{m^{u}}c\sqrt{\ell}=c\sqrt{m^{u}\ell},$
    namely, the minimum distances in $\{\mathcal{C}_{u}\}_{u\geq 2}$ satisfy the square-root-like bound and the result follows.
  • (ii)
    Notice that
    $m^{u-1}d_{1s}\geq c\sqrt{m^{u}\ell}\iff m^{2u-2}d_{1s}^{2}\geq c^{2}m^{u}\ell\iff m^{u-2}\geq\frac{c^{2}\ell}{d_{1s}^{2}}.$
    Hence, $u-2\geq\log_{m}\frac{c^{2}\ell}{d_{1s}^{2}}$, and when $u\geq\left\lceil\log_{m}\frac{c^{2}\ell}{d_{1s}^{2}}\right\rceil+2=:L$, the subfamily $\{\mathcal{C}_{u}\}_{u\geq L}$ of $\{\mathcal{C}_{u}\}_{u\geq 1}$ satisfies the square-root-like bound. $\qed$
Remark 4.35.
Any $c\leq m^{u/2-1}$ will work for square-root-like bound. If we would like a uniform $c$, then $c<\sqrt{m}$ works for all $u\geq 1$.
Remark 4.36.
Additionally, is straightforward to see that the lower-bound for the minimum distance in Theorem 3.24 is asymptotically better than the square-root-like lower bound. Indeed, given that $d_{1s},m,c$, and $\ell$ are positive integers, we have
$\lim_{u\rightarrow+\infty}\frac{m^{u-1}d_{1s}}{c\sqrt{m^{u}\ell}}=\lim_{u\rightarrow+\infty}\frac{m^{\frac{u}{2}-1}d_{1s}}{c\sqrt{\ell}}=+\infty>1.$
Example 4.37.
Let
$R_{5,11}^{6}$ $\cong$ $\left(\frac{\mathbb{F}_{5}[x]}{\langle x^{5}+2x^{4}+4x^{3}+x^{2}+x+4\rangle}\right)^{6}\oplus\left(\frac{\mathbb{F}_{5}[x]}{\langle x^{5}+4x^{4}+4x^{3}+x^{2}+3x+4\rangle}\right)^{6}\oplus\left(\frac{\mathbb{F}_{5}[x]}{\langle x-1\rangle}\right)^{6}$
$=$ $\left(G_{1}^{\prime}\oplus G_{1}^{\prime\prime}\right)^{6}\oplus\left(\mathbb{F}_{5}\right)^{6},$
where $G_{1}^{\prime}=\mathbb{F}_{5}(\alpha)$. Let $C_{1,1}^{\prime}$ be a $6,3,4]_{5^{5}}-$GRS code and $C_{1,1}^{{}^{\prime\prime}}=C_{1,1}^{{}^{\prime}\perp_{E}}$, then $C_{11}^{\prime\prime}$ is also a $6,3,4]_{5^{5}}-$GRS code. In addition, let $C_{1,1}=\langle(1,0,0,2,2,4),(0,1,0,2,4,2),(0,0,1,4,2,2)\rangle$ be a self-dual $6,3,4]_{5}$ code. Observe that all constituent codes are MDS. Therefore, according to Corollary 4.33, it is possible to get a infinite family $\{\mathcal{C}_{u}\}_{u\geq 2}$ of ESD QC codes with parameters $\left[4\cdot 11^{u},2\cdot 11^{u},\geq 4\cdot 11^{u-1}\right]_{5}$ for $u\geq 2$ that satisfy the square-root-like bound for their minimum distances.
Example 4.38.
Let $\alpha$ be a primitive $7$-th root of unity and let
$R_{2,7}^{8}$ $\cong$ $\left(\frac{\mathbb{F}_{2}[x]}{\langle x^{3}+x+1\rangle}\right)^{8}\oplus\left(\frac{\mathbb{F}_{2}[x]}{\langle x^{3}+x^{2}+1\rangle}\right)^{8}\oplus\left(\frac{\mathbb{F}_{2}[x]}{\langle x-1\rangle}\right)^{8}$
$\cong$ $\mathbb{F}_{2}^{8}(\alpha)\oplus\mathbb{F}_{2}^{8}(\alpha^{3})\oplus\mathbb{F}_{2}^{8}.$
Let $C_{1,1}^{\prime}$ be an $8,3,6]_{8}$ MDS code (whose construction can be checked in table), $C_{1,1}^{{}^{\prime\prime}}=C_{1,1}^{{}^{\prime}\perp_{E}}$ be an $8,5,4]_{8}$ (also MDS) code, and $C_{1,1}$ be an EDC $8,6,2]_{2}-$code with basis
$B=\{$ $(1,0,0,1,0,0,0,0),(0,1,0,1,0,0,0,0),(0,0,1,1,0,0,0,0),(0,0,0,0,1,0,0,1),(0,0,0,0,0,1,0,1),$
$(0,0,0,0,0,0,1,1)\}.$
ote that $d(D_{3})=7$, $d(D_{2,3})=3$, and $d(D_{1,2,3})=1$. So from Lemma 2 and Theorem 4, the code $\mathcal{C}_{1}$ is a $56,30,\geq 6]_{2}$ 88-QC code. The current best-known $56,30]_{2}$ code has minimum distance $10$. By Corollary 2.16, we obtain a quantum code $56,4,\geq 6]]_{2}$the best known minimum distance is $13$.
AA: I feel that the phrases I colored in red directly above short-sell our results because our lower-bound for the minimum distance is far away from the best known. G.: I agree
By Corollary 3.26, we obtain a $\left[8\cdot 7^{u},4\cdot 7^{u}+2,\geq 2\cdot 7^{u-1}\right]_{2}$ QC code. When $u=3$, $d\geq\sqrt{8\cdot 7^{u}}$, which satisfies the square-root bound.
G. Gustavo´s suggestion Note that $d(D_{3})=7$, $d(D_{2,3})=3$, and $d(D_{1,2,3})=1$. So from Lemma 2 and Theorem 4, the code $\mathcal{C}_{1}$ is a $56,30,\geq 6]_{2}$ 88-QC code. By Corollary 3.26, we obtain a $\left[8\cdot 7^{u},4\cdot 7^{u}+2,\geq 2\cdot 7^{u-1}\right]_{2}$ QC code.
When $u\geq 3$, $d\geq\sqrt{8\cdot 7^{u}}$, which shows that the infinite family $\{C_{u}\}_{u\geq 3}$ satisfies the square-root bound for the minimum distances. Moreover for $c=5$, we have $\left\lceil\log_{m}\frac{c^{2}\ell}{d_{1s}^{2}}\right\rceil+2=5$. Thus by Corollary 4.33, we then obtain an infinite family of QC codes $\{\mathcal{C}_{u}\}_{u\geq 5}$ whose minimum distances are greater than $c\sqrt{8\cdot 7^{u}}$.
We conclude this section by emphasizing that, alongside the seminal constructions of infinite families of self-dual cyclic codes satisfying the square-root-like bound in infiniteselfdualfamilies, our work introduces a novel construction of infinite families of QC codes that also attain this square-root-like lower bound on minimum distance.
5 Galois Closed QC Codes and Their Duals
Having QC codes whose Hermitian and Euclidean duals coincide is a valuable property, as it allows us to move freely between the two inner products. To achieve this property, we introduce the notion of the Galois closure of a code $C$ using notation from Jin2010: Given a positive integer $m\geq 1$ and $c=(c_{1},c_{2},...,c_{n})\in C\subset\mathbb{F}_{q}^{n}$ (in this case, $C$ is just a subset), we define $c^{m}=(c_{1}^{m},c_{2}^{m},...,c_{n}^{m})$ and $C^{m}=\{c^{m}:c\in C\}\subset\mathbb{F}_{q}^{n}.$ If $C$ is a subspace of $\mathbb{F}_{q}^{n}$ and $m$ is a power of the characteristic of $\mathbb{F}_{q}$, i.e., $m=q^{t}$ for some $t\in\mathbb{Z}^{+}$, then $C^{m}$ is also a subspace of $\mathbb{F}_{q}^{n}$.
Let $\mathbb{F}_{q}\subset\mathbb{F}_{q^{r}}$ and $C$ be an $\mathbb{F}_{q^{r}}$-code. We say that $C$ is Galois closed if $C=C^{q}$, where further on such codes details can be found in Bierbrauer. The Frobenius map $C\mapsto C^{q}$ respects the quasi-cyclic property of $C$ so that $C^{q}$ is also quasi-cyclic whenever $C$ is. Furthermore, if $\phi(C),\phi(C^{q})\subset\left(\frac{\mathbb{F}_{q^{2}}[x]}{\langle x^{m}-1\rangle}\right)^{\ell}$, then we can look at the constituent codes of $C$ and $C^{q}$ and assume that the irreducible factors of $x^{m}-1$ are aligned for both decompositions.
Remark 5.39.
For any code $C$ over $\mathbb{F}_{q^{2}}$, we observe that the Hermitian dual $C^{\perp_{H}}$ is equal to the Euclidean dual $\left(C^{q}\right){{}^{\perp_{E}}}$ of $C^{q}$. Hence, $C$ is HSO if and only if $C\subset(C^{q})^{\perp_{E}}$, i.e., $C^{q}\subset C^{\perp_{E}}$. Moreover, if $C$ is Galois closed, then $C^{\perp_{H}}=C^{\perp_{E}}$. Thus, $C$ is an HSO (resp. HDC, HSD) code if and only if $C$ is ESO (resp. EDC, ESD).
The following two propositions give a classification and construction for the Galois closed codes in $\mathbb{F}_{q^{2}}^{\ell}$, namely that they are direct sums of one-dimensional subspaces spanned by vectors with entries 0 or $v_{i}$ satisfying $v_{i}^{q-1}=\beta$, for some fixed $\beta\in\mathbb{F}_{q^{2}}$. This allows for easy construction of Galois closed codes and demonstrates that they are not rare but rather plentiful, making them indeed useful and desirable.
Proposition 5.40.
If $C\subset\mathbb{F}_{q^{2}}^{\ell}$ is Galois closed and $C=\langle\textbf{v}\rangle$ where $\textbf{v}=(v_{1},\ldots,v_{\ell})$, then $v_{i}=0$ or $v_{i}^{q-1}=\beta$ for some fixed $\beta\in\mathbb{F}_{q^{2}}$ for all $1\leq i\leq s$.
Proposition 5.41.
Let $C\subseteq\mathbb{F}_{q^{2}}^{\ell}$ be a $k$-dimensional code. Then $C$ is Galois closed if and only if $C=\bigoplus_{i}^{k}C_{i}$ where $C_{i}=C_{i}^{q}$ and $\dim(C_{i})=1$, i.e., $C$ is Galois closed if any only if can be decomposed as Galois closed codes of dimension $1$.
proposition Let $C\subset\mathbb{F}_{q^{2}}^{\ell}$ be a Galois closed code. Then $C$ is an HSO code if and only if $C$ is ESO.
Proof 5.42.
$C$ is HSO if and only if $C^{q}\subset C^{\perp_{E}}$. Since $C^{q}=C$, the result follows.
In general $C\neq C^{q}$, but they have the same parameters. According to Remark 5.39, if $C$ is Galois closed, then its Hermitian and Euclidean duals are exactly the same, i.e. $C^{\perp_{H}}=C^{\perp_{E}}$. Furthermore, the following theorem shows that if $C$ is Galois closed, then its Euclidean dual is also Galois closed.
Theorem 5.43.
A code $C$ over $\mathbb{F}_{q^{2}}$ is Galois closed if and only if its Euclidean dual is also Galois closed.
Proof 5.44.
We only prove the forward direction as the reverse direction is nearly identical. Let $x\in C$ and $y\in C^{\perp_{E}}$. Then
$\langle y^{q},x\rangle_{E}=\sum_{i}y_{i}^{q}x_{i}=\bigg(\sum_{i}y_{i}^{q}x_{i}\bigg)^{q}=\sum_{i}y_{i}x_{i}^{q}=\langle y,x^{q}\rangle_{E}.$
As $y\in C^{\perp_{E}}$, $y^{q}$ is in $(C^{\perp_{E}})^{q}$. Likewise, since $x\in C$ and $C=C^{q}$, we have that $\langle y,x^{q}\rangle_{E}=0$ and so $\langle y^{q},x\rangle_{E}=0$. Thus $y^{q}\in C^{\perp_{E}}$ , i.e., $y\in(C^{\perp_{E}})^{q}$, and so $C^{\perp_{E}}\subseteq(C^{\perp_{E}})^{q}$. The reverse direction follows similarly. $\qed$
Proposition 5.45.
Let $S=\left\{(c_{1},...,c_{n}),\left(c_{1}^{q},...,c_{n}^{q}\right),...,\left(c_{1}^{q^{r-1}},...,c_{n}^{q^{r-1}}\right)\right\}\subset\mathbb{F}_{q^{r}}^{n}$, and $C=\langle S\rangle$ be the respective $n,r\geq k]-$code. Then $C$ and its Euclidean dual are Galois closed.
Proof 5.46.
Without loss of generality, let $B=\{(c_{1},...,c_{n}),(c_{1}^{q},...,c_{n}^{q}),...,(c_{1}^{q^{k-1}},...,c_{n}^{q^{k-1}})\}=\{\mathbf{c_{0}},\mathbf{c_{1}},...,\mathbf{c_{k-1}}\}$ be a basis of $C$. Given $\mathbf{c_{\gamma}}=\sum_{i=0}^{k-1}\gamma_{i}\mathbf{c_{i}}$, where $\gamma_{i}\in\mathbb{F}_{q^{r}}$ for $0\leq i\leq k-1$, we have
$\mathbf{c_{\gamma}}^{q}$ $=$ $(\sum_{i=0}^{k-1}\gamma_{i}\mathbf{c_{i}})^{q}=\sum_{i=0}^{k-1}\gamma_{i}^{q}\mathbf{c_{i+1}}=\sum_{i=0}^{k-2}\gamma_{i}^{q}\mathbf{c_{i+1}}+\gamma_{k-1}^{q}\sum_{i=0}^{k-1}\beta_{i}\mathbf{c_{i}},\mbox{ where }\beta_{i}\in\mathbb{F}_{q^{r}},i=0,...,k-1;$ (44)
$=$ $\gamma_{k-1}^{q}\beta_{0}\mathbf{c_{0}}+\sum_{i=1}^{k-1}(\gamma_{i-1}^{q}+\beta_{i})\mathbf{c_{i}}\in C,$
namely, $C^{q}\subset C$. Since $C$ and $C^{q}$ have the same dimension; therefore, $C$ is Galois closed.
Finally, it is enough to prove that $C^{\perp_{E}}\subset(C^{q})^{\perp_{E}}$, since $(C^{q})^{\perp_{E}}\mbox{ and }C^{\perp_{E}}$ have the same dimension. Let $\mathbf{x}=(x_{1},...,x_{n})\in C^{\perp_{E}}$. Therefore,
$\langle\mathbf{x}^{q},\mathbf{c}_{\gamma}\rangle_{E}=\sum_{i=0}^{k-1}\langle\mathbf{x}^{q},\gamma_{i}\mathbf{c_{i}}\rangle_{E}=\left(\sum_{i=0}^{k-1}\langle\mathbf{x},\gamma_{i}^{q^{r-1}}\mathbf{c_{i-1}}\rangle_{E}\right)^{q}=0$ (45)
where $\mathbf{c_{-1}}$ is written as linear combination of $\mathbf{c_{0}},...,\mathbf{c_{k-1}}$. $\qed$
The Trace Representation (lingsole, Theorem 5.1) of $\ell$-QC codes depends on the constituent codes and $q$-cyclotomic cosets. It turns out that the the constituent codes of a QC code being Galois closed directly determines whether $\mathcal{C}\subset\mathbb{F}_{q}^{m\ell}$ is Galois closed.
Theorem 5.47.
An $\ell$-QC code $\mathcal{C}\subset\mathbb{F}_{q^{2}}^{m\ell}$ is Galois closed if and only if its respective constituent codes are Galois closed.
Proof 5.48.
Let $\mathcal{C}$ be an $\ell$-QC code with decomposition in (19). Then
$\mathcal{C}^{q}\cong\Bigg(\bigoplus_{j=1}^{t}({(C^{q})}{{}^{\prime}}_{j}\oplus{{(C^{q})}^{\prime}_{j}}^{\perp_{E}})\Bigg)\oplus\bigg(\bigoplus_{i=1}^{s}(C^{q})_{i}\bigg).$ (46)
As the corresponding constituent codes of $\mathcal{C}^{q}$ and $\mathcal{C}$ are aligned, we have that $(C^{q})^{\prime}_{j}=C^{\prime q}_{j}$ and ${(C^{q})^{\prime}_{j}}^{\perp_{E}}=((C^{\prime}_{j})^{\perp_{E}})^{q}$ for all $1\leq j\leq t$, and $(C^{q})_{i}=C_{i}^{q}$ for each $1\leq i\leq s$. Thus $\mathcal{C}=\mathcal{C}^{q}$ as codes if and only if $(C^{q})_{i}=C_{i}^{q}=C_{i}$ for each $1\leq i\leq s$, and $(C^{q})^{\prime}_{j}=C^{\prime q}_{j}=C^{\prime}_{j}$ and ${(C^{q})^{\prime}_{j}}^{\perp_{E}}=({C^{\prime}_{j}}^{\perp_{E}})^{q}={C^{\prime}_{j}}^{\perp_{E}}$ for all $1\leq j\leq t$. This can also be seen via the Trace Representation (lingsole, Theorem 5.1) in conjunction with properties of the trace map. $\qed$
We use these results to obtain the following family of $q^{2}$-ary QC codes that satisfy the square-root-like bound for their minimum distances.
Theorem 5.49.
Let $R_{q^{2},m}^{\ell}$ be as in (25). Let $\mathcal{C}_{u}$ be the $m^{u}\ell$-length QC code whose constituent codes in  (27) and (28) are all Galois closed and whose minimum distances satisfy (10). Additionally, assume that the constituent codes in (28) are HSO. Then we have an infinite family $\{\mathcal{C}_{u}\}_{u\geq 1}$ of HSO $q^{2}$-ary QC Galois closed codes where $d(\mathcal{C}_{u})\geq m^{u-1}d_{1,s}$ for each $u\geq 1$. Moreover,
  • (i)
    If $d(C_{1,t+1})\geq c\sqrt{\ell}$ for some constant $c$, then all codes in $\{\mathcal{C}_{u}\}_{u\geq 2}$ satisfy the square-root-like bound for their minimum distances.
  • (ii)
    Otherwise, there exist a positive integer $L>1$ so that the all codes in the subfamily $\{\mathcal{C}_{u}\}_{u\geq L}$ of $\{\mathcal{C}_{u}\}_{u\geq 1}$ satisfies the square-root-like bound for their minimum distances.
Remark 5.50.
Analogous results to those of Theorem 5.49 also hold when the constituent codes in (28) are HSD provided that $x^{m}-1$ has only x1x-1 as its self-reciprocal factor.
6 Constructions of QECCs from QC Codes
Given the construction of $\ell$-QC codes and results seen in cem; lingsole, along with the collection of quantum constructions in Subsection 2.4, we present in this section a construction of quantum codes with good parameters.
Example 6.51.
Let $q=4$, $\ell=3$, $m=7$, and $\alpha$ be a primitive $7$-th root of unity. Let
$R_{4,7}^{3}$ $\cong$ $\left(\frac{\mathbb{F}_{4}[x]}{\langle x^{3}+x+1\rangle}\right)^{3}\oplus\left(\frac{\mathbb{F}_{4}[x]}{\langle x^{3}+x^{2}+1\rangle}\right)^{3}\oplus\left(\frac{\mathbb{F}_{4}[x]}{\langle x-1\rangle}\right)^{3}$
$\cong$ $\mathbb{F}_{4}^{3}(\alpha)\oplus\mathbb{F}_{4}^{3}(\alpha^{3})\oplus\mathbb{F}_{4}^{3}.$
For any element $\gamma\in\mathbb{F}_{4}(\alpha)\setminus\mathbb{F}_{4}$, let $C_{1,1}^{\prime}=\left\langle(\gamma,\gamma,\gamma)\right\rangle\subset\mathbb{F}_{4}^{3}(\alpha)$ , $C_{1,1}^{\prime\prime}=(C_{1,1}^{\prime})^{\perp_{E}}\subset\mathbb{F}_{4}^{3}(\alpha^{3})$, and $C_{1,1}=\mathbb{F}_{4}^{3}$ be the constituent codes, whose parameters are $3,1,3]_{64}$, $3,2,2]_{64}$, and $3,3,1]_{4}$, respectively. Notice that $C_{1,1}^{\prime},C_{1,1}^{\prime\prime}$ are Galois closed by Proposition 5.45 and $C_{1,1}$ is trivially HDC and Galois closed. gray The minimum distance of the cyclic code associated to $\mathcal{C}$ and all its associated cyclic subcodes are described below
$\begin{array}[]{lclclclclcl}BZ(D_{1}^{\perp})&=&\{\alpha^{-1}=\alpha^{6}\}&\Rightarrow&D_{1}&=&\left\langle(x^{3}+x^{2}+1)(x+1)\right\rangle&\Rightarrow&d(D_{1})&=&4\\ BZ(D_{2}^{\perp})&=&\{\alpha^{-3}=\alpha^{4}\}&\Rightarrow&D_{2}&=&\left\langle(x^{3}+x+1)(x+1)\right\rangle&\Rightarrow&d(D_{2})&=&4\\ BZ(D_{3}^{\perp})&=&\{1\}&\Rightarrow&D_{3}&=&\left\langle(x^{3}+x+1)(x^{3}+x^{2}+1)\right\rangle&\Rightarrow&d(D_{3})&=&7\\ BZ(D_{1,2}^{\perp})&=&\{\alpha^{4},\alpha^{6}\}&\Rightarrow&D_{1,2}&=&\left\langle x+1\right\rangle&\Rightarrow&d(D_{1,2})&=&2\\ BZ(D_{1,3}^{\perp})&=&\{1,\alpha^{6}\}&\Rightarrow&D_{1,3}&=&\left\langle x^{3}+x^{2}+1\right\rangle&\Rightarrow&d(D_{1,3})&=&3\\ BZ(D_{2,3}^{\perp})&=&\{1,\alpha^{4}\}&\Rightarrow&D_{2,3}&=&\left\langle x^{3}+x+1\right\rangle&\Rightarrow&d(D_{2,3})&=&3\\ BZ(D_{1,2,3}^{\perp})&=&\{1,\alpha^{4},\alpha^{6}\}&\Rightarrow&D_{1,2,3}&=&\left\langle 1\right\rangle&\Rightarrow&d(D_{1,2,3})&=&1\end{array}$
WE NEED TO ADJUST THE REFERENCE BELOW TO JENSEN? There is a considerable difference between GO and Jensen bound for this case
Let $\mathcal{C}\subset\mathbb{F}_{4}^{m\ell}$ be the respective $3$-QC code which, by Theorem 5.47, is also Galois-closed. By Lemma 2 and Theorem 2.8, $\mathcal{C}$ is a $21,12,7]_{4}-$HDC (and also EDC) QC code, where the minimum distance was computed using Magma. By Corollary 2.16, it is possible to obtain a $21,3,d\geq 7]]_{4}-$quantum code $Q$ from the QC code $\mathcal{C}$. According to table, the best-known possible minimum distance for a quaternary QECC with $n=21$ and $k=3$ is $7$.
$16,8]]_{4} [[$$17,7]]_{4} [[$$18,6]]_{4} [[$$19,5]]_{4} [[$$20,4]]_{4}$
Ex. 6.51 and Lemma 2.17 - item (2) $\geq 2$ $\geq 3$ $\geq 4$ $\geq 5$ $\geq 6$
Code Table table 4-5 5-6 5-7 6-8 7-8
$22,3]]_{4} [[$$23,3]]_{4} [[$$24,3]]_{4} [[$$25,3]]_{4} [[$$26,3]]_{4}$
Ex. 6.51 and Lemma 2.17 - item (1) $\geq 7$ $\geq 7$ $\geq 7$ $\geq 7$ $\geq 7$
Code Table table 7-10 8-11 8-11 9-11 9-12
Table 4: Parameters of good QECCs from our code construction
By Lemma 2.17, it is still possible to construct more good QECCs from the code $\mathcal{C}$ which are listed in Table 4. Although there is no significant improvement in the minimum distances for any of the above parameters, most of our codes achieve values that are within one or two of the best-known lower bounds. It is also worth noting that our example exhibits lower bounds for the minimum distance, so it may be possible to obtain codes with larger minimum distances.
Example 6.52.
Still working with the decomposition of $R_{4,7}^{3}$ provided in Example 6.51, assume the same constituent codes $C_{1,1}^{\prime}$ and $C_{1,1}^{{}^{\prime\prime}}$ but let $C_{1,1}=\langle(0,1,\beta),(1,0,\beta^{2})\rangle$ be the last constituent code, where $\beta\in\mathbb{F}_{4}$ so that $\beta^{2}=\beta+1$. The code $C_{1,1}$ parameters $3,2,2]_{4}$ and is EDC.
Let $\mathcal{C}\subset\mathbb{F}_{4}^{m\ell}$ be the respective $3$-QC code. By Corollary 3.26, we have an infinity family $\{\mathcal{C}_{u}\}_{u\geq 1}$ of EDC QC codes whose their parameters are
$\left[3\cdot 7^{u},\frac{3\cdot 7^{u}+1}{2},\geq 2\cdot 7^{u-1}\right]_{4}.$}
From Corollary 2.16, it is possible to obtain the infinity family of quantum codes $\{Q_{u}\}_{u\geq 1}$ with parameters $\left[\left[3\cdot 7^{u},1,\geq 2\cdot 7^{u-1}\right]\right]_{4}$.
example Let $q=2$, $m=7$, $\ell=8$, and $\alpha$ be a primitive $7$-th root of unity. Assume the following decomposition of $R_{2,7}^{8}=\left(\frac{\mathbb{F}_{2}[x]}{\langle x^{7}-1\rangle}\right)^{8}$. Thus
$\left(\frac{\mathbb{F}_{2}[x]}{\langle x^{7}-1\rangle}\right)^{8}$ $\cong$ $\left(\frac{\mathbb{F}_{2}[x]}{\langle x^{3}+x+1\rangle}\right)^{8}\oplus\left(\frac{\mathbb{F}_{2}[x]}{\langle x^{3}+x^{2}+1\rangle}\right)^{8}\oplus\left(\frac{\mathbb{F}_{2}[x]}{\langle x-1\rangle}\right)^{8}$
$\cong$ $\mathbb{F}_{2}^{8}(\alpha)\oplus\mathbb{F}_{2}^{8}(\alpha^{3})\oplus\mathbb{F}_{2}^{8}.$
Let $C_{11}^{\prime}$ be a $8,3,6]_{8}$ MDS code (its construction can be checked in table), $C_{11}^{{}^{\prime\prime}}=C_{11}^{{}^{\prime}\perp_{E}}$ a $8,5,4]_{8}$ (also MDS) code, and $C_{11}$ be a EDC $8,6,2]_{2}-$code with basis
$B=\{$ $(1,0,0,1,0,0,0,0),(0,1,0,1,0,0,0,0),(0,0,1,1,0,0,0,0),(0,0,0,0,1,0,0,1),(0,0,0,0,0,1,0,1),$
$(0,0,0,0,0,0,1,1)\}$
ote that $d(D_{3})=7$, $d(D_{2,3})=3$, and $d(D_{1,2,3})=1$. So from Lemma 2 and Theorem 4, the code $\mathcal{C}_{1}$ is a $56,30,\geq 6]_{2}$ 88-QC code. The current best-known $56,30]_{2}$ code has minimum distance $10$. By Corollary 2.16, we obtain a quantum code $56,4,\geq 6]]_{2}$the best known minimum distance is $13$. AA: I feel that the phrases I colored in red directly above short-sell our results because our lower-bound for the minimum distance is far away from the best known. G.: I agree By Corollary 3.26, we obtain a $\left[8\cdot 7^{u},4\cdot 7^{u}+2,\geq 2\cdot 7^{u-1}\right]_{2}$ QC code. When $u=3$, $d\geq\sqrt{8\cdot 7^{u}}$, which satisfies the square-root bound. G. Gustavo´s suggestion ote that $d(D_{3})=7$, $d(D_{2,3})=3$, and $d(D_{1,2,3})=1$. So from Lemma 2 and Theorem 4, the code $\mathcal{C}_{1}$ is a $56,30,\geq 6]_{2}$ 88-QC code By Corollary 3.26, we obtain a $\left[8\cdot 7^{u},4\cdot 7^{u}+2,\geq 2\cdot 7^{u-1}\right]_{2}$ QC code. When $u=3$, $d\geq\sqrt{8\cdot 7^{u}}$, which satisfies the square-root bound. By Theorem 4.33, for $c=5$, we obtain $\left\lceil\log_{m}\frac{c^{2}\ell}{d_{1s}^{2}}\right\rceil+2=5$ and an infinite family $\{\mathcal{C}_{u}\}_{u\geq 5}$ with parameters
$\left[8\cdot 7^{u},4\cdot 7^{u}+2,\geq 2\cdot 7^{u-1}\right]_{2},$
where this bound to minimum distance is greater than $c\sqrt{n}$. magentaAA: I moved the last Example to the end of Section 4 as what’s impressive about it is that it satisfies the square-root and square-root-like lower bounds. As the classical code yields a family of quantum codes $\{\mathcal{Q}_{u}\}_{u\geq 2}$ with parameters
$\left[\left[8\cdot 7^{u},4,\geq 2\cdot 7^{u-1}\right]\right]_{2}.$
, I feel that it is not too impressive. as when $u=1$, we have a $\left[\left[56,4,2\right]\right]_{2}$ code, and the best known QECC for these pramaters has lower bound $13$. G.: Our first classical/quantum code is always poor due u1u-1, but it turns better from $u\geq 2$. By Corollary 2.16, we obtain a family of quantum codes $\{\mathcal{Q}_{u}\}_{u\geq 2}$ with parameters
$\left[\left[8\cdot 7^{u},4,\geq 2\cdot 7^{u-1}\right]\right]_{2}.$
7 Conclusion
In this work, we introduced a recursive construction of infinite families of quasi-cyclic codes whose parameters are determined by their constituent codes. By carefully selecting these constituents, we obtained families that are self-orthogonal, dual-containing, or self-dual with respect to the Euclidean and Hermitian inner products. A key feature of our construction is that both the dimension and a lower bound for the minimum distance can be explicitly determined from the parameters of the initial constituent codes, allowing us to study how the parameters evolve throughout the family. In particular, we showed that the resulting codes satisfy a square-root-like lower bound on the minimum distance which ensures a nontrivial distance growth as the code length increases. Moreover, we can use the CSS construction to establish the existence of quantum error-correcting codes with good parameters that arise from our families.
For future work, we aim to gain a deeper understanding of how using MDS codes as constituent codes affects the minimum distance of the respective QC Code. We also plan to explore sharper lower bounds for the minimum distance to improve the parameters of both the resulting QC code and its respective quantum code. Additionally, another plan is to relax the restriction in Section 3.2 that the constituent codes $C_{u,i}$ are given by $m^{u-1}$-copies of $C_{1,i}$. While this assumption simplified the analysis, allowing more general choices of $C_{u,i}$ may yield families with improved parameters. In each of these future directions, we aim to find the existence of quantum codes with good parameters.
n general, describing the parameters of infinity families of codes is not an easy task. In this paper, we have presented families of QC codes over $\mathbb{F}_{q}$, making explicit their dimension and a lower bound to their minimum distances. In addition, based on their constituent code decomposition, we also explored their dualities according to Euclidean and Hermitian inner products. We have also addressed when this lower bound to minimum distance is greater than the square-root-like bound, which there are only a few constructions in the literature achieving this goal. Finally, some quantum codes have been reported, which are comparable with those ones depicted on the Code Table table.
AA: this paragraph needs to be changed as we are no longer use GO bound. For future work, we aim to gain a deeper understanding of how using MDS codes as constituent codes affects the minimum distance of the respective QC code. It is also important to explore whether the GO-bound can be improved, since doing so would lead to better lower-bound estimates for the minimum distances of both our QC codes and their corresponding quantum codes.
8 Acknowledgements
The authors would like to thank the Institute for Computational and Experimental Research in Mathematics (ICERM) for providing funding and space for us to work in a mathematically rich environment during our week-long visit in Summer 2025. The first two authors would also like to thank Simons Laufer Mathematical Sciences Institute (SLMath) for providing funding for post-programmatic travel in 2024 following our Summer Research in Mathematics visit in 2023. The first author was supported by project FAPEMIG RED-00133-21. The second author was partially supported by the AMS-Simons Research Enhancement Grants for PUI Faculty while working on this project. The three authors also thank Dr. Lara Ismert for her conversations that inspired us to use Galois closed codes to move freely between the Hermitian and Euclidean inner products and Dr. Markus Grassl for reading the first version of this manuscript and pointing out some inconsistencies.
[…]